### v1

# ReZero: Boosting MCTS-based Algorithms by Just-in-Time and Speedy Reanalyze

Abstract
MCTS-based algorithms, such as MuZero and its derivatives, have achieved widespread success in various decision-making domains. These algorithms employ the reanalyze process to enhance sample efficiency, albeit at the expense of significant wall-clock time consumption. To address this issue, we propose a general approach named ReZero to boost MCTS-based algorithms. Specifically, we propose a new scheme that simplifies data collecting and reanalyzing, which significantly reduces the search cost while guarantees the performance as well. Furthermore, to accelerate each search process, we conceive a method to reuse the subsequent information in the trajectory. The corresponding analysis conducted on the bandit model also provides auxiliary theoretical substantiation for our design. Experiments conducted on Atari environments and board games demonstrates that ReZero substantially improves training speed while maintaining high sample efficiency. The code is available as part of the LightZero benchmark at https://github.com/opendilab/LightZero.
1 Introduction
As a pivotal subset of artificial intelligence, Reinforcement Learning (RL) (Sutton & Barto, 1988) has acquired achievements and applications across diverse fields, including interactive gaming (Vinyals et al., 2019), autonomous vehicles (Li et al., 2022), and natural language processing (Rafailov et al., 2023). Despite its successes, a fundamental challenge plaguing traditional model-free RL algorithms is their low sample efficiency. These algorithms typically require a large amount of data to learn effectively, making them often infeasible in the real world. In response to this problem, numerous model-based RL(Janner et al., 2019) (Hafner et al., 2020) methods have emerged. These methods involve learning an additional model of the environment from data and utilizing the model to assist the agent’s learning, thereby improving sample efficiency significantly.
Within this field, Monte Carlo Tree Search (Świechowski et al., 2023) (MCTS) has been proven to be a efficient method for utilizing models for planning. It incorporates the UCB1 algorithm (Auer et al., 2002) into the tree search process and has achieved promising results in in a wide range of application scenarios. Specifically, AlphaZero (Silver et al., 2017) plays a big role in combining deep RL with MCTS, achieving notable accomplishments that can beat top-level human players. While AlphaZero can only be applied to environments with perfect simulators, MuZero (Schrittwieser et al., 2019) extended the algorithm to cases without known environment models, resulting in good performances in a wider range of tasks. Following MuZero, many successor algorithms have emerged, enabling MuZero to be applied in continuous action spaces (Hubert et al., 2021), environments with stochastic dynamics (Antonoglou et al., 2021), offline RL training scenarios (Schrittwieser et al., 2021), and etc. All these MCTS-based algorithms (Silver et al., 2016, 2017; Schrittwieser et al., 2019; Ye et al., 2021) made valuable contributions to the universal applicability of the MCTS+RL framework.
However, the extensive tree search computations incur additional time overhead for these algorithms: During the data collection phase, the agent needs to execute MCTS to select an action every time it receives a new state. Furthermore, due to the characteristics of tree search, it is challenging to parallelize it using commonly used vectorized environments methods, further amplifying the speed disadvantage. On the other hand, during the reanalyze (Schrittwieser et al., 2021) process, in order to obtain higher-quality update targets, the latest model parameters are used to re-run MCTS on samples in buffer. The time required to process samples thus increases as a trade-off for high sample efficiency. The excessive time cost has become a bottleneck hindering the further promotion of these algorithms.
Recently, a segment of research endeavors are directed toward mitigating the above wall-clock time overhead about MCTS-based algorithms. On the one hand, SpeedyZero (Mei et al., 2023) greatly diminishes algorithms’ time overheads by deploying a parallel system; however, it demands more computational resources. On the other hand, it remains imperative to identify methodologies that accelerate these algorithms without imposing extra computational demands. PTSAZero (Fu et al., 2023), for instance, compresses the search space via state abstraction, decreasing the time cost per search. Nevertheless, it is widely acknowledged that MCTS-based algorithms necessitate search times to assure their performance. From this perspective, our research try to discover strategies that preserve algorithmic performance while minimizing the search times as much as possible. Moreover, given the observation that the tree search trajectories for states at adjacent time-steps frequently exhibit a significant degree of overlap, it becomes feasible to leverage the reuse of search information across different time-steps as a method to further expedite each search process.
In this paper, we introduce a new approach designed to boost the MCTS-based algorithms. Firstly, drawing inspiration from the utilization of a periodically updated target network in DQN (Mnih et al., 2013), we find that it is not essential to reanalyze the mini-batch at each iteration. Therefore, we propose a simplified framework which periodically reanalyzes the buffer to update policy targets in a just-in-time manner. Such an approach markedly diminishes the number of calls to MCTS. Secondly, within the reanalyze process, given the collected game trajectory, we propose to reuse search information from subsequent state to accelerate the search of the current state. In addition, we provide theoretical analysis on the potential implications of these modifications about the quality of target policy. Our approach can be easily applied to the MCTS-based algorithm family and can be synergistically implemented alongside existing acceleration methods. In addition, it will not bring about an increase in hardware overhead. Empirical experiments demonstrate that our approach yields good results in both single-agent environment (Atari (Bellemare et al., 2013)) and two-player board games (Silver et al., 2016), greatly improving the training speed while maintaining high sample efficiency. Ablation experiments explore the impact of different reanalyze periods and the acceleration effect of information reuse on a single search. The main contributions of this paper can be summarized as follows:
  • We propose a simplified framework that greatly reduces the number of calls to MCTS, thus speeding up MCTS-based algorithms.
  • We propose a method to speed up a single search by reusing search information. Theoretical support for the effectiveness of the method is also provided.
  • We conduct serveal experiments on diverse environments and further investigate our method through ablations.
2 Related Work
2.1 MCTS-based Algorithms
AlphaGo (Silver et al., 2016) and AlphaZero (Silver et al., 2017) combined MCTS with deep reinforcement learning, achieving significant results in board games, repeatedly defeating world-champion players, and revealing their remarkable capabilities. Following these achievements, MuZero (Schrittwieser et al., 2019) algorithm combines tree search with a learned value equivalent model (Grimm et al., 2020b), successfully promoting the algorithm’s application to scenarios without known models, such as Atari. Subsequent to MuZero’s development, several new works based on the MuZero framework emerged. Ye et al. (2021) further improved MuZero’s sample efficiency, using very little data for training and still achieving outstanding performance. Hubert et al. (2021) and Antonoglou et al. (2021) extended the MuZero agent to environments with large action spaces and stochastic environments, further broadening the algorithm’s usage scenario. Schrittwieser et al. (2021) applied MuZero’s reanalyze operation as a policy enhancement operator and extended it to offline settings. Danihelka et al. (2022) changed the node selection method during tree search, so the policy obtained from the search shows a monotonic improvement, ensuring algorithm performance even with low simulation numbers. Despite the aforementioned excellent improvement techniques, the current time overhead of MCTS-based RL remains significant, which is the main problem this paper aims to address.
2.2 MCTS Acceleration
Recent research has focused on accelerating MCTS-based algorithms. The work by Mei et al. (2023) significantly reduces the algorithm’s time overhead by designing a parallel system; however, this approach requires more computational resources and involves some adjustments for large batch training. Fu et al. (2023) narrows the search space through state abstraction, which amalgamates redundant spatial information, thereby reducing the time cost per search. Silver et al. (2016); Wu (2019) use information reuse to speed up the search process. They save sub-trees of the search tree and serve as initialization for the next search. However, this naive front-to-back information reuse is significantly different from our approach. To our knowledge, we are the first to enhance MCTS by reusing information in a backwards manner. Our proposed approach can seamlessly integrate with various MCTS-based algorithms, many of which (Danihelka et al., 2022; Mei et al., 2023; Fu et al., 2023; Leurent & Maillard, 2020) are orthogonal to our contributions.
3 Preliminaries
3.1 MuZero
MuZero (Schrittwieser et al., 2019) is a fundamental model-based RL algorithm that incorporates a value-equivalent model (Grimm et al., 2020a) and leverages Monte Carlo Tree Search for planning within a learned latent space. The model consists of three core components: a representation model $h_{\theta}$, a dynamics model $g_{\theta}$, and a prediction model $f_{\theta}$:
$\begin{array}[]{lll}\text{Representation:}&&s_{t}=h_{\theta}(o_{t-l:t})\\ \text{Dynamics:}&&s_{t+1},r_{t}=g_{\theta}(s_{t},a_{t})\\ \text{Prediction: }&&v_{t},p_{t}=f_{\theta}(s_{t})\\ \end{array}$ (1)
During the inference phase, the representation model transforms a sequence of the last $l$ observations $o_{t-l:t}$ into a corresponding latent state representation $s_{t}$. The dynamics model processes this latent state alongside an action $a_{t}$, yielding the subsequent latent state $s_{t+1}$ and an estimated reward $r_{t}$. Finally, the prediction model accepts a latent state and produces both the predicted policy $p_{t}$ and the state’s value estimate $v_{t}$. These outputs are instrumental in guiding the agent’s action selection process throughout its MCTS. Lastly the agent selects or samples the best action $a_{t}$ following the searched visit count distribution. During the training phase, given a training sequence $\{o_{t-l:t+K},a_{t+1:t+K},u_{t+1:t+K},\pi_{t+1:t+K},z_{t+1:t+K}\}$ at time $t$ sampled from the replay buffer, where $u_{t+k}$ denotes the actual reward obtained from the environment, $\pi_{t+k}$ represents the target policy obtained through MCTS during the agent-environment interaction, and $z_{t+k}$ is the value target computed using n-step bootstrapping (Hessel et al., 2018). The representation model initially converts the sequence of observations $o_{t-l:t}$ into the latent state $s_{t}^{0}$. Subsequently, the dynamic model executes $K$ latent space rollouts based on the sequence of actions $a_{t+1:t+K}$. The latent state derived after the $k$-th rollout is denoted as $s_{t}^{k}$, with the corresponding predicted reward indicated as $r_{t}^{k}$. Upon receiving $s_{t}^{k}$, the prediction model generates a predicted policy $p_{t}^{k}$ and a estimated value $v_{t}^{k}$. The final training loss encompasses three components: the policy loss ($l_{p}$), the value loss ($l_{v}$), and the reward loss ($l_{r}$):
$L_{\text{MuZero}}=$ $\sum_{k=0}^{K}l_{p}(\pi_{t+k},p_{t}^{k})+\sum_{k=0}^{K}l_{v}(z_{t+k},v_{t}^{k})+\sum_{k=1}^{K}l_{r}(u_{t+k},r_{t}^{k})$ (2)
MuZero Reanalyze, as introduced in (Schrittwieser et al., 2021), is an advanced iteration of the original MuZero algorithm. This variant enhances the model’s accuracy by conducting a fresh Monte Carlo Tree Search on sampled states with the most recent version of the model, subsequently utilizing the refined policy from this search to update the policy targets. Such reanalysis yields targets of superior quality compared to those obtained during the initial data collection phase. The Schrittwieser et al. (2021) expands upon this approach, formalizing it as a standalone method for policy refinement. This innovation opens avenues for its application in offline settings, where interactions with the environment are not possible. Traditional uses of the algorithm have intertwined reanalysis with training, periodically revisiting a subset of samples within a mini-batch for reevaluation. Furthermore, we suggests a paradigm shift, advocating for the decoupling of the reanalyze process from training routines, thus providing a more flexible and potentially more efficient methodology.
3.2 Stochastic Bandits
In the upcoming Section 4.2.2, we plan to offer an in-depth look at the theoretical assurances behind the performance of our newly designed ReZero agent, which is rooted in the fundamentals of bandit theory. Before diving into that, we’ll start with an easy-to-understand overview of the key concepts surrounding Stochastic Bandits.
A stochastic bandit $\nu$ with $K$ arms (Simchi-Levi et al., 2023) can be formulated as $\nu=\{P_{i}|i\in{1,2,...,K}\}$, where $P_{i}$ denotes the reward distribution corresponding to the arm $i$. The expected reward of arm $i$ is denoted as $\mu_{i}(\nu)$. Given an algorithm $\pi$ interacting with the bandit $\nu$ for $n$ rounds, we define the regret $R_{n}$ as follows:
$R_{n}(\pi,\nu)=n\mu^{*}(\nu)-\mathbb{E}[\sum_{t=1}^{n}X_{t}]$ (3)
where $\mu^{*}(\nu)=\max_{a\in[K]}\mu_{a}(\nu)$, $X_{t}$ denotes the random reward achieved at time $t$.
The regret quantifies the loss of reward incurred by $\pi$ compared to the optimal policy on $\nu$ during $n$ rounds of interactions. In practical problems, we aim for $\pi$ to exhibit sub-linear regret growth across a category of bandits $\Omega$ , i.e.:
$\lim_{n\to\infty}\frac{R_{n}(\pi,\nu)}{n}=0,\ \ \ \text{for all $\nu\in\Omega$ }$ (4)
where sub-linear growth rate indicates $\pi$ is choosing the optimal arm almost all of the time with $n$ tends to infinity. UCB (Auer et al., 2002) is a classical bandit algorithm that achieves regret growth rate of $O(\log(n))$.
AlphaZero (Silver et al., 2017) considers the child nodes in tree search as bandit arms and applies a variant of the UCB in the tree search process:
$UCB_{score}(s,a)=Q(s,a)+cP(s,a)\frac{\sqrt{{\sum}_{b}N(s,b)}}{1+N(s,a)}$ (5)
where $s$ is the state corresponding to the current node, $a$ is the action corresponding to a child node. $Q(s,a)$ is the mean return of choosing action $a$, $P(s,a)$ is the prior score of action $a$, $N(s,a)$ is the total time that $a$ has been chosen, ${\sum}_{b}N(s,b)$ is the total time that $s$ has been visited.
4 Method
Figure 1: Execution workflow and runtime cycle graph about MuZero and ReZero in both single and multiple worker cases. The number inside the modules represent the number of iterations, and the number under the modules represent the time required for module execution. The model is updated $n$ iterations between two collections. MuZero reanalyzes the mini-batch before each model update. ReZero reanalyzes the entire buffer after certain update iterations($\frac{n}{3}$ for example), which not only reduces the total number of MCTS calls, but also takes advantage of the processing speed of large batches.
In this section, we introduce two techniques to boost MCTS-based algorithms. Section 4.1 describes the design of just-in-time reanalyze, which achieves sufficient training on policy target meanwhile significantly reduces the number of MCTS executions. Furthermore, section 4.2 introduces speedy reanalyze: an acceleration method based on information reuse, which utilizes search results from the subsequent time-step to speed up the search at the previous time-step, thereby saving the time required for a single MCTS execution.
4.1 Just-in-Time Reanalyze: A Simplified Framework
Reanalyze, as a method to enhance the quality of training targets, was previously bound together with the training process in earlier implementations (Schrittwieser et al., 2021) (Figure 1). However, recomputing the target policy at each training iteration incurs massive MCTS computations. This dilemma motivates us to find a method to reduce the amount of MCTS calculations while ensuring performance.
We begin our detailed analysis with the following perspective: consider the reanalyze process as a policy improvement operator $\mathcal{R}_{\theta^{t}}$, which takes the current policy $P_{\theta^{t}}$ and transforms it into the target policy $T_{\theta^{t}}$ for training.
$T_{\theta^{t}}(a|s)=(\mathcal{R}_{\theta^{t}}P_{\theta^{t}})(a|s)=\frac{E_{T_{a}\sim\theta^{t}}[T_{a}]}{N}$ (6)
where $N$ is the number of simulations during the tree search and $E[T_{a}]$ is the expected total number that action $a$ was chosen. During the training process, we update the model parameter $\theta$ according to the following equation:
$\theta^{t+1}=\arg\min_{\theta}E_{s\sim\mathcal{D}}[\sum_{k=0}^{K}l^{p}(T_{\theta^{t}}(\cdot|s^{k})),P_{\theta}(\cdot|s^{k}))+l_{reg}]$ (7)
where $s$ is the state sampled from buffer $\mathcal{D}$, and $s^{k}$ is the state after $k$ unroll steps. We consider $l_{r}$ and $l_{v}$ in Equation 2 as a regularizer for $l_{p}$, and mark them as $l_{reg}$. In this way, the training can be seen as a projection operator $\mathcal{P}$ (Grill et al., 2020), projecting $T_{\theta^{t}}$ onto the policy space. The alternation of the reanalyze process and the training process constitutes policy iteration (Sutton & Barto, 1988), which can be seen as the composition of two operators:
$P_{\theta^{t+1}}=(\mathcal{P}\circ\mathcal{R}_{\theta^{t}})P_{\theta^{t}}$ (8)
In previous work (Schrittwieser et al., 2021), the reanalyze process is executed on the mini-batch of each training iteration. Changing the projection target frequently may result in an insufficient optimization, thereby degrading the quality of $\mathcal{P}$. In this way, even if the improvement operator $\mathcal{R}_{\theta}$ can produce high-quality training targets, the composite operator may not perform well. Therefore, we can adopt the practice of periodically updating the target network in DQN (Mnih et al., 2013), using Just-in-Time Reanalyze to change the projection targets only after several rounds of training.
Specifically, we propose the periodical reanalyze mechanism, i.e., we reanalyze the whole buffer after a fixed number of iterations. For each training iteration, we only need to sample the mini-batch and execute the gradient descent. Compared with previous methods, our method greatly reduces the number of MCTS, and can improve the parallelism due to the bigger search batch (Keskar et al., 2016), further enhanced the overall speed. Since the reanalyze process provides the target policy, the tree search during the collection phase is no longer necessary in the entire pipeline. Thus, we can bypass the tree search and directly select actions based on the policy, further reducing MCTS executions.
Based on our simplified framework, there are several potential advantages that may be explored in future work:
  • When directly using policy for data collection, action selection is no longer constrained by the tree search. As a result, previous vectorized environments such as (Weng et al., 2022) can be integrated to speed up. Besides, this design also makes MCTS-based algorithms compatible to existing RL exploration methods like (Badia et al., 2020).
  • Our method no longer needs to reanalyze the mini-batch for each iteration, thus decoupling the process of reanalyze and training. This provides greater scope for parallelization. In the case of multiple workers, we can design efficient parallelization paradigms as shown in Figure 1.
4.2 Speedy Reanalyze: Temporal Information Reuse
The above framework reduces the total number of MCTS. Next, we propose an temporal information reuse method to speed up the single tree search (Hereafter, termed information reuse for brevity). Section 4.2.1 introduces motivations and specific details. Section 4.2.2 carries out theoretical analysis of the method based on the bandit model.
4.2.1 Reuse the Information from Backward
As a local planning approach (Świechowski et al., 2023), MCTS is executed when receives each environment’s state. The agent, by performing tree search on the current state, can gain the root value and the visit counts: the former tells the quality of the state, and the later tells which actions are more likely to yield good returns. These can be seen as a kind of information. The trajectories traversed by the search of adjacent time-steps often have many overlaps, indicating that it involves similar knowledge. If the information gained at one time-step can be passed on to another, the search process can be accelerated: for example, if the search conducted at time-step $A$ reveals the quality of state $s$, then at time-step $B$, the agent need not to spend extra search cost to explore state $s$. Therefore, we designed a method to speed up the reanalyze process by reusing search information.
Figure 2: An illustration about the information reuse between adjacent time-steps. We search in the reverse order of the trajectory. After performing search on $s_{t+1}$, we obtain the search tree $T_{t+1}$ and the root value $m_{t+1}$. When searching on $s_{t}$, we set the value of the node corresponding to $s_{t+1}$ in search tree $T_{t}$ as $m_{t+1}$. If action $a_{t}$ is selected, the traverse process will be terminated, otherwise the traverse will proceed normally.
Algorithm 1 shows the specific design with Python-like code. During reanalyze, since the trajectories was already collected in the replay buffer, we can perform the search in a reverse order. That is to say, we first perform tree search on the state of the last time-step in the trajectory, and then search the second to last, and so on. After we search on state $s_{t+1}$  [^1] To facilitate the explanation, we have omitted details such as the latent space and do not make a deliberate distinction between states and nodes. (the state of the $t+1$-th time-step in the trajectory), the root value $m_{t+1}$ is obtained. When performing search on $s_{t}$, we set the value of $s_{t+1}$ to the fixed value $m_{t+1}$. During traverse in the tree, we select the action $a_{root}$ for root node $s_{t}$ according to the following equation:
$a_{root}=\arg\max_{a}I_{t}(a)$ (9)
$I_{t}(a)=\begin{cases}UCB_{score}(s_{t},a),&a\neq a_{t}\\ r_{t}+\gamma m_{t+1},&a=a_{t}\end{cases}$ (10)
where $a$ is an action corresponding to a child node, $a_{t}$ is the action corresponding to $s_{t+1}$, and $r_{t}$ is the reward predicted by the dynamic model. If an action other than $a_{t}$ is selected, the simulation just traverses following the original setting in MCTS (Schrittwieser et al., 2019). If action $a_{t}$ is selected, this simulation is terminated immediately. Since the time used to search for node $s_{t+1}$ is saved, the modified search process is faster than the original version, achieving Speedy Reanalyze.
Algorithm 1 Python-like code for information reuse
# trajectory_segment: a segment with length K
def search_backwards(trajectory_segment):
# prepare search context from the segment
roots, actions = prepare(trajectory_segment)
policy_targets = []
# search the roots backwards
for i in range(K, 1, -1):
if i == K:
# origin MCTS for Kth root
policy, value = origin_MCTS(roots[i])
else:
# reuse information from previous search
policy, value = reuse_MCTS(
roots[i], actions[i], value
policy_targets.append(policy)
# N: simulation numbers during one search
def reuse_MCTS(root, action, values)
for i in range(N):
# select an action for root node
a = select_root_child(root, action, value):
# early stop the simulation
if a == action:
backpropagate()
break
# traverse to the leaf node
else:
traverse()
backpropagate()
4.2.2 Theoretical Analysis
AlphaZero employs a variant of the UCB algorithm during MCTS search. Subsequently, MuZero proposes further improvements to this variant formula. Practical evidence has demonstrated the effectiveness of both these node selection methods; however, they have not been theoretically analyzed in their original paper. In earlier studies, UCT analyzed the convergence properties of the UCB1 algorithm in a tree structure, albeit without introducing a prior policy. PUCB (Rosin, 2011) incorporates relevant analysis of a prior policy, but it uses the prior in a penalty function-like approach which is different from in the MCTS-based algorithm family. In recent work, Gumbel MuZero (Danihelka et al., 2022) approximated the tree search model as a stochastic bandit, and on this basis, proved the monotonic improvement of their designs. In this section, we will draw upon such a simplification and demonstrate within the context of a stochastic bandit model why the node selection formula and policy updating methods used in the MCTS-based algorithm family are effective, and how information reuse can influence the policy target.
Definition 4.1.
After $n$ simulations with a stochastic bandit $\nu$ with $K$ arms, let $T_{i}$ denote the number of visits to arm $i$, then the policy generated based on the visit distribution is a random vector $P=[\frac{T_{1}}{n},...,\frac{T_{K}}{n}]$ The value $V_{\nu}(P)$ of policy $P$ on $\nu$ is defined as:
$V_{\nu}(P)=\sum_{i}\frac{T_{i}\mu_{i}(\nu)}{n}$ (11)
In cases where it does not cause any ambiguity, we will omit the mark $\nu$ in the following text. The next theorem explains the rationale behind using distribution as a target policy:
Theorem 4.2.
If an algorithm $\pi$ exhibits sub-linear growth in regret $R_{n}$ with respect to a bandit $\nu$, then the expectation $E[V(P)]$ converges to $\mu^{*}$ as $n$ tends to infinity.
$E[V(P)]$ $=\sum_{i}\frac{E[T_{i}]\mu_{i}}{n}$ (12)
$=\sum_{i}\frac{E[T_{i}](\mu^{*}-\Delta_{i})}{n}$ (13)
$=\mu^{*}-\frac{\sum_{i}E[T_{i}]\Delta_{i}}{n}$ (14)
where $\Delta_{i}=\mu^{*}-\mu_{i}$, and $\sum_{i}E[T_{i}]\Delta_{i}$ is the regret $R_{n}$ of the algorithm (see Appendix A).
Since $R_{n}$ growths in a sub-linear rate, so we can get
$\lim_{n\to\infty}E[V(P)]=\mu^{*}-\lim_{n\to\infty}\frac{R_{n}}{n}=\mu^{*}$ (15)
The following theorem provides an upper bound for the growth rate of $R_{n}$ for choosing arm in AlphaZero’s way , thereby demonstrating that the regret is sub-linearly increasing.
Theorem 4.3.
Denote the prior for arm $i$ as $P_{i}$. Without loss of generality, suppose $\mu^{*}=\mu_{1}$. Assume that the reward distribution of each arm $i$ is $1$-subgaussian and $P_{1}>\mu_{1}$, then the regret $R_{n}$ of AlphaZero satisfies (Appendix A)
$R_{n}\leq\sum_{i:\Delta_{i}>0}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(3+\frac{16}{\varepsilon^{4}P_{1}^{2}}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})$ (16)
The upper bound given by Theorem 4.3 is $O(\sqrt{n})$, which, while not reaching the level of $O(\log{n})$, is still sub-linear. Inequality (16) also indicates that a better policy which has higher probability to choose the best arm can lower the upper bound. Combining 4.2 and 4.3, we can conclude that with the increase in the number of simulations, the visit distribution will increasingly concentrate on the optimal arm, and the value corresponding to the policy will also converge to the value of the optimal arm. If the assumption that $P_{1}>\mu_{1}$ doesn’t hold, an upper bound with $O(\sqrt{n})$ rate can still be obtained by similar proof as long as there exists a $t$ that $P_{1}\sqrt{t-1}>\mu_{1}$, which indicates unlike some algorithms that are sensitive to prior policies, MCTS has strong robustness to prior policies. As long as the number of simulations is sufficient, it can still achieve a good target policy.
When reusing information in the search process, we can obtain a reference value for an arm. The existence of this reference value can eliminate the need for exploration of known arms, thereby further reducing the algorithm’s regret. The following theorem proves this point.
Theorem 4.4.
For a bandit which satisfies the assumptions in 4.3 and an already known arm value, the regret upper bound can be reduced if we use the way of 4.2.1 to choose arm. Specifically, if the known arm is the optimal arm with value $\mu_{1}$, then the regret $R_{n}$ satisfies:
$R_{n}\leq\sum_{i:\Delta_{i}>0}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(2+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})$ (17)
In other case, with the known value $\mu_{l}$ of a sub-optimal arm $l$, the regret $R_{n}$ satisfies:
$R_{n}$ $\leq\sum_{i:\Delta_{i}>0,i\neq l}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(3+\frac{16}{\varepsilon^{4}P_{1}^{2}}$ (18)
$+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})$ (19)
$+\inf_{\varepsilon\in(0,\Delta_{l})}\Delta_{l}(1+\frac{16}{\varepsilon^{4}P_{1}^{2}})$ (20)
The complete proof can be found in Appendix A.
Figure 3: Time-efficiency of ReZero-M vs. MuZero on six representative Atari games and two board games (Connect4, Gomoku). The horizontal axis represents Wall-clock Time (hours), while the vertical axis indicates the Episode Return over 5 episodes. ReZero-M achieves higher time-efficiency than the baseline on most Atari games and Gomoku, and achieve similar time-efficiency on some simpler game like Connect4. Mean of 5 runs, shaded areas are 95% confidence intervals.
Atari Board Games
avg. wall time (h) to 100k env. steps $\downarrow$ Pong Breakout MsPacman Qbert Seaquest UpNDown Connet4 Gomoku
ReZero-M (ours) $\mathbf{1}\scriptstyle{\pm 0.6}$ $\mathbf{7.7}\scriptstyle{\pm 0.3}$ $\mathbf{3.8}\scriptstyle{\pm 0.2}$ $\mathbf{5.1\scriptstyle{\pm 0.3}}$ $\mathbf{6.1}\scriptstyle{\pm 0.5}$ $\mathbf{6.9}\scriptstyle{\pm 0.5}$ $\mathbf{10}\scriptstyle{\pm 0.8}$ $\mathbf{9\scriptstyle{\pm 0.8}}$
MuZero (Schrittwieser et al., 2019) $4.1\scriptstyle{\pm 0.2}$ $12.4\scriptstyle{\pm 0.8}$ $12.0\scriptstyle{\pm 0.9}$ $10.2\scriptstyle{\pm 0.5}$ $11.75\scriptstyle{\pm 0.9}$ $12.4\scriptstyle{\pm 0.9}$ $11\scriptstyle{\pm 0.9}$ $23\scriptstyle{\pm 2}$
Table 1: Average wall-time of ReZero-M vs. MuZero on various tasks. (left) Six Atari games, (right) two board games. The time represents the average total wall-time to 100k environment steps for each algorithm. Mean and standard deviation over 5 runs. ReZero used two to four times less wall-time per 100k steps than the baseline method.
From the proof, it can be seen that if the real value $\mu_{1}$ of the optimal arm is known in advance, the regret caused by the underestimation of the optimal arm can be eliminated. However, the part of the regret that is reduced is $O(1)$. If the real value $\mu_{i}$ of a sub-optimal arm $l$ is known in advance, the regret caused by continuous exploration of this sub-optimal arm can be eliminated. The part of the reduced regret is $O(\sqrt{n})$. Intuitively, it can be understood this way: If the value of a sub-optimal arm $l$ is known, then after sufficient exploration of the optimal arm, the upper confidence bound will converge to be higher than $\mu_{l}$ with probability $1$, thus making the regret caused by arm $l$ no longer increases with $n$. While if the value of the optimal arm is known, then as the number of simulations $n$ increases, there will always be a moment when the upper confidence bound of the sub-optimal arm $i$ is greater than $\mu_{1}$, so arm $i$ will still be repeatedly selected, leading to the regret caused by arm $i$ increases with $n$. However, in this case, since the optimal arm is chosen the most time as $n$ increases, we can save more time by early stopping the search.
5 Experiment
Efficiency in RL usually refers to two aspects: sample efficiency, the agent’s ability to learn effectively from a limited number of environmental interactions, gauged by the samples needed to reach a successful policy using equal compute resources; time efficiency, which is how swiftly an algorithm learns to make optimal decisions, indicated by the wall-clock time taken to achieve a successful policy. Our primary goal is to improve time efficiency without compromising sample efficiency. To empirically validate the efficiency of the two boosting techniques introduced in Section 3.2 for MCTS-based algorithms, we have incorporated the ReZero with two prominent algorithms detailed in (Niu et al., 2023): MuZero with a Self-Supervised Learning loss (SSL) and EfficientZero, yielding the enhanced variants ReZero-M and ReZero-E respectively. For brevity, when the context is unambiguous, we simply refer to them as ReZero, omitting the symbol representing the baseline algorithm.
In order to validate the applicability of our method across multiple types of decision-making environments, we opted for six representative Atari environments characterized by classic image-input observation and discrete action spaces, in addition to the strategic board games Connect4 and Gomoku with special state spaces. For a baseline comparison, we employed the original implementations of MuZero and EfficientZero as delineated in the LightZero benchmark (Niu et al., 2023). The full implementation details available in Appendix B. Besides, we emphasize that to ensure an fair comparison of wall-clock time, all experimental trials were executed on a fixed single worker hardware settings.
In the next subsections, we sequentially explore and try to answer the following three questions:
  • How much can ReZero-M/ReZero-E improve time efficiency compared to MuZero/EfficientZero, while maintaining equivalent levels of high sample efficiency?
  • What is the concrete effect and hyper-parameter sensitivity of the Just-in-Time Reanalyze technique in ReZero?
  • How much search budgets and time costs can be saved in practice by the Temporal Information Reuse technique?
5.1 Time Efficiency
Setup In this section, we aim to evaluate the performance of ReZero against classical MCTS-based algorithms MuZero and EfficientZero, focusing on the wall-clock time reduction required to achieve the comparable performance level. In order to facilitate a fair comparison in terms of wall-time, for each run, we not only utilize identical computational resources but also maintain consistent hyper-parameter settings across the algorithms (unless otherwise specified cases). Key parameters are aligned with those from original papers. Specifically, we set the replay ratio (the ratio between environment steps and training steps) (D’Oro et al., 2022; Schwarzer et al., 2023) to 0.25, and the reanalyze ratio (the ratio between targets computed from interactions with the environment and by reanalysing existing data) (Schrittwieser et al., 2021) is set to 1. For detailed hyper-paramater configurations, please refer to the Appendix B.
Results Figure 3 illustrates the performance comparison in terms of wall-clock time between ReZero-M and the original MuZero across a variety of RL environments. It is apparent that ReZero-M achieves a significant improvement in time efficiency, reaching a near-optimal policy in substantially less time for the games Pong, Breakout and Seaquest, and obtained a better policy than MuZero on the remaining tasks. Additional results about ReZero-E and EfficientZero are detailed in Appendix C.2, specifically in Figure 6 and 7. For a more comprehensive understanding of our findings, we provide an in-depth comparison of the wall-clock training time up to 100k environment steps for ReZero-M against MuZero in Table 1. The data reveals that, on most games, ReZero require between two to four times less wall-clock time per 100k steps compared to the baseline methods.
5.2 Effect of Reanalyze Frequency
In next two sub-sectons, we will delve into details to explain how ReZero works. Firstly, we will analyze how the Just-in-Time Reanalyze technique manages to simplify the original iterative mini-batch reanalyze scheme to the periodical updated version while maintaining high sample efficiency. Here we adjust the periodical reanalyze frequency(reanalyze the buffer 1 time every $n$ training iterations) in ReZero to a lower or higher value on the MsPacman environment, specifically reanalyze frequency=$\{2,1,1/3,0\}$, where $1$ denotes the setting $n=1000$. The original MuZero variants with the reanalyze ratio of 1 is also included in this ablation experiment as a baseline. Figure 4 shows entire training curves in terms of Wall-time or Env Steps and validates that appropriate reanalyze frequency can save the time overhead without causing any obvious performance loss.
Figure 4: The effect of Reanalyze Frequency in ReZero-M on MsPacman. The horizontal axis in left represents Wall-clock Time (hours) while right side uses Env Steps. The vertical axis indicates Return. Proper reanalyze frequency can improve efficiency while obtaining the comparable Return with MuZero (reanalyze ratio=1).
5.3 Effect of Information Reuse
Indicators Avg. time (ms) tree search (num calls) dynamics model (num calls) data process (num calls)
ReZero-M $\textbf{0.69}\pm\textbf{0.02}$ 6089 122 277
MuZero $1.08\pm 0.09$ $13284$ $256$ $455$
Table 2: Comparisons about detailed indicators between ReZero-M and MuZero inside the tree search. The number of function calls is the cumulative value of 100 training iterations on Pong. Avg. time is the average time of a MCTS across all calls.
To further validate and understand the advantages and significance of the Information Reuse technique proposed in ReZero, we meticulously document a suite of statistical indicators of the tree search process in Table 2. Comparative analysis between ReZero-M and MuZero reveals that the Information Reuse technique reduces the invocation frequency of the dynamics model, the search tree, and other operations like data process transformations. Consequently, this advanced technique in leveraging subsequent time step data contributes to save the tree search time in various MCTS-based algorithms, further leading to overall wall-clock time gains. Besides, Information Reuse also incrementally reallocates computations (tree search simulations) towards previously non-visited nodes rather than repeatedly accessing extant nodes, which maybe yield some performance gains in some environments (e.g. Breakout in Figure 3).
6 Conclusion and Limitation
In this paper, we have delved into the efficiency and scalability of MCTS-based algorithms. Unlike most existing works, we incorporates information reuse and periodic reanalyze techniques to reduces wall-clock time costs while preserving sample efficiency. Our theoretical analysis and experimental results confirm that ReZero efficiently reduces the time cost and maintains or even improves performance across different decision-making domains. However, our current experiments are mainly conducted on the single worker setting, there exists considerable optimization scope to apply our approach into distributed RL training, and our design harbors the potential of better parallel acceleration and more stable convergence in large-scale training tasks. Also, the combination between ReZero and frameworks akin to AlphaZero, or its integration with some recent offline datasets such as RT-X (Padalkar et al., 2023), constitutes a fertile avenue for future research. These explorations could broadening the application horizons of MCTS-based algorithms.
7 Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Proof Materials
To verify the theoretical convergence guarantee of ReZero, we have provided some parts of detailed theoretical proofs in the main paper. In this section, we will present the complete supplementary proofs.
Lemma A.1.
Regret decomposition lemma:Denote $X_{t}$ as the reward at time $t$, $A_{t}$ as the arm chosen at time $t$, and $S_{n}=\sum_{t}X_{t}$, then the regret can be decomposed by regret caused by every sub-optimal arm:
$R_{n}=\sum_{i}E[T_{i}]\Delta_{i}$ (21)
Proof.
Note that
$R_{n}$ $=n\mu^{*}-E[S_{n}]$ (22)
$=\sum_{i}\sum_{t=1}^{n}E[(\mu^{*}-X_{t})\mathbb{I}\{A_{t}=i\}]$ (23)
and
$E[(\mu^{*}-X_{t})\mathbb{I}\{A_{t}=i\}|A_{t}]$ (24)
$=\mathbb{I}\{A_{t}=i\}E[\mu^{*}-X_{t}|A_{t}]$ (25)
$=\mathbb{I}\{A_{t}=i\}(\mu^{*}-\mu_{A_{t}})$ (26)
$=\mathbb{I}\{A_{t}=i\}(\mu^{*}-\mu_{i})$ (27)
$=\mathbb{I}\{A_{t}=i\}\Delta_{i}$ (28)
The proof can be completed by combining the two equations and using the definition of $T_{i}$. ∎
Lemma A.2.
Let $X_{1},...,X_{n}$ be a sequence of independent 1-subgaussian random variables, $\hat{\mu_{t}}=\frac{1}{t}\sum^{t}_{s=1}X_{s}$, $\varepsilon>0$, $a>0$ and
$\kappa=\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{t}+\sqrt{\frac{a}{t^{2}}}\geq\varepsilon\}$ (30)
, then it holds that $E[\kappa]\leq 1+\frac{2\sqrt{a}}{\varepsilon}+\frac{2}{\varepsilon^{2}}$.
Proof.
Take $u$ as $\frac{2\sqrt{a}}{\varepsilon}$, then
$E[\kappa]$ $\leq u+\sum_{t=\lceil u\rceil}^{n}\mathbb{P}(\hat{\mu}_{t}\sqrt{\frac{a}{t^{2}}}\geq\varepsilon)$ (31)
$\leq u+\sum_{t=\lceil u\rceil}^{n}\exp(-\frac{t(\varepsilon-\sqrt{\frac{a}{t^{2}}})^{2}}{2})$ (32)
[…]
$=1+u+\frac{2}{\varepsilon^{2}}=1+\frac{2\sqrt{a}}{\varepsilon}+\frac{2}{\varepsilon^{2}}$ (35)
Proof for Theorem 4.3:
Proof.
Denote $T_{i}(k)$ as the times that arm $i$ has been chosen until time $k$, $\hat{\mu}_{is}$ as the empirical mean based on the first $s$ samples from distribution $P_{i}$, and $\hat{\mu}_{i}(k)=\hat{\mu}_{iT_{i}(k)}$. Then we have:
$T_{i}(n)=\sum_{t=1}^{n}\mathbb{I}\{A_{t}=i\}$ (36)
$\leq\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}$ (37)
$+\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}$ (38)
for 37, we have
$E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]$ (39)
$\leq 1+\sum_{t=2}^{n}\sum_{s=0}^{t-1}\mathbb{P}(\hat{\mu}_{1s}+P_{1}\frac{\sqrt{t-1}}{1+s}\leq\mu_{1}-\varepsilon)$ (40)
$=1+\sum_{t=2}^{n}\sum_{s=1}^{t-1}\mathbb{P}(\hat{\mu}_{1s}+P_{1}\frac{\sqrt{t-1}}{1+s}\leq\mu_{1}-\varepsilon)$ (41)
$\leq 1+\sum_{t=2}^{n}\sum_{s=1}^{t-1}\exp(-\frac{s(\varepsilon+P_{1}\frac{\sqrt{t-1}}{1+s})^{2}}{2})$ (42)
$\leq 1+\sum_{t=2}^{n}\exp(-\frac{1}{2}\varepsilon P_{1}\sqrt{t-1})\sum_{s=1}^{t-1}\exp(-\frac{s\varepsilon^{2}}{2})$ (43)
$\leq 1+\sum_{t=2}^{n}\exp(-\frac{1}{2}\varepsilon P_{1}\sqrt{t-1})\frac{2}{\varepsilon^{2}}$ (44)
$\leq 1+\frac{16}{\varepsilon^{4}P_{1}^{2}}$ (45)
Notes: In 40, since we assume $P_{1}\geq\mu_{1}$, the probability of the term $s=0$ would be 00. Thus, we can discard it. If this assumption doesn’t hold, we can choose to accumulate $t$ starting from a larger $t_{0}$(which satisfies $P_{1}\sqrt{t_{0}-1}>\mu_{1}$ as mentioned in the article). Starting the summation from such a $t_{0}$ ensures all terms of $s=0$ can still be discarded, and all add terms that $t\leq t_{0}$ can be bounded to $1$. This only changes the constant term and won’t affect the growth rate of regret. From 42 to 43, we just need to expand the quadratic term and do some simple inequality scaling. From 43 to 44, we need to notice that $\sum_{s=1}^{t-1}\exp(-\frac{s\varepsilon^{2}}{2})$ is a geometric sequence and scale it to $\frac{2}{\varepsilon^{2}}$. From 44 to 45, we use the inequality $\sum_{t=2}^{n}\frac{1}{e^{a\sqrt{t-1}}}\leq\int_{1}^{\infty}\frac{1}{e^{a\sqrt{t-1}}}$.
for 38, we have
$E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (46)
$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\sqrt{\frac{n-1}{(1+T_{i}(t-1))^{2}}}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (47)
$\leq 1+E[\sum_{s=1}^{n-1}\mathbb{I}\{\hat{\mu}_{is}+P_{i}\sqrt{\frac{n-1}{(1+s)^{2}}}\geq\mu_{1}-\varepsilon\}]$ (48)
$\leq 1+E[\sum_{s=1}^{n}\mathbb{I}\{\hat{\mu}_{is}-\mu_{i}+P_{i}\sqrt{\frac{n-1}{s^{2}}}\geq\Delta_{i}-\varepsilon\}]$ (49)
with Lemma A.2, we can have
[…]
so
$E[T_{i}(n)]\leq 3+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{16}{\varepsilon^{4}P_{1}^{2}}$ (51)
combined with Lemma A.1, the proof is completed. ∎
Proof for Theorem 4.4:
Proof.
The proof can be obtained by making slight modifications to the proof of 4.3. In case that $\mu_{1}$ is already known,
$T_{i}(n)=\sum_{t=1}^{n}\mathbb{I}\{A_{t}=i\}$ (52)
$\leq\sum_{t=1}^{n}\mathbb{I}\{\mu_{1}\leq\mu_{1}-\varepsilon\}$ (53)
$+\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}$ (54)
$=\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}$ (55)
$\leq 2+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}$ (56)
In case that a sub-optimal value $\mu_{l}$ is known,
$T_{l}(n)=\sum_{t=1}^{n}\mathbb{I}\{A_{t}=l\}$ (57)
$\leq\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}$ (58)
$+\sum_{t=1}^{n}\mathbb{I}\{\mu_{l}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}$ (59)
$=\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}$ (60)
$\leq 1+\frac{16}{\varepsilon^{4}P_{1}^{2}}$ (61)
and the bound of $T_{i}(n)$ for $i\neq l$ keeps unchanged.
In both cases, combined with Lemma A.1, the proof is completed. ∎
Appendix B Implementation Details
B.1 Environments
In this section, we first introduce various types of reinforcement learning environments evaluated in the main paper and their respective characteristics, including different observation/action/reward space and transition functions.
Atari: This category includes sub-environments like Pong, Qbert, Ms.Pacman, Breakout, UpNDown, and Seaquest. In these environments, agents control game characters and perform tasks based on pixel input, such as hitting bricks in Breakout. With their high-dimensional visual input and discrete action space features, Atari environments are widely used to evaluate the capability of reinforcement learning algorithms in handling visual inputs and discrete control problems.
Board Games: This types of environment includes Connect4, Gomoku, where uniquely marked boards and explicitly defined rules for placement, movement, positioning, and attacking are employed to achieve the game’s ultimate objective. These environments feature a variable discrete action space, allowing only one player’s piece per board position. In practice, algorithms utilize action mask to indicate reasonable actions.
B.2 Algorithm Implementation Details
Our algorithm’s implementation is based on the open-source code of LightZero (Niu et al., 2023). Given that our proposed theoretical improvements are applicable to any MCTS-based RL method, we have chosen MuZero and EfficientZero as case studies to investigate the practical improvements in time efficiency achieved by integrating the ReZero boosting techniques: just-in-time reanalyze and speedy reanalyze (temporal information reuse).
To ensure an equitable comparison of wall-clock time, all experimental trials were executed on a fixed single worker hardware setting consisting of a single NVIDIA A100 GPU with 30 CPU cores and 120 GiB memory. Besides, we emphasize that to ensure a fair comparison of time efficiency and sample efficiency, the model architecture and hyper-parameters used in the experiments of Section 5 are essentially consistent with the settings in LightZero (Niu et al., 2023). For specific hyper-parameters of ReZero-M and MuZero on Atari, please refer to the Table 3. The main different hyper-parameters for the ReZero-M algorithm in the Connect4 and Gomoku environment are set out in Tables 4. In addition to employing an LSTM network with a hidden state dimension of 512 to predict the value prefix (Ye et al., 2021), all hyperparameters of ReZero-E are essentially identical to those of ReZero-M in Table 3.
Wall-time statistics Note that all our current tests are conducted in the single-worker case. Therefore, the wall-time reported in Table 1 and Appendix 5 for reaching 100k env steps includes:
  • collect time: The total time spent by an agent interacting with the environment to gather experience data.(Weng et al., 2022) can be integrated to speed up. Besides, this design also makes MCTS-based algorithms compatible to existing RL exploration methods like (Burda et al., 2018).
  • reanalyze time: The time used to reanalyze collected data with the current policy or value function for more accurate learning targets (Schrittwieser et al., 2021).
  • train time: The duration for performing updates to the agent’s policy, value functions and model based on collected data.
  • evaluation time: The period during which the agent’s policy is tested against the environment, separate from training, to assess performance.
Currently, we have set collect_max_episode_steps to 10,000 and eval_max_episode_steps to 20,000 to mitigate the impact of anomalously long evaluation episodes on time. In the future, we will consider conducting offline evaluations to avoid the influence of evaluation time on our measurement of time efficiency. Furthermore, the ReZero methodology represents a pure algorithmic enhancement, eliminating the need for supplementary computational resources or additional overhead. This approach is versatile, enabling seamless integration with single-worker serial execution environments as well as multi-worker asynchronous frameworks. The exploration of ReZero’s extensions and its evaluations in a multi-worker (Mei et al., 2023) paradigm are earmarked for future investigation.
Board games settings Given that our primary objective is to test the proposed techniques for improvements in time efficiency, we consider a simplified version of single-player mode in all the board games. This involves setting up a fixed but powerful expert bot and treating this opponent as an integral part of the environment. Exploration of our proposed techniques in the context of learning through self-play training pipeline is reserved for our future work.
Hyperparameter Value
Replay Ratio (Schwarzer et al., 2023) 0.25
Reanalyze ratio (Schrittwieser et al., 2021) 1
Num of frames stacked 4
Num of frames skip 4
Reward clipping (Mnih et al., 2013) True
Optimizer type Adam (Kingma & Ba, 2015)
Learning rate $3\times 10^{-3}$
Discount factor 0.997
Weight of policy loss 1
Weight of value loss 0.25
Weight of reward loss 1
Weight of policy entropy loss 0
Weight of SSL (self-supervised learning) loss (Ye et al., 2021) 2
Batch size 256
Model update ratio 0.25
Frequency of target network update 100
Weight decay $10^{-4}$
Max gradient norm 10
Length of game segment 400
Replay buffer size (in transitions) 1e6
TD steps 5
Number of unroll steps 5
Use augmentation True
Discrete action encoding type One Hot
Normalization type Batch Normalization (Ioffe & Szegedy, 2015)
Priority exponent coefficient (Schaul et al., 2015) 0.6
Priority correction coefficient 0.4
Dirichlet noise alpha 0.3
Dirichlet noise weight 0.25
Number of simulations in MCTS (sim) 50
Categorical distribution in value and reward modeling True
The scale of supports used in categorical distribution (Pohlen et al., 2018) 300
Table 3: Key hyperparameters of ReZero-M on Atari environments.
Hyperparameter Value
Replay Ratio (Schwarzer et al., 2023) 0.25
Reanalyze ratio (Schrittwieser et al., 2021) 1
Board size 6x7; 6x6
Num of frames stacked 1
Discount factor 1
Weight of SSL (self-supervised learning) loss 0
Length of game segment 18
TD steps 21; 18
Use augmentation False
Number of simulations in MCTS (sim) 50
The scale of supports used in categorical distribution 10
Table 4: Key hyperparameters of ReZero-M on Connect4 and Gomoku environments. If the parameter settings of these two environments are different, they are separated by a semicolon.
Appendix C Additional Experiments
C.1 Sample-efficiency of ReZero-M
Except for the results reported in main paper for time efficiency, we also shows the detailed sample efficiency results in this section. Figure 5 displays the performance over environment interaction steps of the ReZero-M algorithm compared with the original MuZero algorithm across six representative Atari environments and two board games. We can find that ReZero-M obtained similar sample efficiency than MuZero on the most tasks.
Figure 5: Sample-efficiency of ReZero-M vs. MuZero on six representative Atari games and two board games. The horizontal axis represents Env Steps, while the vertical axis indicates the Episode Return over 5 assessed episodes. ReZero achieves similar sample-efficiency than the baseline method. Mean of 5 runs; shaded areas are 95% confidence intervals.
C.2 ReZero-E
The enhancements of ReZero we have proposed are universally applicable to any MCTS-based reinforcement learning approach theoretically. In this section, we integrate ReZero with EfficientZero to obtain the enhanced ReZero-E algorithm. We present the empirical results comparing ReZero-E with the standard EfficientZero across four Atari environments.
Figure 6 shows that ReZero-E is better than EfficientZero in terms of time efficiency. Figure 7 indicates that ReZero-E matches EfficientZero’s sample efficiency across most tasks. Additionally, Table 5 details training times to 100k environment steps, revealing that ReZero-E is 0.5 to 2 times faster than baseline methods on most games.
Figure 6: Time-efficiency of ReZero-E vs. EfficientZero on four representative Atari games. The horizontal axis represents Wall-time (hours), while the vertical axis indicates the Episode Return over 5 assessed episodes. ReZero-E achieves higher time-efficiency than the baseline method. Mean of 3 runs; shaded areas are 95% confidence intervals.
Figure 7: Sample-efficiency of ReZero-E vs. EfficientZero on four representative Atari games. The horizontal axis represents Env Steps, while the vertical axis indicates the Episode Return over 5 assessed episodes. ReZero-E achieves similar sample-efficiency than the baseline method. Mean of 3 runs; shaded areas are 95% confidence intervals.
avg. wall time (h) to 100k env. steps $\downarrow$ Pong MsPacman Seaquest UpNDown
ReZero-E (ours) $\mathbf{2.3}\scriptstyle{\pm 1.4}$ $\mathbf{3}\scriptstyle{\pm 0.3}$ $\mathbf{3.1}\scriptstyle{\pm 0.1}$ $\mathbf{3.6}\scriptstyle{\pm 0.2}$
EfficientZero (Ye et al., 2021) $10\scriptstyle{\pm 0.2}$ $12\scriptstyle{\pm 1.3}$ $15\scriptstyle{\pm 2.3}$ $15\scriptstyle{\pm 0.7}$
Table 5: Average wall-time of ReZero-E vs. EfficientZero on four Atari games.The time represents the average total wall-time to 100k environment steps for each algorithm. Mean and standard deviation over 5 runs. ReZero-E used $\mathbf{0.5-2\times}$ less wall-time per 100k steps than baseline methods on most games.

### v2

@@
 Following MuZero, many successor algorithms have emerged, enabling MuZero to be applied in continuous action spaces (Hubert et al., 2021), environments with stochastic dynamics (Antonoglou et al., 2021), offline RL training scenarios (Schrittwieser et al., 2021), and etc.
-All these MCTS-based algorithms (Silver et al., 2016, 2017; Schrittwieser et al., 2019; Ye et al., 2021) made valuable contributions to the universal applicability of the MCTS+RL framework.
+All these MCTS-based algorithms (Silver et al., 2016, 2017; Schrittwieser et al., 2019) made valuable contributions to the universal applicability of the MCTS+RL framework.
 However, the extensive tree search computations incur additional time overhead for these algorithms: During the data collection phase, the agent needs to execute MCTS to select an action every time it receives a new state.
@@
 2.1 MCTS-based Algorithms
-AlphaGo (Silver et al., 2016) and AlphaZero (Silver et al., 2017) combined MCTS with deep reinforcement learning, achieving significant results in board games, repeatedly defeating world-champion players, and revealing their remarkable capabilities.
-Following these achievements, MuZero (Schrittwieser et al., 2019) algorithm combines tree search with a learned value equivalent model (Grimm et al., 2020b), successfully promoting the algorithm’s application to scenarios without known models, such as Atari.
+AlphaGo (Silver et al., 2016) and AlphaZero (Silver et al., 2017) combined MCTS with deep RL, achieving significant results in board games, defeating world-champion players, and revealing their remarkable capabilities.
+Following these achievements, MuZero (Schrittwieser et al., 2019) combines tree search with a learned value equivalent model (Grimm et al., 2020b), successfully promoting the algorithm’s application to scenarios without known models, such as Atari.
 Subsequent to MuZero’s development, several new works based on the MuZero framework emerged.
@@
 Combining 4.2 and 4.3, we can conclude that with the increase in the number of simulations, the visit distribution will increasingly concentrate on the optimal arm, and the value corresponding to the policy will also converge to the value of the optimal arm.
-If the assumption that $P_{1}>\mu_{1}$ doesn’t hold, an upper bound with $O(\sqrt{n})$ rate can still be obtained by similar proof as long as there exists a $t$ that $P_{1}\sqrt{t-1}>\mu_{1}$, which indicates unlike some algorithms that are sensitive to prior policies, MCTS has strong robustness to prior policies.
+If the assumption that $P_{1}>\mu_{1}$ doesn’t hold, an upper bound with $O(\sqrt{n})$ rate can still be obtained by similar proof as long as there exists a $t$ that $P_{1}\sqrt{t-1}>\mu_{1}$, which indicates unlike some algorithms that are sensitive to prior policies, MCTS has strong robustness to priors.
 As long as the number of simulations is sufficient, it can still achieve a good target policy.
@@
 Mean and standard deviation over 5 runs.
-ReZero used two to four times less wall-time per 100k steps than the baseline method.
 From the proof, it can be seen that if the real value $\mu_{1}$ of the optimal arm is known in advance, the regret caused by the underestimation of the optimal arm can be eliminated.
@@
 The data reveals that, on most games, ReZero require between two to four times less wall-clock time per 100k steps compared to the baseline methods.
+Figure 4: The effect of Reanalyze Frequency in ReZero-M on MsPacman.
+The horizontal axis in left represents Wall-clock Time (hours) while right side uses Env Steps.
+The vertical axis indicates Return.
+Proper reanalyze frequency can improve efficiency while obtaining the comparable Return with MuZero (reanalyze ratio=1).
 5.2 Effect of Reanalyze Frequency
@@
 Figure 4 shows entire training curves in terms of Wall-time or Env Steps and validates that appropriate reanalyze frequency can save the time overhead without causing any obvious performance loss.
-Figure 4: The effect of Reanalyze Frequency in ReZero-M on MsPacman.
-The horizontal axis in left represents Wall-clock Time (hours) while right side uses Env Steps.
-The vertical axis indicates Return.
-Proper reanalyze frequency can improve efficiency while obtaining the comparable Return with MuZero (reanalyze ratio=1).
 5.3 Effect of Information Reuse
@@
 This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+8 Acknowledgements
+Thanks to each of the authors for their efforts in the work of ReZero.
+We spent several memorable mid-nights and early mornings together in the lab.
+Thanks to everyone at OpenDILab and their pioneering work, especially the LightZero benchmark (https://github.com/opendilab/LightZero) for providing the base code for ReZero.
+Finally, we would like to thank Shanghai AI Laboratory and professor Jing Yang of Xi’an Jiaotong University for providing a platform for this work.
 Appendix A Proof Materials

### v3

@@
-ReZero: Boosting MCTS-based Algorithms by Just-in-Time and Speedy Reanalyze
+ReZero: Boosting MCTS-based Algorithms by Backward-view and Entire-buffer Reanalyze
 Abstract
-MCTS-based algorithms, such as MuZero and its derivatives, have achieved widespread success in various decision-making domains.
-These algorithms employ the reanalyze process to enhance sample efficiency, albeit at the expense of significant wall-clock time consumption.
-To address this issue, we propose a general approach named ReZero to boost MCTS-based algorithms.
-Specifically, we propose a new scheme that simplifies data collecting and reanalyzing, which significantly reduces the search cost while guarantees the performance as well.
-Furthermore, to accelerate each search process, we conceive a method to reuse the subsequent information in the trajectory.
-The corresponding analysis conducted on the bandit model also provides auxiliary theoretical substantiation for our design.
-Experiments conducted on Atari environments and board games demonstrates that ReZero substantially improves training speed while maintaining high sample efficiency.
+Monte Carlo Tree Search (MCTS)-based algorithms, such as MuZero and its derivatives, have achieved widespread success in various decision-making domains.
+These algorithms employ the reanalyze process to enhance sample efficiency from stale data, albeit at the expense of significant wall-clock time consumption.
+To address this issue, we propose a general approach named ReZero to boost tree search operations for MCTS-based algorithms.
+Specifically, drawing inspiration from the one-armed bandit model, we reanalyze training samples through a backward-view reuse technique which obtains the value estimation of a certain child node in advance.
+To further adapt to this design, we periodically reanalyze the entire buffer instead of frequently reanalyzing the mini-batch.
+The synergy of these two designs can significantly reduce the search cost and meanwhile guarantee or even improve performance, simplifying both data collecting and reanalyzing.
+Experiments conducted on Atari environments and board games demonstrate that ReZero substantially improves training speed while maintaining high sample efficiency.
 The code is available as part of the LightZero benchmark at https://github.com/opendilab/LightZero.
 1 Introduction
-As a pivotal subset of artificial intelligence, Reinforcement Learning (RL) (Sutton & Barto, 1988) has acquired achievements and applications across diverse fields, including interactive gaming (Vinyals et al., 2019), autonomous vehicles (Li et al., 2022), and natural language processing (Rafailov et al., 2023).
+As a pivotal subset of artificial intelligence, Reinforcement Learning (RL) [1] has acquired achievements and applications across diverse fields, including interactive gaming [2], autonomous vehicles [3], and natural language processing [4].
 Despite its successes, a fundamental challenge plaguing traditional model-free RL algorithms is their low sample efficiency.
-These algorithms typically require a large amount of data to learn effectively, making them often infeasible in the real world.
-In response to this problem, numerous model-based RL(Janner et al., 2019) (Hafner et al., 2020) methods have emerged.
+These algorithms typically require a large amount of data to learn effectively, making them often infeasible in the real-world scenarios.
+In response to this problem, numerous model-based RL [5, 6] methods have emerged.
 These methods involve learning an additional model of the environment from data and utilizing the model to assist the agent’s learning, thereby improving sample efficiency significantly.
-Within this field, Monte Carlo Tree Search (Świechowski et al., 2023) (MCTS) has been proven to be a efficient method for utilizing models for planning.
-It incorporates the UCB1 algorithm (Auer et al., 2002) into the tree search process and has achieved promising results in in a wide range of application scenarios.
-Specifically, AlphaZero (Silver et al., 2017) plays a big role in combining deep RL with MCTS, achieving notable accomplishments that can beat top-level human players.
-While AlphaZero can only be applied to environments with perfect simulators, MuZero (Schrittwieser et al., 2019) extended the algorithm to cases without known environment models, resulting in good performances in a wider range of tasks.
-Following MuZero, many successor algorithms have emerged, enabling MuZero to be applied in continuous action spaces (Hubert et al., 2021), environments with stochastic dynamics (Antonoglou et al., 2021), offline RL training scenarios (Schrittwieser et al., 2021), and etc.
-All these MCTS-based algorithms (Silver et al., 2016, 2017; Schrittwieser et al., 2019) made valuable contributions to the universal applicability of the MCTS+RL framework.
+Within this field, Monte Carlo Tree Search [7] (MCTS) has been proven to be a efficient method for utilizing models for planning.
+It incorporates the UCB1 algorithm [8] into the tree search process and has achieved promising results in in a wide range of scenarios.
+Specifically, AlphaZero [9] plays a big role in combining deep reinforcement learning with MCTS, achieving notable accomplishments that can beat top-level human players.
+While it can only be applied to environments with perfect simulators, MuZero [10] extended the algorithm to cases without known environment models, resulting in good performances in a wider range of tasks.
+Following MuZero, many successor algorithms have emerged, enabling MuZero to be applied in continuous action spaces [11], environments with stochastic dynamics [12], offline RL training scenarios [13], and etc.
+All these MCTS-based algorithms [9, 10, 14] made valuable contributions to the universal applicability of the MCTS+RL paradigm.
 However, the extensive tree search computations incur additional time overhead for these algorithms: During the data collection phase, the agent needs to execute MCTS to select an action every time it receives a new state.
-Furthermore, due to the characteristics of tree search, it is challenging to parallelize it using commonly used vectorized environments methods, further amplifying the speed disadvantage.
-On the other hand, during the reanalyze (Schrittwieser et al., 2021) process, in order to obtain higher-quality update targets, the latest model parameters are used to re-run MCTS on samples in buffer.
+Furthermore, due to the characteristics of tree search, it is challenging to parallelize it using commonly used vectorized environments [15], further amplifying the speed disadvantage.
+On the other hand, during the reanalyze [13] process, in order to obtain higher-quality update targets, the latest model parameters are used to re-run MCTS on the training mini-batch.
 The time required to process samples thus increases as a trade-off for high sample efficiency.
@@
 Recently, a segment of research endeavors are directed toward mitigating the above wall-clock time overhead about MCTS-based algorithms.
-On the one hand, SpeedyZero (Mei et al., 2023) greatly diminishes algorithms’ time overheads by deploying a parallel system; however, it demands more computational resources.
-On the other hand, it remains imperative to identify methodologies that accelerate these algorithms without imposing extra computational demands.
-PTSAZero (Fu et al., 2023), for instance, compresses the search space via state abstraction, decreasing the time cost per search.
-Nevertheless, it is widely acknowledged that MCTS-based algorithms necessitate search times to assure their performance.
+On the one hand, SpeedyZero [16] greatly diminishes algorithms’ time overheads by deploying a parallel execution training pipeline; however, it demands additional computational resources.
+On the other hand, it remains imperative to identify methodologies that accelerate these algorithms without imposing extra demands.
+PTSAZero [17], for instance, compresses the search space via state abstraction, decreasing the time cost per search.
+Nevertheless, it is widely acknowledged that MCTS-based algorithms necessitate search times (the number of simulations) to assure their performance.
 From this perspective, our research try to discover strategies that preserve algorithmic performance while minimizing the search times as much as possible.
-Moreover, given the observation that the tree search trajectories for states at adjacent time-steps frequently exhibit a significant degree of overlap, it becomes feasible to leverage the reuse of search information across different time-steps as a method to further expedite each search process.
-In this paper, we introduce a new approach designed to boost the MCTS-based algorithms.
-Firstly, drawing inspiration from the utilization of a periodically updated target network in DQN (Mnih et al., 2013), we find that it is not essential to reanalyze the mini-batch at each iteration.
-Therefore, we propose a simplified framework which periodically reanalyzes the buffer to update policy targets in a just-in-time manner.
-Such an approach markedly diminishes the number of calls to MCTS.
-Secondly, within the reanalyze process, given the collected game trajectory, we propose to reuse search information from subsequent state to accelerate the search of the current state.
-In addition, we provide theoretical analysis on the potential implications of these modifications about the quality of target policy.
-Our approach can be easily applied to the MCTS-based algorithm family and can be synergistically implemented alongside existing acceleration methods.
-In addition, it will not bring about an increase in hardware overhead.
-Empirical experiments demonstrate that our approach yields good results in both single-agent environment (Atari (Bellemare et al., 2013)) and two-player board games (Silver et al., 2016), greatly improving the training speed while maintaining high sample efficiency.
-Ablation experiments explore the impact of different reanalyze periods and the acceleration effect of information reuse on a single search.
+In this paper, we introduce ReZero, a new approach designed to boost the MCTS-based algorithms.
+Firstly, inspired by the one-armed bandit model [18], we propose a backward-view reanalyze technique that proceeds in the reverse direction of the trajectories, utilizing previously searched root values to bypass the exploration of specific child nodes, thereby saving time.
+We present several propositions to articulate our insights and analyzed the convergence of the algorithm based on the bandit model.
+Secondly, to better adapt to our proposed backward-view reanalyze technique, we have devised a novel pipeline that concentrates MCTS calls within the reanalyze process and periodically reanalyzes the entire buffer after a fixed number of training iterations.
+This entire-buffer reanalyze not only reduces the number of MCTS calls but also better leverages the speed advantages of parallelization.
+Our approach can be easily applied to the MCTS-based algorithm family and synergistically implemented alongside existing acceleration methods.
+In addition, it will not bring about any overhead in computation resources.
+Empirical experiments (Section 5) show that our approach yields good results in both single-agent environment (Atari [19]) and two-player board games [14], greatly improving the training speed while maintaining or even improving sample efficiency.
+Ablation experiments explore the impact of different reanalyze frequencies and the acceleration effect of backward-view reuse on a single search.
 The main contributions of this paper can be summarized as follows:
-  • We propose a simplified framework that greatly reduces the number of calls to MCTS, thus speeding up MCTS-based algorithms.
-  • We propose a method to speed up a single search by reusing search information. Theoretical support for the effectiveness of the method is also provided.
-  • We conduct serveal experiments on diverse environments and further investigate our method through ablations.
+  • We design a method to speed up a single search by the backward-view reanalyze technique. Theoretical support for the effectiveness of our proposed method is also provided.
+  • We propose a simplified framework with the entire-buffer reanalyze mechanims that further reduces the number of MCTS calls and enhance its parallelization, thus boosting MCTS-based algorithms.
+  • We conduct several experiments on diverse environments and investigate ReZero through ablations.
 2 Related Work
 2.1 MCTS-based Algorithms
-AlphaGo (Silver et al., 2016) and AlphaZero (Silver et al., 2017) combined MCTS with deep RL, achieving significant results in board games, defeating world-champion players, and revealing their remarkable capabilities.
-Following these achievements, MuZero (Schrittwieser et al., 2019) combines tree search with a learned value equivalent model (Grimm et al., 2020b), successfully promoting the algorithm’s application to scenarios without known models, such as Atari.
+AlphaGo [14] and AlphaZero [9] combined MCTS with deep RL, achieving significant results in board games, defeating world-champion players, and revealing their remarkable capabilities.
+Following these achievements, MuZero [10] combines tree search with a learned value equivalent model [20], successfully promoting the algorithm’s application to scenarios without known models, such as Atari.
 Subsequent to MuZero’s development, several new works based on the MuZero framework emerged.
-Ye et al.
-(2021) further improved MuZero’s sample efficiency, using very little data for training and still achieving outstanding performance.
-Hubert et al.
-(2021) and Antonoglou et al.
-(2021) extended the MuZero agent to environments with large action spaces and stochastic environments, further broadening the algorithm’s usage scenario.
-Schrittwieser et al.
-(2021) applied MuZero’s reanalyze operation as a policy enhancement operator and extended it to offline settings.
-Danihelka et al.
-(2022) changed the node selection method during tree search, so the policy obtained from the search shows a monotonic improvement, ensuring algorithm performance even with low simulation numbers.
+EfficientZero [21] further improved MuZero’s sample efficiency, using very little data for training and still achieving outstanding performance.
+Another works [11, 12] extended the MuZero agent to environments with large action spaces and stochastic environments, further broadening the algorithm’s usage scenario.
+MuZero Unplugged [13] applied MuZero’s reanalyze operation as a policy enhancement operator and extended it to offline settings.
+Gumbel MuZero [22] changed the node selection method during tree search, so the policy obtained from the search shows a monotonic improvement, ensuring algorithm performance even with low simulation numbers.
 Despite the aforementioned excellent improvement techniques, the current time overhead of MCTS-based RL remains significant, which is the main problem this paper aims to address.
@@
 Recent research has focused on accelerating MCTS-based algorithms.
-The work by Mei et al.
-(2023) significantly reduces the algorithm’s time overhead by designing a parallel system; however, this approach requires more computational resources and involves some adjustments for large batch training.
-Fu et al.
-(2023) narrows the search space through state abstraction, which amalgamates redundant spatial information, thereby reducing the time cost per search.
-Silver et al.
-(2016); Wu (2019) use information reuse to speed up the search process.
+The work [16] reduces the algorithm’s time overhead by designing a parallel system; however, this method requires more computational resources and involves some adjustments for large batch training.
+PTSAZero [17] narrows the search space through state abstraction, which amalgamates redundant information, thereby reducing the time cost per search.
+KataGo [23] use a naive information reuse trick.
 They save sub-trees of the search tree and serve as initialization for the next search.
-However, this naive front-to-back information reuse is significantly different from our approach.
-To our knowledge, we are the first to enhance MCTS by reusing information in a backwards manner.
-Our proposed approach can seamlessly integrate with various MCTS-based algorithms, many of which (Danihelka et al., 2022; Mei et al., 2023; Fu et al., 2023; Leurent & Maillard, 2020) are orthogonal to our contributions.
+However, our proposed backward-view reanalyze technique is fundamentally different from this naive forward-view reuse and can enhance search results while saving time.
+To our knowledge, we are the first to enhance MCTS by reusing information in a backward-view.
+Our proposed approach can seamlessly integrate with various MCTS-based algorithms, many of which [16, 17, 22, 24] are orthogonal to our contributions.
 3 Preliminaries
 3.1 MuZero
-MuZero (Schrittwieser et al., 2019) is a fundamental model-based RL algorithm that incorporates a value-equivalent model (Grimm et al., 2020a) and leverages Monte Carlo Tree Search for planning within a learned latent space.
+MuZero [10] is a fundamental model-based RL algorithm that incorporates a value-equivalent model [25] and leverages MCTS for planning within a learned latent space.
 The model consists of three core components: a representation model $h_{\theta}$, a dynamics model $g_{\theta}$, and a prediction model $f_{\theta}$:
 $\begin{array}[]{lll}\text{Representation:}&&s_{t}=h_{\theta}(o_{t-l:t})\\ \text{Dynamics:}&&s_{t+1},r_{t}=g_{\theta}(s_{t},a_{t})\\ \text{Prediction: }&&v_{t},p_{t}=f_{\theta}(s_{t})\\ \end{array}$ (1)
-During the inference phase, the representation model transforms a sequence of the last $l$ observations $o_{t-l:t}$ into a corresponding latent state representation $s_{t}$.
+The representation model transforms a last observation sequences $o_{t-l:t}$ into a corresponding latent state $s_{t}$.
 The dynamics model processes this latent state alongside an action $a_{t}$, yielding the subsequent latent state $s_{t+1}$ and an estimated reward $r_{t}$.
-Finally, the prediction model accepts a latent state and produces both the predicted policy $p_{t}$ and the state’s value estimate $v_{t}$.
-These outputs are instrumental in guiding the agent’s action selection process throughout its MCTS.
-Lastly the agent selects or samples the best action $a_{t}$ following the searched visit count distribution.
-During the training phase, given a training sequence $\{o_{t-l:t+K},a_{t+1:t+K},u_{t+1:t+K},\pi_{t+1:t+K},z_{t+1:t+K}\}$ at time $t$ sampled from the replay buffer, where $u_{t+k}$ denotes the actual reward obtained from the environment, $\pi_{t+k}$ represents the target policy obtained through MCTS during the agent-environment interaction, and $z_{t+k}$ is the value target computed using n-step bootstrapping (Hessel et al., 2018).
-The representation model initially converts the sequence of observations $o_{t-l:t}$ into the latent state $s_{t}^{0}$.
-Subsequently, the dynamic model executes $K$ latent space rollouts based on the sequence of actions $a_{t+1:t+K}$.
-The latent state derived after the $k$-th rollout is denoted as $s_{t}^{k}$, with the corresponding predicted reward indicated as $r_{t}^{k}$.
-Upon receiving $s_{t}^{k}$, the prediction model generates a predicted policy $p_{t}^{k}$ and a estimated value $v_{t}^{k}$.
-The final training loss encompasses three components: the policy loss ($l_{p}$), the value loss ($l_{v}$), and the reward loss ($l_{r}$):
-$L_{\text{MuZero}}=$ $\sum_{k=0}^{K}l_{p}(\pi_{t+k},p_{t}^{k})+\sum_{k=0}^{K}l_{v}(z_{t+k},v_{t}^{k})+\sum_{k=1}^{K}l_{r}(u_{t+k},r_{t}^{k})$ (2)
-MuZero Reanalyze, as introduced in (Schrittwieser et al., 2021), is an advanced iteration of the original MuZero algorithm.
-This variant enhances the model’s accuracy by conducting a fresh Monte Carlo Tree Search on sampled states with the most recent version of the model, subsequently utilizing the refined policy from this search to update the policy targets.
-Such reanalysis yields targets of superior quality compared to those obtained during the initial data collection phase.
-The Schrittwieser et al.
-(2021) expands upon this approach, formalizing it as a standalone method for policy refinement.
-This innovation opens avenues for its application in offline settings, where interactions with the environment are not possible.
-Traditional uses of the algorithm have intertwined reanalysis with training, periodically revisiting a subset of samples within a mini-batch for reevaluation.
-Furthermore, we suggests a paradigm shift, advocating for the decoupling of the reanalyze process from training routines, thus providing a more flexible and potentially more efficient methodology.
+Finally, the prediction model accepts a latent state and produces both the predicted policy $p_{t}$ and the state’s value $v_{t}$.
+These outputs are instrumental in guiding the agent’s action selection throughout its MCTS.
+Lastly the agent samples the best action $a_{t}$ following the searched visit count distribution.
+MuZero Reanalyze is an advanced version of the original MuZero algorithm.
+This variant enhances the model’s accuracy by conducting a fresh MCTS on sampled states with the most recent version of the model, subsequently utilizing the refined policy from this search to update the policy targets.
+Such reanalysis yields targets of superior quality compared to those obtained during the initial data collection.
+Traditional uses of the algorithm have intertwined reanalysis with training, while we suggests a novel paradigm, advocating for the decoupling of the reanalyze process from training iterations, thus providing a more flexible and efficient methodology.
 3.2 Stochastic Bandits
-In the upcoming Section 4.2.2, we plan to offer an in-depth look at the theoretical assurances behind the performance of our newly designed ReZero agent, which is rooted in the fundamentals of bandit theory.
-Before diving into that, we’ll start with an easy-to-understand overview of the key concepts surrounding Stochastic Bandits.
-A stochastic bandit $\nu$ with $K$ arms (Simchi-Levi et al., 2023) can be formulated as $\nu=\{P_{i}|i\in{1,2,...,K}\}$, where $P_{i}$ denotes the reward distribution corresponding to the arm $i$.
+In the upcoming Section 4.1, we plan to provide an inspiration case study behind our method, which is rooted in the fundamentals of bandit algorithms.
+Before diving into that, we’ll start with an overview of the key concepts surrounding stochastic bandits.
+A stochastic bandit $\nu$ with $K$ arms [26] can be formulated as $\nu=\{P_{i}|i\in{1,2,...,K}\}$, where $P_{i}$ denotes the reward distribution corresponding to the arm $i$.
 The expected reward of arm $i$ is denoted as $\mu_{i}(\nu)$.
-Given an algorithm $\pi$ interacting with the bandit $\nu$ for $n$ rounds, we define the regret $R_{n}$ as follows:
+Given an algorithm $\pi$ interacting with the bandit $\nu$ for $n$ rounds, we define the regret $R_{n}$:
 $R_{n}(\pi,\nu)=n\mu^{*}(\nu)-\mathbb{E}[\sum_{t=1}^{n}X_{t}]$ (2)
-where $\mu^{*}(\nu)=\max_{a\in[K]}\mu_{a}(\nu)$, $X_{t}$ denotes the random reward achieved at time $t$.
+where $\mu^{*}(\nu)=\max_{a\in[K]}\mu_{a}(\nu)$, $X_{t}$ denotes the random reward at time $t$.
 The regret quantifies the loss of reward incurred by $\pi$ compared to the optimal policy on $\nu$ during $n$ rounds of interactions.
-In practical problems, we aim for $\pi$ to exhibit sub-linear regret growth across a category of bandits $\Omega$ , i.e.:
+In practice, we aim for $\pi$ to exhibit sub-linear regret growth across a category of bandits $\Omega$ , i.e.:
 $\lim_{n\to\infty}\frac{R_{n}(\pi,\nu)}{n}=0,\ \ \ \text{for all $\nu\in\Omega$ }$ (3)
 where sub-linear growth rate indicates $\pi$ is choosing the optimal arm almost all of the time with $n$ tends to infinity.
-UCB (Auer et al., 2002) is a classical bandit algorithm that achieves regret growth rate of $O(\log(n))$.
-AlphaZero (Silver et al., 2017) considers the child nodes in tree search as bandit arms and applies a variant of the UCB in the tree search process:
+UCB [8] is a classical bandit algorithm that achieves regret growth rate of $O(\log(n))$.
+AlphaZero [9] considers the child nodes in tree search as bandit arms and applies a variant of the UCB in the tree search process:
 $UCB_{score}(s,a)=Q(s,a)+cP(s,a)\frac{\sqrt{{\sum}_{b}N(s,b)}}{1+N(s,a)}$ (4)
@@
 4 Method
-Figure 1: Execution workflow and runtime cycle graph about MuZero and ReZero in both single and multiple worker cases.
-The number inside the modules represent the number of iterations, and the number under the modules represent the time required for module execution.
-The model is updated $n$ iterations between two collections.
-MuZero reanalyzes the mini-batch before each model update.
-ReZero reanalyzes the entire buffer after certain update iterations($\frac{n}{3}$ for example), which not only reduces the total number of MCTS calls, but also takes advantage of the processing speed of large batches.
-In this section, we introduce two techniques to boost MCTS-based algorithms.
-Section 4.1 describes the design of just-in-time reanalyze, which achieves sufficient training on policy target meanwhile significantly reduces the number of MCTS executions.
-Furthermore, section 4.2 introduces speedy reanalyze: an acceleration method based on information reuse, which utilizes search results from the subsequent time-step to speed up the search at the previous time-step, thereby saving the time required for a single MCTS execution.
-4.1 Just-in-Time Reanalyze: A Simplified Framework
-Reanalyze, as a method to enhance the quality of training targets, was previously bound together with the training process in earlier implementations (Schrittwieser et al., 2021) (Figure 1).
-However, recomputing the target policy at each training iteration incurs massive MCTS computations.
-This dilemma motivates us to find a method to reduce the amount of MCTS calculations while ensuring performance.
-We begin our detailed analysis with the following perspective: consider the reanalyze process as a policy improvement operator $\mathcal{R}_{\theta^{t}}$, which takes the current policy $P_{\theta^{t}}$ and transforms it into the target policy $T_{\theta^{t}}$ for training.
-$T_{\theta^{t}}(a|s)=(\mathcal{R}_{\theta^{t}}P_{\theta^{t}})(a|s)=\frac{E_{T_{a}\sim\theta^{t}}[T_{a}]}{N}$ (6)
-where $N$ is the number of simulations during the tree search and $E[T_{a}]$ is the expected total number that action $a$ was chosen.
-During the training process, we update the model parameter $\theta$ according to the following equation:
-$\theta^{t+1}=\arg\min_{\theta}E_{s\sim\mathcal{D}}[\sum_{k=0}^{K}l^{p}(T_{\theta^{t}}(\cdot|s^{k})),P_{\theta}(\cdot|s^{k}))+l_{reg}]$ (7)
-where $s$ is the state sampled from buffer $\mathcal{D}$, and $s^{k}$ is the state after $k$ unroll steps.
-We consider $l_{r}$ and $l_{v}$ in Equation 2 as a regularizer for $l_{p}$, and mark them as $l_{reg}$.
-In this way, the training can be seen as a projection operator $\mathcal{P}$ (Grill et al., 2020), projecting $T_{\theta^{t}}$ onto the policy space.
-The alternation of the reanalyze process and the training process constitutes policy iteration (Sutton & Barto, 1988), which can be seen as the composition of two operators:
-$P_{\theta^{t+1}}=(\mathcal{P}\circ\mathcal{R}_{\theta^{t}})P_{\theta^{t}}$ (8)
-In previous work (Schrittwieser et al., 2021), the reanalyze process is executed on the mini-batch of each training iteration.
-Changing the projection target frequently may result in an insufficient optimization, thereby degrading the quality of $\mathcal{P}$.
-In this way, even if the improvement operator $\mathcal{R}_{\theta}$ can produce high-quality training targets, the composite operator may not perform well.
-Therefore, we can adopt the practice of periodically updating the target network in DQN (Mnih et al., 2013), using Just-in-Time Reanalyze to change the projection targets only after several rounds of training.
-Specifically, we propose the periodical reanalyze mechanism, i.e., we reanalyze the whole buffer after a fixed number of iterations.
-For each training iteration, we only need to sample the mini-batch and execute the gradient descent.
-Compared with previous methods, our method greatly reduces the number of MCTS, and can improve the parallelism due to the bigger search batch (Keskar et al., 2016), further enhanced the overall speed.
-Since the reanalyze process provides the target policy, the tree search during the collection phase is no longer necessary in the entire pipeline.
-Thus, we can bypass the tree search and directly select actions based on the policy, further reducing MCTS executions.
-Based on our simplified framework, there are several potential advantages that may be explored in future work:
-  • When directly using policy for data collection, action selection is no longer constrained by the tree search.
-As a result, previous vectorized environments such as (Weng et al., 2022) can be integrated to speed up.
-Besides, this design also makes MCTS-based algorithms compatible to existing RL exploration methods like (Badia et al., 2020).
-  • Our method no longer needs to reanalyze the mini-batch for each iteration, thus decoupling the process of reanalyze and training.
-This provides greater scope for parallelization.
-In the case of multiple workers, we can design efficient parallelization paradigms as shown in Figure 1.
-4.2 Speedy Reanalyze: Temporal Information Reuse
-The above framework reduces the total number of MCTS.
-Next, we propose an temporal information reuse method to speed up the single tree search (Hereafter, termed information reuse for brevity).
-Section 4.2.1 introduces motivations and specific details.
-Section 4.2.2 carries out theoretical analysis of the method based on the bandit model.
-4.2.1 Reuse the Information from Backward
-As a local planning approach (Świechowski et al., 2023), MCTS is executed when receives each environment’s state.
-The agent, by performing tree search on the current state, can gain the root value and the visit counts: the former tells the quality of the state, and the later tells which actions are more likely to yield good returns.
-These can be seen as a kind of information.
-The trajectories traversed by the search of adjacent time-steps often have many overlaps, indicating that it involves similar knowledge.
-If the information gained at one time-step can be passed on to another, the search process can be accelerated: for example, if the search conducted at time-step $A$ reveals the quality of state $s$, then at time-step $B$, the agent need not to spend extra search cost to explore state $s$.
-Therefore, we designed a method to speed up the reanalyze process by reusing search information.
-Figure 2: An illustration about the information reuse between adjacent time-steps.
-We search in the reverse order of the trajectory.
-After performing search on $s_{t+1}$, we obtain the search tree $T_{t+1}$ and the root value $m_{t+1}$.
-When searching on $s_{t}$, we set the value of the node corresponding to $s_{t+1}$ in search tree $T_{t}$ as $m_{t+1}$.
-If action $a_{t}$ is selected, the traverse process will be terminated, otherwise the traverse will proceed normally.
+4.1 An Inspiring Case Study
+As a classic algorithm in the realm of stochastic bandits, Upper Confidence Bound 1 (UCB1) [8] has been deployed to select nodes within a search tree.
+Figure 1: The connection between MCTS and bandits.
+Left shows the tree search in a bandit-view.
+When action $A$ is selected, a reward $r_{A}$ will be returned, where $r_{A}=\sum_{t=1}^{3}\gamma^{t-1}r_{t}^{a}+\gamma^{3}v_{a}$.
+For the root node, the traversal, evaluation and back-propagation occurring in the gray box can be approximated as sampling from an non-stationary distribution.
+Thus the root node can be seen as a non-stationary bandit.
+Right shows the one-armed bandit case.
+Once the reward $\mu_{A}$ is known, we can evaluate arm A using $\mu_{A}$, thereby eliminating the need to rely on subsequent tree search processes.
+As explained in Figure 1, each node in the tree can be regarded as a multi-armed bandit Notice that in the context of stochastic bandits, the reward for each arm is sampled from a stationary probability distribution.
+However, in the process of tree search, the value obtained from selecting a node depends on subsequent search process.
+The UCT [27] algorithm models this process as sampling from a non-stationary distribution which converges.
+For the sake of analytical convenience in this paper, we approximate the root node as a stochastic bandit and consider each child node as an arm with a deterministic reward distribution.
+Based on this modeling, we can analyze the Eq.
+4 employed in AlphaZero [9] and derive the following proposal:
+Proposal 1
+Denote the prior for arm $i$ as $P_{i}$.
+Without loss of generality, suppose $\mu^{*}=\mu_{1}$.
+Assume that the reward distribution of each arm $i$ is $1$-subgaussian and $P_{1}>\mu_{1}$, then the regret $R_{n}$ of using Eq.
+4 to choose arm satisfies
+$R_{n}\leq\sum_{i:\Delta_{i}>0}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(3+\frac{16}{\varepsilon^{4}P_{1}^{2}}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})$ (5)
+The assumption that $P_{1}>\mu_{1}$ is made purely for the sake of formality; in fact, Appendix A indicates that, as long as there exists a $t$ that $P_{1}\sqrt{t-1}>\mu_{1}$, a similar upper bound on regret can always be obtained.
+This elucidates that selecting actions according to Eq.
+4 can yield a sublinear growth of regret, indicating that the final visit distribution will be concentrated on the optimal arm (see analysis for Lemma A.1 in Appendix A).
+This also confirms the rationality of selecting the action with the largest number of visits.
+The inspiration of our enhanced method comes from the one-armed bandit model: if the expected reward of a particular arm is known as prior, the necessity for exploratory trials on this arm is obviated, thereby reducing the regret upper bound.
+To concretize this insight, we can extend Eq.
+4 and deduce the following proposition:
+Proposal 2
+For a bandit which satisfies the assumptions in Proposal 1, the regret upper bound can be reduced if we use the known expectation value of an arm to replace its score in Eq.
+4.
+Specifically, if the known arm is the optimal arm with value $\mu_{1}$, then the regret $R_{n}$ satisfies:
+$R_{n}\leq\sum_{i:\Delta_{i}>0}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(2+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})$ (6)
+In other case, with the known value $\mu_{l}$ of a sub-optimal arm $l$, the regret $R_{n}$ satisfies:
+$R_{n}$ $\leq\sum_{i:\Delta_{i}>0,i\neq l}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(3+\frac{16}{\varepsilon^{4}P_{1}^{2}}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})+\inf_{\varepsilon\in(0,\Delta_{l})}\Delta_{l}(1+\frac{16}{\varepsilon^{4}P_{1}^{2}})$ (7)
+Proposal 2 demonstrates that, given the known expected reward of a suboptimal arm, the asymptotic upper bound of regret can be further reduced, thereby concentrating the visit distribution more heavily on the optimal arm.
+Meanwhile, as shown in Figure 1, when the expectation $\mu_{A}$ is known, we no longer need to evaluate arm A through real-time search results.
+Instead, we can directly substitute $\mu_{A}$ for the score in Eq.
+4, thus saving the time spent on traverse and back-propagation processes.
+This inspires us to evaluate a particular child node using existing information, thereby eliminating the exploration of the corresponding sub-tree and accelerating the search process.
+See Appendix A for detailed proof and analysis of the two proposals.
+4.2 Backward-view Reanalyze
+Figure 2: An illustration about the backward-view reanalyze in both node and batch view.
+We sample $n+1$ trajectories of length $k+1$ to form a batch and conduct the search in the reverse direction of trajectories.
+From the node view, we would first search $S_{0}^{1}$ and then pass root value $m_{0}^{1}$ to $S_{0}^{0}$ to evaluate the value of a child node.
+$T_{0}^{1}$ and $T_{0}^{0}$ are the corresponding search trees.
+From the batch view, we would group all $S^{1}$s into a sub batch to search together and pass the root values to the $S^{0}$s.
+Inspired by the aforementioned motivation, we aspire to obtain the reward expectation of a child node in advance to assess it.
+However, the true expectation is always unknown.
+Consequently, we resort to using the root value obtained from MCTS as an approximate substitute.
+Specifically, for two adjacent time steps $S_{0}^{1}$ and $S_{0}^{0}$ in Figure 2, when searching for state $S_{0}^{1}$, the root node corresponds to a child node of $S_{0}^{0}$.
+Therefore, the root value obtained can be utilized to assist the search for $S_{0}^{0}$.
+However, there is a temporal contradiction.
+$S_{0}^{1}$ is the successor state of $S_{0}^{0}$, yet we need to complete the search for $S_{0}^{1}$ first.
+Fortunately, this is possible during the reanalyze process.
 Algorithm 1 shows the specific design with Python-like code.
-During reanalyze, since the trajectories was already collected in the replay buffer, we can perform the search in a reverse order.
-That is to say, we first perform tree search on the state of the last time-step in the trajectory, and then search the second to last, and so on.
-After we search on state $s_{t+1}$  [^1] To facilitate the explanation, we have omitted details such as the latent space and do not make a deliberate distinction between states and nodes.
-(the state of the $t+1$-th time-step in the trajectory), the root value $m_{t+1}$ is obtained.
-When performing search on $s_{t}$, we set the value of $s_{t+1}$ to the fixed value $m_{t+1}$.
-During traverse in the tree, we select the action $a_{root}$ for root node $s_{t}$ according to the following equation:
-$a_{root}=\arg\max_{a}I_{t}(a)$ (9)
-$I_{t}(a)=\begin{cases}UCB_{score}(s_{t},a),&a\neq a_{t}\\ r_{t}+\gamma m_{t+1},&a=a_{t}\end{cases}$ (10)
-where $a$ is an action corresponding to a child node, $a_{t}$ is the action corresponding to $s_{t+1}$, and $r_{t}$ is the reward predicted by the dynamic model.
-If an action other than $a_{t}$ is selected, the simulation just traverses following the original setting in MCTS (Schrittwieser et al., 2019).
-If action $a_{t}$ is selected, this simulation is terminated immediately.
-Since the time used to search for node $s_{t+1}$ is saved, the modified search process is faster than the original version, achieving Speedy Reanalyze.
+During reanalyze, since the trajectories was already collected in the replay buffer, we can perform tree search in a backward-view, which searches the trajectory in a reverse order.
+Figure 2 illustrates a batch containing $n+1$ trajectories, each of length $k+1$.
+We denote $S_{l}^{t}$ as the $t$-th state in the $l$-th trajectory.
+We first conduct a search for all $S^{k}$s, followed by a search for all $S^{k-1}$s, and so on.
+After we search on state $S_{l}^{t+1}$  [^1] To facilitate the explanation, we have omitted details such as the latent space and do not make a deliberate distinction between states and nodes., the root value $m_{l}^{t+1}$ is obtained.
+When engaging in search on $S_{l}^{t}$, we assign the value of $S_{l}^{t+1}$ to the fixed value $m_{l}^{t+1}$.
+During traverse in the tree, we select the action $a_{root}$ for root node $S_{l}^{t}$ with the following equation:
+$a_{root}=\arg\max_{a}I_{l}^{t}(a)$ (8)
+$I_{t}(a)=\begin{cases}UCB_{score}(S^{t}_{l},a),&a\neq a_{l}^{t}\\ r_{l}^{t}+\gamma m_{l}^{t+1},&a=a_{l}^{t}\end{cases}$ (9)
+where $a$ refers to the action associated with a child node, $a_{l}^{t}$ is the action corresponding to $S_{l}^{t+1}$, and $r_{l}^{t}$ signifies the reward predicted by the dynamic model.
+If an action distinct from $a_{l}^{t}$ is selected, the simulation continues its traversal in accordance with the original setting [10].
+If action $a_{l}^{t}$ is selected, this simulation is terminated immediately.
+Since the time used to search for node $S^{t+1}_{l}$ is saved, this enhanced search process is faster than the original version.
+Notice that the analysis in Proposal 2 assumes that we know the expectation value.
+However, in our design, we use the root value as an approximate estimation.
+Still based on the stochastic bandit assumption, the following theorem analyzes the impact of the error between these two values:
+Theorem 1
+For a bandit which satisfies assumptions in Proposal 1 and a sample result for an arm, the regret upper bound can be reduced if we use the way of Equation 8 to choose arm.
+Specifically, if the sampled arm is the optimal arm with $n$ times sample mean $\hat{\mu}_{1n}$, then the regret $R_{n}$ satisfies:
+$R_{n}\leq\sum_{i:\Delta_{i}>0}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(2+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+n\exp(-\frac{n\varepsilon^{2}}{2}))$ (10)
+In other case, with the $n$ times sample mean $\hat{\mu}_{ln}$ of a sub-optimal arm $l$, the regret $R_{n}$ satisfies:
+$R_{n}$ $\leq\sum_{i:\Delta_{i}>0,i\neq l}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(3+\frac{16}{\varepsilon^{4}P_{1}^{2}}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})$ (11)
+$+$ $\inf_{\varepsilon\in(0,\Delta_{l})}\Delta_{l}(1+\frac{16}{\varepsilon^{4}P_{1}^{2}}+n\exp(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{2}))$ (12)
+This theorem demonstrates that using the sampled mean as a substitute for the expectation value can also reduce the asymptotic upper bound of regret and maintain sub linear growth rate.
+This can be understood as follows: although there is an error between the sampled mean and the expectation value, since we use the average of $n$ samples (as the root value comes from a tree search which has $n$ simulations), this sampled mean will increasingly approximate the expectation value as $n$ grows, thereby preventing a linear increase.
+Refer to Appendix A for the proof.
 Algorithm 1 Python-like code for information reuse
@@
 backpropagate()
-4.2.2 Theoretical Analysis
-AlphaZero employs a variant of the UCB algorithm during MCTS search.
-Subsequently, MuZero proposes further improvements to this variant formula.
-Practical evidence has demonstrated the effectiveness of both these node selection methods; however, they have not been theoretically analyzed in their original paper.
-In earlier studies, UCT analyzed the convergence properties of the UCB1 algorithm in a tree structure, albeit without introducing a prior policy.
-PUCB (Rosin, 2011) incorporates relevant analysis of a prior policy, but it uses the prior in a penalty function-like approach which is different from in the MCTS-based algorithm family.
-In recent work, Gumbel MuZero (Danihelka et al., 2022) approximated the tree search model as a stochastic bandit, and on this basis, proved the monotonic improvement of their designs.
-In this section, we will draw upon such a simplification and demonstrate within the context of a stochastic bandit model why the node selection formula and policy updating methods used in the MCTS-based algorithm family are effective, and how information reuse can influence the policy target.
-Definition 4.1.
-After $n$ simulations with a stochastic bandit $\nu$ with $K$ arms, let $T_{i}$ denote the number of visits to arm $i$, then the policy generated based on the visit distribution is a random vector $P=[\frac{T_{1}}{n},...,\frac{T_{K}}{n}]$ The value $V_{\nu}(P)$ of policy $P$ on $\nu$ is defined as:
-$V_{\nu}(P)=\sum_{i}\frac{T_{i}\mu_{i}(\nu)}{n}$ (11)
-In cases where it does not cause any ambiguity, we will omit the mark $\nu$ in the following text. The next theorem explains the rationale behind using distribution as a target policy:
-Theorem 4.2.
-If an algorithm $\pi$ exhibits sub-linear growth in regret $R_{n}$ with respect to a bandit $\nu$, then the expectation $E[V(P)]$ converges to $\mu^{*}$ as $n$ tends to infinity.
-$E[V(P)]$ $=\sum_{i}\frac{E[T_{i}]\mu_{i}}{n}$ (12)
-$=\sum_{i}\frac{E[T_{i}](\mu^{*}-\Delta_{i})}{n}$ (13)
-$=\mu^{*}-\frac{\sum_{i}E[T_{i}]\Delta_{i}}{n}$ (14)
-where $\Delta_{i}=\mu^{*}-\mu_{i}$, and $\sum_{i}E[T_{i}]\Delta_{i}$ is the regret $R_{n}$ of the algorithm (see Appendix A).
-Since $R_{n}$ growths in a sub-linear rate, so we can get
-$\lim_{n\to\infty}E[V(P)]=\mu^{*}-\lim_{n\to\infty}\frac{R_{n}}{n}=\mu^{*}$ (15)
-The following theorem provides an upper bound for the growth rate of $R_{n}$ for choosing arm in AlphaZero’s way , thereby demonstrating that the regret is sub-linearly increasing.
-Theorem 4.3.
-Denote the prior for arm $i$ as $P_{i}$.
-Without loss of generality, suppose $\mu^{*}=\mu_{1}$.
-Assume that the reward distribution of each arm $i$ is $1$-subgaussian and $P_{1}>\mu_{1}$, then the regret $R_{n}$ of AlphaZero satisfies (Appendix A)
-$R_{n}\leq\sum_{i:\Delta_{i}>0}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(3+\frac{16}{\varepsilon^{4}P_{1}^{2}}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})$ (16)
-The upper bound given by Theorem 4.3 is $O(\sqrt{n})$, which, while not reaching the level of $O(\log{n})$, is still sub-linear.
-Inequality (16) also indicates that a better policy which has higher probability to choose the best arm can lower the upper bound.
-Combining 4.2 and 4.3, we can conclude that with the increase in the number of simulations, the visit distribution will increasingly concentrate on the optimal arm, and the value corresponding to the policy will also converge to the value of the optimal arm.
-If the assumption that $P_{1}>\mu_{1}$ doesn’t hold, an upper bound with $O(\sqrt{n})$ rate can still be obtained by similar proof as long as there exists a $t$ that $P_{1}\sqrt{t-1}>\mu_{1}$, which indicates unlike some algorithms that are sensitive to prior policies, MCTS has strong robustness to priors.
-As long as the number of simulations is sufficient, it can still achieve a good target policy.
-When reusing information in the search process, we can obtain a reference value for an arm.
-The existence of this reference value can eliminate the need for exploration of known arms, thereby further reducing the algorithm’s regret.
-The following theorem proves this point.
-Theorem 4.4.
-For a bandit which satisfies the assumptions in 4.3 and an already known arm value, the regret upper bound can be reduced if we use the way of 4.2.1 to choose arm.
-Specifically, if the known arm is the optimal arm with value $\mu_{1}$, then the regret $R_{n}$ satisfies:
-$R_{n}\leq\sum_{i:\Delta_{i}>0}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(2+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})$ (17)
-In other case, with the known value $\mu_{l}$ of a sub-optimal arm $l$, the regret $R_{n}$ satisfies:
-$R_{n}$ $\leq\sum_{i:\Delta_{i}>0,i\neq l}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(3+\frac{16}{\varepsilon^{4}P_{1}^{2}}$ (18)
-$+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})$ (19)
-$+\inf_{\varepsilon\in(0,\Delta_{l})}\Delta_{l}(1+\frac{16}{\varepsilon^{4}P_{1}^{2}})$ (20)
-The complete proof can be found in Appendix A.
+4.3 The ReZero Framework
+The technique introduced in Section 4.2 will bring a new problem in practice.
+The experiments in the Appendix C.3 demonstrate that batching MCTS allows for parallel model inference and data processing, thereby accelerating the average search speed.
+However, as shown in the Figure 2, to conduct the backward-view reanalyze, we need to divide the batch into $\frac{1}{k+1}$ of its original size.
+This diminishes the benefits of parallelized search and, on the contrary, makes the algorithm slower.
+Figure 3: Execution workflow and runtime cycle graph about MuZero and ReZero in both single and multiple worker cases.
+The number inside the modules represent the number of iterations, and the number under the modules represent the time required for module execution.
+The model is updated $n$ iterations between two collections.
+MuZero reanalyzes the mini-batch before each model update.
+ReZero reanalyzes the entire buffer after certain iterations($\frac{n}{3}$ for example), which not only reduces the total number of MCTS calls, but also takes advantage of the processing speed of large batches.
+We propose a new pipeline that is more compatible with the method introduced in Section 4.2.
+In particular, during the collect phase, we have transitioned from using MCTS to select actions to directly sampling actions based on the policy network’s output.
+This shift can be interpreted as an alternative method to augment exploration in the collect phase, akin to the noise injection at the root node in MCTS-based algorithms.
+Figure 10 in Appendix C.3 indicates that this modification does not significantly compromise performance during evaluation.
+For the reanalyze process, we introduce the periodical entire-buffer reanalyze mechanism.
+As shown in Figure 3, we reanalyze the whole buffer after a fixed number of training iterations.
+For each training iteration, we don’t need to run MCTS to reanalyze the mini-batch, only need to sample the mini-batch and execute the gradient descent.
+Overall, this design offers two significant advantages: 1 The entire buffer reanalyze is akin to the fixed target net mechanism in DQN [28], maintaining a constant policy target for a certain number of training iterations.
+Reducing the frequency of policy target updates correspondingly decreases the number of MCTS calls.
+Concurrently, this does not result in a decrease in performance.
+2 Since we no longer invoke MCTS during the collect phase, all MCTS calls are concentrated in the reanalyze process.
+And during the entire-buffer reanalyze, we are no longer constrained by the size of the mini-batch, allowing us to freely adjust the batch size to leverage the advantages of large batches.
+Figure 11 shows both excessively large and excessively small batch sizes can lead to a decrease in search speed.
+We choose the batch size of $2000$ according to the experiment in Appendix C.3.
+Experiments show that our algorithm maintains high sample efficiency and greatly save the running time of the algorithm. Our pipeline also has the following potential improvement directions:
+  • When directly using policy for data collection, action selection is no longer bound by tree search.
+Thus, previous vectorized environments like [15] can be seamlessly integrated.
+Besides, this design makes MCTS-based algorithms compatible to existing RL exploration methods like [29].
+  • Our method no longer needs to reanalyze the mini-batch for each iteration, thus decoupling the process of reanalyze and training.
+This provides greater scope for parallelization.
+In the case of multiple workers, we can design efficient parallelization paradigms as shown in Figure 3.
+  • We can use a more reasonable way, such as weighted sampling to preferentially reanalyze a part of the samples in the buffer, instead of simply reanalyzing all samples in the entire buffer.
+This is helpful to further reduce the computational overhead.
+5 Experiment
 Figure 4: Time-efficiency of ReZero-M vs.
@@
 Mean of 5 runs, shaded areas are 95% confidence intervals.
+Efficiency in RL usually refers to two aspects: sample efficiency, the agent’s ability to learn effectively from a limited number of environmental interactions, gauged by the samples needed to reach a successful policy using equal compute resources; time efficiency, which is how swiftly an algorithm learns to make optimal decisions, indicated by the wall-clock time taken to achieve a successful policy.
+Our primary goal is to improve time efficiency without compromising sample efficiency.
+To empirically validate the efficiency of our introduced two boosting techniques for MCTS-based algorithms, we have incorporated the ReZero with two prominent algorithms detailed in [30]: MuZero with a Self-Supervised Learning loss (SSL) and EfficientZero, yielding the enhanced variants ReZero-M and ReZero-E respectively.
+For brevity, when the context is unambiguous, we simply refer to them as ReZero, omitting the symbol representing the baseline algorithm.
+In order to validate the applicability of our method across multiple types of decision-making environments, we opted for six representative Atari environments characterized by classic image-input observation and discrete action spaces, in addition to the strategic board games Connect4 and Gomoku with special state spaces.
+For a baseline comparison, we employed the original implementations of MuZero and EfficientZero as delineated in the LightZero benchmark [30].
+The full implementation details available in Appendix B.
+Besides, we emphasize that to ensure an fair comparison of wall-clock time, all experimental trials were executed on a fixed single worker hardware settings.
+In the next subsections, we sequentially explore and try to answer the following three questions:
+  • How much can ReZero-M/ReZero-E improve time efficiency compared to MuZero/EfficientZero, while maintaining equivalent levels of high sample efficiency?
+  • What is the effect and hyper-parameter sensitivity of the Entire-buffer Reanalyze technique?
+  • How much search budgets can be saved in practice by the Backward-view Reanalyze technique?
 Atari Board Games
@@
 ReZero-M (ours) $\mathbf{1}\scriptstyle{\pm 0.6}$ $\mathbf{7.7}\scriptstyle{\pm 0.3}$ $\mathbf{3.8}\scriptstyle{\pm 0.2}$ $\mathbf{5.1\scriptstyle{\pm 0.3}}$ $\mathbf{6.1}\scriptstyle{\pm 0.5}$ $\mathbf{6.9}\scriptstyle{\pm 0.5}$ $\mathbf{10}\scriptstyle{\pm 0.8}$ $\mathbf{9\scriptstyle{\pm 0.8}}$
-MuZero (Schrittwieser et al., 2019) $4.1\scriptstyle{\pm 0.2}$ $12.4\scriptstyle{\pm 0.8}$ $12.0\scriptstyle{\pm 0.9}$ $10.2\scriptstyle{\pm 0.5}$ $11.75\scriptstyle{\pm 0.9}$ $12.4\scriptstyle{\pm 0.9}$ $11\scriptstyle{\pm 0.9}$ $23\scriptstyle{\pm 2}$
+MuZero [10] $4.1\scriptstyle{\pm 0.2}$ $12.4\scriptstyle{\pm 0.8}$ $12.0\scriptstyle{\pm 0.9}$ $10.2\scriptstyle{\pm 0.5}$ $11.75\scriptstyle{\pm 0.9}$ $12.4\scriptstyle{\pm 0.9}$ $11\scriptstyle{\pm 0.9}$ $23\scriptstyle{\pm 2}$
 Table 1: Average wall-time of ReZero-M vs.
@@
 Mean and standard deviation over 5 runs.
-From the proof, it can be seen that if the real value $\mu_{1}$ of the optimal arm is known in advance, the regret caused by the underestimation of the optimal arm can be eliminated.
-However, the part of the regret that is reduced is $O(1)$.
-If the real value $\mu_{i}$ of a sub-optimal arm $l$ is known in advance, the regret caused by continuous exploration of this sub-optimal arm can be eliminated.
-The part of the reduced regret is $O(\sqrt{n})$.
-Intuitively, it can be understood this way: If the value of a sub-optimal arm $l$ is known, then after sufficient exploration of the optimal arm, the upper confidence bound will converge to be higher than $\mu_{l}$ with probability $1$, thus making the regret caused by arm $l$ no longer increases with $n$.
-While if the value of the optimal arm is known, then as the number of simulations $n$ increases, there will always be a moment when the upper confidence bound of the sub-optimal arm $i$ is greater than $\mu_{1}$, so arm $i$ will still be repeatedly selected, leading to the regret caused by arm $i$ increases with $n$.
-However, in this case, since the optimal arm is chosen the most time as $n$ increases, we can save more time by early stopping the search.
-5 Experiment
-Efficiency in RL usually refers to two aspects: sample efficiency, the agent’s ability to learn effectively from a limited number of environmental interactions, gauged by the samples needed to reach a successful policy using equal compute resources; time efficiency, which is how swiftly an algorithm learns to make optimal decisions, indicated by the wall-clock time taken to achieve a successful policy.
-Our primary goal is to improve time efficiency without compromising sample efficiency.
-To empirically validate the efficiency of the two boosting techniques introduced in Section 3.2 for MCTS-based algorithms, we have incorporated the ReZero with two prominent algorithms detailed in (Niu et al., 2023): MuZero with a Self-Supervised Learning loss (SSL) and EfficientZero, yielding the enhanced variants ReZero-M and ReZero-E respectively.
-For brevity, when the context is unambiguous, we simply refer to them as ReZero, omitting the symbol representing the baseline algorithm.
-In order to validate the applicability of our method across multiple types of decision-making environments, we opted for six representative Atari environments characterized by classic image-input observation and discrete action spaces, in addition to the strategic board games Connect4 and Gomoku with special state spaces.
-For a baseline comparison, we employed the original implementations of MuZero and EfficientZero as delineated in the LightZero benchmark (Niu et al., 2023).
-The full implementation details available in Appendix B.
-Besides, we emphasize that to ensure an fair comparison of wall-clock time, all experimental trials were executed on a fixed single worker hardware settings.
-In the next subsections, we sequentially explore and try to answer the following three questions:
-  • How much can ReZero-M/ReZero-E improve time efficiency compared to MuZero/EfficientZero, while maintaining equivalent levels of high sample efficiency?
-  • What is the concrete effect and hyper-parameter sensitivity of the Just-in-Time Reanalyze technique in ReZero?
-  • How much search budgets and time costs can be saved in practice by the Temporal Information Reuse technique?
 5.1 Time Efficiency
@@
 Key parameters are aligned with those from original papers.
-Specifically, we set the replay ratio (the ratio between environment steps and training steps) (D’Oro et al., 2022; Schwarzer et al., 2023) to 0.25, and the reanalyze ratio (the ratio between targets computed from interactions with the environment and by reanalysing existing data) (Schrittwieser et al., 2021) is set to 1.
+Specifically, we set the replay ratio (the ratio between environment steps and training steps) [31, 32] to 0.25, and the reanalyze ratio (the ratio between targets computed from interactions with the environment and by reanalysing existing data) [13] is set to 1.
 For detailed hyper-paramater configurations, please refer to the Appendix B.
@@
 It is apparent that ReZero-M achieves a significant improvement in time efficiency, reaching a near-optimal policy in substantially less time for the games Pong, Breakout and Seaquest, and obtained a better policy than MuZero on the remaining tasks.
-Additional results about ReZero-E and EfficientZero are detailed in Appendix C.2, specifically in Figure 6 and 7.
+Additional results about ReZero-E and EfficientZero are detailed in Appendix C.2, specifically in Figure 7 and 8.
 For a more comprehensive understanding of our findings, we provide an in-depth comparison of the wall-clock training time up to 100k environment steps for ReZero-M against MuZero in Table 1.
 The data reveals that, on most games, ReZero require between two to four times less wall-clock time per 100k steps compared to the baseline methods.
-Figure 4: The effect of Reanalyze Frequency in ReZero-M on MsPacman.
-The horizontal axis in left represents Wall-clock Time (hours) while right side uses Env Steps.
-The vertical axis indicates Return.
-Proper reanalyze frequency can improve efficiency while obtaining the comparable Return with MuZero (reanalyze ratio=1).
+Figure 5: The effect of Reanalyze Frequency in ReZero-M on the MsPacman game.
+The proper reanalyze frequency can improve time and sample efficiency while obtaining the comparable return with MuZero (reanalyze ratio=1).
 5.2 Effect of Reanalyze Frequency
 In next two sub-sectons, we will delve into details to explain how ReZero works.
-Firstly, we will analyze how the Just-in-Time Reanalyze technique manages to simplify the original iterative mini-batch reanalyze scheme to the periodical updated version while maintaining high sample efficiency.
+Firstly, we will analyze how the entire-buffer reanalyze technique manages to simplify the original iterative mini-batch reanalyze scheme to the periodical updated version while maintaining high sample efficiency.
 Here we adjust the periodical reanalyze frequency(reanalyze the buffer 1 time every $n$ training iterations) in ReZero to a lower or higher value on the MsPacman environment, specifically reanalyze frequency=$\{2,1,1/3,0\}$, where $1$ denotes the setting $n=1000$.
@@
 Figure 5 shows entire training curves in terms of Wall-time or Env Steps and validates that appropriate reanalyze frequency can save the time overhead without causing any obvious performance loss.
-5.3 Effect of Information Reuse
-Indicators Avg. time (ms) tree search (num calls) dynamics model (num calls) data process (num calls)
+5.3 Effect of Backward-view Reanalyze
+Indicators Avg.time (ms) tree search (num calls) dynamics (num calls) data process (num calls)
 ReZero-M $\textbf{0.69}\pm\textbf{0.02}$ 6089 122 277
 MuZero $1.08\pm 0.09$ $13284$ $256$ $455$
-Table 2: Comparisons about detailed indicators between ReZero-M and MuZero inside the tree search.
+Table 2: Comparisons about detailed time indicators between MuZero and ReZero-M inside the tree search.
+To further validate and understand the advantages and significance of the backward-view reanalyze technique proposed in ReZero, we meticulously document a suite of statistical indicators of the tree search process in Table 2.
 The number of function calls is the cumulative value of 100 training iterations on Pong.
@@
 time is the average time of a MCTS across all calls.
-To further validate and understand the advantages and significance of the Information Reuse technique proposed in ReZero, we meticulously document a suite of statistical indicators of the tree search process in Table 2.
-Comparative analysis between ReZero-M and MuZero reveals that the Information Reuse technique reduces the invocation frequency of the dynamics model, the search tree, and other operations like data process transformations.
+Comparative analysis between ReZero-M and MuZero reveals that the backward-view reanalyze technique reduces the invocation frequency of the dynamics model, the search tree, and other operations like data process transformations.
 Consequently, this advanced technique in leveraging subsequent time step data contributes to save the tree search time in various MCTS-based algorithms, further leading to overall wall-clock time gains.
-Besides, Information Reuse also incrementally reallocates computations (tree search simulations) towards previously non-visited nodes rather than repeatedly accessing extant nodes, which maybe yield some performance gains in some environments (e.g.
-Breakout in Figure 3).
+Additionally, Theorem 1 demonstrates that the backward-view reanalyze can reduce the regret upper bound, which indicates a better search result.
+Figure 9 in Appendix C.3 shows a comparison of sample efficiency between using backward-view reanalyze and origin reanalyze process in MsPacman.
+The experimental results reveal that our method not only enhances the speed of individual searches but also improves sample efficiency.
+This aligns with the theoretical analysis.
 6 Conclusion and Limitation
@@
 However, our current experiments are mainly conducted on the single worker setting, there exists considerable optimization scope to apply our approach into distributed RL training, and our design harbors the potential of better parallel acceleration and more stable convergence in large-scale training tasks.
-Also, the combination between ReZero and frameworks akin to AlphaZero, or its integration with some recent offline datasets such as RT-X (Padalkar et al., 2023), constitutes a fertile avenue for future research.
+Also, the combination between ReZero and frameworks akin to AlphaZero, or its integration with some recent offline datasets such as RT-X [33], constitutes a fertile avenue for future research.
 These explorations could broadening the application horizons of MCTS-based algorithms.
-7 Impact Statements
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
-8 Acknowledgements
-Thanks to each of the authors for their efforts in the work of ReZero.
-We spent several memorable mid-nights and early mornings together in the lab.
-Thanks to everyone at OpenDILab and their pioneering work, especially the LightZero benchmark (https://github.com/opendilab/LightZero) for providing the base code for ReZero.
-Finally, we would like to thank Shanghai AI Laboratory and professor Jing Yang of Xi’an Jiaotong University for providing a platform for this work.
 Appendix A Proof Materials
-To verify the theoretical convergence guarantee of ReZero, we have provided some parts of detailed theoretical proofs in the main paper.
 In this section, we will present the complete supplementary proofs.
@@
 The proof can be completed by combining the two equations and using the definition of $T_{i}$. ∎
+This lemma demonstrates that for a certain bandit, if the regret grows sub linearly, then the expected times each sub optimal arm is chosen also grows sub linearly.
+Consequently, as $n$ increases, the proportion of times the sub optimal arm is selected converges to 00.
+In other words, the visit count distribution concentrates on the optimal arm as $n$ grows.
 Lemma A.2.
@@
 $=1+u+\frac{2}{\varepsilon^{2}}=1+\frac{2\sqrt{a}}{\varepsilon}+\frac{2}{\varepsilon^{2}}$ (27)
-Proof for Theorem 4.3:
+Proof for Proposal 1:
 Proof.
@@
 $+\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}$ (30)
-for 37, we have
+for 29, we have
 $E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]$ (31)
@@
 $\leq 1+\frac{16}{\varepsilon^{4}P_{1}^{2}}$ (37)
-Notes: In 40, since we assume $P_{1}\geq\mu_{1}$, the probability of the term $s=0$ would be 00.
+Notes: In 32, since we assume $P_{1}\geq\mu_{1}$, the probability of the term $s=0$ would be 00.
 Thus, we can discard it.
@@
 This only changes the constant term and won’t affect the growth rate of regret.
-From 42 to 43, we just need to expand the quadratic term and do some simple inequality scaling.
-From 43 to 44, we need to notice that $\sum_{s=1}^{t-1}\exp(-\frac{s\varepsilon^{2}}{2})$ is a geometric sequence and scale it to $\frac{2}{\varepsilon^{2}}$.
-From 44 to 45, we use the inequality $\sum_{t=2}^{n}\frac{1}{e^{a\sqrt{t-1}}}\leq\int_{1}^{\infty}\frac{1}{e^{a\sqrt{t-1}}}$.
-for 38, we have
+From 34 to 35, we just need to expand the quadratic term and do some simple inequality scaling.
+From 35 to 36, we need to notice that $\sum_{s=1}^{t-1}\exp(-\frac{s\varepsilon^{2}}{2})$ is a geometric sequence and scale it to $\frac{2}{\varepsilon^{2}}$.
+From 36 to 37, we use the inequality $\sum_{t=2}^{n}\frac{1}{e^{a\sqrt{t-1}}}\leq\int_{1}^{\infty}\frac{1}{e^{a\sqrt{t-1}}}$.
+for 30, we have
 $E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (38)
@@
 combined with Lemma A.1, the proof is completed. ∎
-Proof for Theorem 4.4:
+[…]
+Proof for Proposal 2:
 Proof.
-The proof can be obtained by making slight modifications to the proof of 4.3. In case that $\mu_{1}$ is already known,
-$T_{i}(n)=\sum_{t=1}^{n}\mathbb{I}\{A_{t}=i\}$ (52)
-$\leq\sum_{t=1}^{n}\mathbb{I}\{\mu_{1}\leq\mu_{1}-\varepsilon\}$ (53)
-$+\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}$ (54)
-$=\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}$ (55)
+The proof can be obtained by making slight modifications to the proof of Proposal 1. In case that $\mu_{1}$ is already known,
+$E[T_{i}(n)]=E\sum_{t=1}^{n}\mathbb{I}\{A_{t}=i\}]$ (44)
+$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\mu_{1}\leq\mu_{1}-\varepsilon\}]$ (45)
+$+E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (46)
+$=E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (47)
 $\leq 2+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}$ (48)
 In case that a sub-optimal value $\mu_{l}$ is known,
-$T_{l}(n)=\sum_{t=1}^{n}\mathbb{I}\{A_{t}=l\}$ (57)
-$\leq\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}$ (58)
-$+\sum_{t=1}^{n}\mathbb{I}\{\mu_{l}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}$ (59)
-$=\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}$ (60)
+$E[T_{l}(n)]=E[\sum_{t=1}^{n}\mathbb{I}\{A_{t}=l\}]$ (49)
+$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]$ (50)
+$+E[\sum_{t=1}^{n}\mathbb{I}\{\mu_{l}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (51)
+$=E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]$ (52)
 $\leq 1+\frac{16}{\varepsilon^{4}P_{1}^{2}}$ (53)
-and the bound of $T_{i}(n)$ for $i\neq l$ keeps unchanged.
+and the bound of $T_{i}(n)$ for $i\neq l$ keeps unchanged. In both cases, combined with Lemma A.1, the proof is completed. ∎
+From the proof, it can be seen that if the real value $\mu_{1}$ of the optimal arm is known in advance, the regret caused by the underestimation of the optimal arm can be eliminated.
+However, the part of the regret that is reduced is $O(1)$.
+If the real value $\mu_{i}$ of a sub-optimal arm $l$ is known in advance, the regret caused by continuous exploration of this sub-optimal arm can be eliminated.
+The part of the reduced regret is $O(\sqrt{n})$.
+Intuitively, it can be understood this way: If the value of a sub-optimal arm $l$ is known, then after sufficient exploration of the optimal arm, the upper confidence bound will converge to be higher than $\mu_{l}$ with probability $1$, thus making the regret caused by arm $l$ no longer increases with $n$.
+While if the value of the optimal arm is known, then as the number of simulations $n$ increases, there will always be a moment when the upper confidence bound of the sub-optimal arm $i$ is greater than $\mu_{1}$, so arm $i$ will still be repeatedly selected, leading to the regret caused by arm $i$ increases with $n$.
+However, in this case, since the optimal arm is chosen the most time as $n$ increases, we can save more time by early stopping the search.
+Proof for Theorem 1:
+Proof.
+The proof can be obtained by making slight modifications to the proof of Proposal 2.
+In case we have $\hat{\mu}_{1n}=\frac{1}{n}\sum_{t=1}^{n}\hat{\mu}_{t}$ where $\hat{\mu}_{t}$ is sampled from the same reward distribution of arm $1$,
+$E[T_{i}(n)]=E[\sum_{t=1}^{n}\mathbb{I}\{A_{t}=i\}]$ (54)
+$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1n}\leq\mu_{1}-\varepsilon\}]$ (55)
+$+E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (56)
+$\leq n\exp{(-\frac{n\varepsilon^{2}}{2})}+E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (57)
+$\leq 2+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+n\exp{(-\frac{n\varepsilon^{2}}{2})}$ (58)
+In case we have $\hat{\mu}_{ln}=\frac{1}{n}\sum_{t=1}^{n}\hat{\mu}_{t}$ where $\hat{\mu}_{t}$ is sampled from the same reward distribution of arm $l$,
+$E[T_{l}(n)]=E[\sum_{t=1}^{n}\mathbb{I}\{A_{t}=l\}]$ (59)
+$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]$ (60)
+$+E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{ln}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (61)
+$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{2})}$ (62)
+$\leq 1+\frac{16}{\varepsilon^{4}P_{1}^{2}}+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{2})}$ (63)
+and the bound of $E[T_{i}(n)]$ for $i\neq l$ keeps unchanged.
 In both cases, combined with Lemma A.1, the proof is completed. ∎
@@
 B.2 Algorithm Implementation Details
-Our algorithm’s implementation is based on the open-source code of LightZero (Niu et al., 2023).
+Our algorithm’s implementation is based on the open-source code of LightZero [30].
 Given that our proposed theoretical improvements are applicable to any MCTS-based RL method, we have chosen MuZero and EfficientZero as case studies to investigate the practical improvements in time efficiency achieved by integrating the ReZero boosting techniques: just-in-time reanalyze and speedy reanalyze (temporal information reuse).
 To ensure an equitable comparison of wall-clock time, all experimental trials were executed on a fixed single worker hardware setting consisting of a single NVIDIA A100 GPU with 30 CPU cores and 120 GiB memory.
-Besides, we emphasize that to ensure a fair comparison of time efficiency and sample efficiency, the model architecture and hyper-parameters used in the experiments of Section 5 are essentially consistent with the settings in LightZero (Niu et al., 2023).
+Besides, we emphasize that to ensure a fair comparison of time efficiency and sample efficiency, the model architecture and hyper-parameters used in the experiments of Section 5 are essentially consistent with the settings in LightZero [30].
 For specific hyper-parameters of ReZero-M and MuZero on Atari, please refer to the Table 3.
 The main different hyper-parameters for the ReZero-M algorithm in the Connect4 and Gomoku environment are set out in Tables 4.
-In addition to employing an LSTM network with a hidden state dimension of 512 to predict the value prefix (Ye et al., 2021), all hyperparameters of ReZero-E are essentially identical to those of ReZero-M in Table 3.
-Wall-time statistics Note that all our current tests are conducted in the single-worker case. Therefore, the wall-time reported in Table 1 and Appendix 5 for reaching 100k env steps includes:
-  • collect time: The total time spent by an agent interacting with the environment to gather experience data.(Weng et al., 2022) can be integrated to speed up.
-Besides, this design also makes MCTS-based algorithms compatible to existing RL exploration methods like (Burda et al., 2018).
-  • reanalyze time: The time used to reanalyze collected data with the current policy or value function for more accurate learning targets (Schrittwieser et al., 2021).
+In addition to employing an LSTM network with a hidden state dimension of 512 to predict the value prefix [21], all hyperparameters of ReZero-E are essentially identical to those of ReZero-M in Table 3.
+Wall-time statistics Note that all our current tests are conducted in the single-worker case. Therefore, the wall-time reported in Table 1 and Table 5 for reaching 100k env steps includes:
+  • collect time: The total time spent by an agent interacting with the environment to gather experience data.[15] can be integrated to speed up.
+Besides, this design also makes MCTS-based algorithms compatible to existing RL exploration methods like [34].
+  • reanalyze time: The time used to reanalyze collected data with the current policy or value function for more accurate learning targets [13].
   • train time: The duration for performing updates to the agent’s policy, value functions and model based on collected data.
@@
 This approach is versatile, enabling seamless integration with single-worker serial execution environments as well as multi-worker asynchronous frameworks.
-The exploration of ReZero’s extensions and its evaluations in a multi-worker (Mei et al., 2023) paradigm are earmarked for future investigation.
+The exploration of ReZero’s extensions and its evaluations in a multi-worker [16] paradigm are earmarked for future investigation.
 Board games settings Given that our primary objective is to test the proposed techniques for improvements in time efficiency, we consider a simplified version of single-player mode in all the board games.
@@
 Hyperparameter Value
-Replay Ratio (Schwarzer et al., 2023) 0.25
-Reanalyze ratio (Schrittwieser et al., 2021) 1
+Replay Ratio [32] 0.25
+Reanalyze ratio [13] 1
 Num of frames stacked 4
 Num of frames skip 4
-Reward clipping (Mnih et al., 2013) True
-Optimizer type Adam (Kingma & Ba, 2015)
+Reward clipping [28] True
+Optimizer type Adam [35]
 Learning rate $3\times 10^{-3}$
@@
 Weight of policy entropy loss 0
-Weight of SSL (self-supervised learning) loss (Ye et al., 2021) 2
+Weight of SSL (self-supervised learning) loss [21] 2
 Batch size 256
@@
 Discrete action encoding type One Hot
-Normalization type Batch Normalization (Ioffe & Szegedy, 2015)
-Priority exponent coefficient (Schaul et al., 2015) 0.6
+Normalization type Batch Normalization [36]
+Priority exponent coefficient [37] 0.6
 Priority correction coefficient 0.4
@@
 Categorical distribution in value and reward modeling True
-The scale of supports used in categorical distribution (Pohlen et al., 2018) 300
+The scale of supports used in categorical distribution [38] 300
 Table 3: Key hyperparameters of ReZero-M on Atari environments.
 Hyperparameter Value
-Replay Ratio (Schwarzer et al., 2023) 0.25
-Reanalyze ratio (Schrittwieser et al., 2021) 1
+Replay Ratio [32] 0.25
+Reanalyze ratio [13] 1
 Board size 6x7; 6x6
@@
 Figure 8 indicates that ReZero-E matches EfficientZero’s sample efficiency across most tasks.
-Additionally, Table 5 details training times to 100k environment steps, revealing that ReZero-E is 0.5 to 2 times faster than baseline methods on most games.
+Additionally, Table 5 details training times to 100k environment steps, revealing that ReZero-E is significantly faster than baseline methods on most games.
 Figure 7: Time-efficiency of ReZero-E vs.
@@
 ReZero-E (ours) $\mathbf{2.3}\scriptstyle{\pm 1.4}$ $\mathbf{3}\scriptstyle{\pm 0.3}$ $\mathbf{3.1}\scriptstyle{\pm 0.1}$ $\mathbf{3.6}\scriptstyle{\pm 0.2}$
-EfficientZero (Ye et al., 2021) $10\scriptstyle{\pm 0.2}$ $12\scriptstyle{\pm 1.3}$ $15\scriptstyle{\pm 2.3}$ $15\scriptstyle{\pm 0.7}$
+EfficientZero [21] $10\scriptstyle{\pm 0.2}$ $12\scriptstyle{\pm 1.3}$ $15\scriptstyle{\pm 2.3}$ $15\scriptstyle{\pm 0.7}$
 Table 5: Average wall-time of ReZero-E vs.
@@
 Mean and standard deviation over 5 runs.
-ReZero-E used $\mathbf{0.5-2\times}$ less wall-time per 100k steps than baseline methods on most games.
+C.3 Ablations
+Figure 9: The results of the ablation study comparing the use of backward-view reanalyze versus its absence. Mean of 3 runs; shaded areas are 95% confidence intervals.
+Figure 10: The comparison of the outcomes of sampling actions based on the policy network’s output against selecting actions using MCTS. Mean of 3 runs; shaded areas are 95% confidence intervals.
+Figure 11: The depiction of the variation in the average search speed of MCTS as the batch size increases. Mean of 3 runs; shaded areas are 95% confidence intervals.
+This section presents the results of three supplementary ablation experiments.
+Figure 9 illustrates the impact of using backward-view reanalyze within the ReZero framework on sample efficiency.
+The results indicate that backward-view reanalyze, by introducing root value as auxiliary information, achieves higher sample efficiency, aligning with the theoretical analysis regarding regret upper bound.
+Figure 10 demonstrates the effects of different action selection methods during the collect phase.
+The findings reveal that sampling actions directly from the distribution output by the policy network does not significantly degrade the experimental results compared to using MCTS for action selection.
+Figure 11 depicts the relationship between the average MCTS search duration and the batch size.
+We set a baseline batch size of 256 and experimented with search sizes ranging from 1 to 20 times the baseline, calculating the average time required to search 256 samples by dividing the total search time by the multiplier.
+The results suggest that larger batch sizes can better leverage the advantages of parallelized model inference and data processing.
+However, when the batch size becomes excessively large, constrained by the limits of hardware resources (memory, CPU), the search speed cannot increase further and may even slightly decrease.
+We ultimately set the batch size to 2000, which yields the fastest average search speed on our device.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The claims made matches theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: Section 6 discusses the limitations of the paper.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory Assumptions and Proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: The proof of the theorems and relative analysis can be found in appendix.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental Result Reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: The experiment settings can be found in Section 5.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: Yes, We will release the experiment codes and documentations in GitHub (anonymous in the review phase).
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental Setting/Details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: Yes, the specific settings can be found in Section 5 and Appendix.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment Statistical Significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: Yes, the error bars are shown in the experiment results.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments Compute Resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: The computer resources are provided in the Appendix.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code Of Ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: Yes, it does.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader Impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [N/A]
+  49. Justification: This work does not involve issues like that.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: This work does not involve this kind of issues.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We used public benchmark and quoted it.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New Assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: We will release the experiment codes and documentations in GitHub (anonymous in the review phase).
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and Research with Human Subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: This work does not involve this kind of issues.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: This work does not involve this kind of issues.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v4

@@
 Traditional uses of the algorithm have intertwined reanalysis with training, while we suggests a novel paradigm, advocating for the decoupling of the reanalyze process from training iterations, thus providing a more flexible and efficient methodology.
-3.2 Stochastic Bandits
-In the upcoming Section 4.1, we plan to provide an inspiration case study behind our method, which is rooted in the fundamentals of bandit algorithms.
-Before diving into that, we’ll start with an overview of the key concepts surrounding stochastic bandits.
-A stochastic bandit $\nu$ with $K$ arms [26] can be formulated as $\nu=\{P_{i}|i\in{1,2,...,K}\}$, where $P_{i}$ denotes the reward distribution corresponding to the arm $i$.
-The expected reward of arm $i$ is denoted as $\mu_{i}(\nu)$.
-Given an algorithm $\pi$ interacting with the bandit $\nu$ for $n$ rounds, we define the regret $R_{n}$:
-$R_{n}(\pi,\nu)=n\mu^{*}(\nu)-\mathbb{E}[\sum_{t=1}^{n}X_{t}]$ (2)
-where $\mu^{*}(\nu)=\max_{a\in[K]}\mu_{a}(\nu)$, $X_{t}$ denotes the random reward at time $t$.
-The regret quantifies the loss of reward incurred by $\pi$ compared to the optimal policy on $\nu$ during $n$ rounds of interactions.
-In practice, we aim for $\pi$ to exhibit sub-linear regret growth across a category of bandits $\Omega$ , i.e.:
-$\lim_{n\to\infty}\frac{R_{n}(\pi,\nu)}{n}=0,\ \ \ \text{for all $\nu\in\Omega$ }$ (3)
-where sub-linear growth rate indicates $\pi$ is choosing the optimal arm almost all of the time with $n$ tends to infinity.
-UCB [8] is a classical bandit algorithm that achieves regret growth rate of $O(\log(n))$.
-AlphaZero [9] considers the child nodes in tree search as bandit arms and applies a variant of the UCB in the tree search process:
+3.2 Bandit-view Tree Search
+A stochastic bandit has $K$ arms, and playing each arm means sampling a reward from the corresponding distribution.
+For a search tree in MCTS, the root node can be seen as a bandit, with each child node as an arm.
+The left side of Figure 1 illustrates this idea.
+However, as the policy is continuously improved during the search process, the reward distribution for the arms should change over time.
+Therefore, UCT [27] modeled the root node as a non-stationary stochastic bandit with a drift condition:
+$\mathbb{P}(\overline{X}_{is}-\mu_{i}\geq\varepsilon)\leq\exp(-\frac{\varepsilon^{2}s}{C^{2}})\>\text{and}\>\mathbb{P}(\overline{X}_{is}-\mu_{i}\leq-\varepsilon)\leq\exp(-\frac{\varepsilon^{2}s}{C^{2}})$ (2)
+Where $\overline{X}_{is}$ is the average reward of the first $s$ samples of arm $i$.
+$\mu_{i}$ is the limit of $\mathbb{E}[\overline{X}_{is}]$ as $s$ approaches infinity, which indicates that the expectation of the node value converges.
+$C$ is an appropriate constant characterizes the rate of concentration.
+Based on this modeling, UCT uses the bandit algorithm UCB1 [8] to select child nodes. AlphaZero [9] inherits this concept and employs a variant formula:
 $UCB_{score}(s,a)=Q(s,a)+cP(s,a)\frac{\sqrt{{\sum}_{b}N(s,b)}}{1+N(s,a)}$ (3)
@@
 4 Method
-4.1 An Inspiring Case Study
-As a classic algorithm in the realm of stochastic bandits, Upper Confidence Bound 1 (UCB1) [8] has been deployed to select nodes within a search tree.
+4.1 Backward-view Reanalyze
+Our algorithm design stems from a simple inspiration: if we could know the true state-value of a child node in advance, we could save the search for it, thus conserving search time.
+As shown on the right side of Figure 1, we directly use the true expected value $\mu_{A}$ to evaluate the quality of Arm A, thereby eliminating the need for the back-propagated value to calculate the score in 3.
+This results in the process occurring within the gray box being omitted.
+Indeed, this situation can be well modeled as an one-armed bandit[18], which also facilitates our subsequent theoretical analysis.
 Figure 1: The connection between MCTS and bandits.
@@
 When action $A$ is selected, a reward $r_{A}$ will be returned, where $r_{A}=\sum_{t=1}^{3}\gamma^{t-1}r_{t}^{a}+\gamma^{3}v_{a}$.
-For the root node, the traversal, evaluation and back-propagation occurring in the gray box can be approximated as sampling from an non-stationary distribution.
+For the root node, the traversal, evaluation and back-propagation occurring in the sub-tree can be approximated as sampling from an non-stationary distribution.
 Thus the root node can be seen as a non-stationary bandit.
@@
 Once the reward $\mu_{A}$ is known, we can evaluate arm A using $\mu_{A}$, thereby eliminating the need to rely on subsequent tree search processes.
-As explained in Figure 1, each node in the tree can be regarded as a multi-armed bandit Notice that in the context of stochastic bandits, the reward for each arm is sampled from a stationary probability distribution.
-However, in the process of tree search, the value obtained from selecting a node depends on subsequent search process.
-The UCT [27] algorithm models this process as sampling from a non-stationary distribution which converges.
-For the sake of analytical convenience in this paper, we approximate the root node as a stochastic bandit and consider each child node as an arm with a deterministic reward distribution.
-Based on this modeling, we can analyze the Eq.
-4 employed in AlphaZero [9] and derive the following proposal:
-Proposal 1
-Denote the prior for arm $i$ as $P_{i}$.
-Without loss of generality, suppose $\mu^{*}=\mu_{1}$.
-Assume that the reward distribution of each arm $i$ is $1$-subgaussian and $P_{1}>\mu_{1}$, then the regret $R_{n}$ of using Eq.
-4 to choose arm satisfies
-$R_{n}\leq\sum_{i:\Delta_{i}>0}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(3+\frac{16}{\varepsilon^{4}P_{1}^{2}}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})$ (5)
-The assumption that $P_{1}>\mu_{1}$ is made purely for the sake of formality; in fact, Appendix A indicates that, as long as there exists a $t$ that $P_{1}\sqrt{t-1}>\mu_{1}$, a similar upper bound on regret can always be obtained.
-This elucidates that selecting actions according to Eq.
-4 can yield a sublinear growth of regret, indicating that the final visit distribution will be concentrated on the optimal arm (see analysis for Lemma A.1 in Appendix A).
-This also confirms the rationality of selecting the action with the largest number of visits.
-The inspiration of our enhanced method comes from the one-armed bandit model: if the expected reward of a particular arm is known as prior, the necessity for exploratory trials on this arm is obviated, thereby reducing the regret upper bound.
-To concretize this insight, we can extend Eq.
-4 and deduce the following proposition:
-Proposal 2
-For a bandit which satisfies the assumptions in Proposal 1, the regret upper bound can be reduced if we use the known expectation value of an arm to replace its score in Eq.
-4.
-Specifically, if the known arm is the optimal arm with value $\mu_{1}$, then the regret $R_{n}$ satisfies:
-$R_{n}\leq\sum_{i:\Delta_{i}>0}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(2+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})$ (6)
-In other case, with the known value $\mu_{l}$ of a sub-optimal arm $l$, the regret $R_{n}$ satisfies:
-$R_{n}$ $\leq\sum_{i:\Delta_{i}>0,i\neq l}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(3+\frac{16}{\varepsilon^{4}P_{1}^{2}}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})+\inf_{\varepsilon\in(0,\Delta_{l})}\Delta_{l}(1+\frac{16}{\varepsilon^{4}P_{1}^{2}})$ (7)
-Proposal 2 demonstrates that, given the known expected reward of a suboptimal arm, the asymptotic upper bound of regret can be further reduced, thereby concentrating the visit distribution more heavily on the optimal arm.
-Meanwhile, as shown in Figure 1, when the expectation $\mu_{A}$ is known, we no longer need to evaluate arm A through real-time search results.
-Instead, we can directly substitute $\mu_{A}$ for the score in Eq.
-4, thus saving the time spent on traverse and back-propagation processes.
-This inspires us to evaluate a particular child node using existing information, thereby eliminating the exploration of the corresponding sub-tree and accelerating the search process.
-See Appendix A for detailed proof and analysis of the two proposals.
-4.2 Backward-view Reanalyze
-Figure 2: An illustration about the backward-view reanalyze in both node and batch view.
-We sample $n+1$ trajectories of length $k+1$ to form a batch and conduct the search in the reverse direction of trajectories.
-From the node view, we would first search $S_{0}^{1}$ and then pass root value $m_{0}^{1}$ to $S_{0}^{0}$ to evaluate the value of a child node.
-$T_{0}^{1}$ and $T_{0}^{0}$ are the corresponding search trees.
-From the batch view, we would group all $S^{1}$s into a sub batch to search together and pass the root values to the $S^{0}$s.
-Inspired by the aforementioned motivation, we aspire to obtain the reward expectation of a child node in advance to assess it.
+Driven by the aforementioned motivation, we aspire to obtain the reward expectation of a child node in advance.
 However, the true expectation is always unknown.
 Consequently, we resort to using the root value obtained from MCTS as an approximate substitute.
-Specifically, for two adjacent time steps $S_{0}^{1}$ and $S_{0}^{0}$ in Figure 2, when searching for state $S_{0}^{1}$, the root node corresponds to a child node of $S_{0}^{0}$.
+Specifically, for two adjacent time steps $S_{0}^{1}$ and $S_{0}^{0}$  [^1] The subscript denotes the trajectory, the superscript denotes the time step in the trajectory.
+in Figure 2, when searching for state $S_{0}^{1}$, the root node corresponds to a child node of $S_{0}^{0}$.
 Therefore, the root value obtained can be utilized to assist the search for $S_{0}^{0}$.
@@
 During reanalyze, since the trajectories was already collected in the replay buffer, we can perform tree search in a backward-view, which searches the trajectory in a reverse order.
-Figure 2 illustrates a batch containing $n+1$ trajectories, each of length $k+1$.
-We denote $S_{l}^{t}$ as the $t$-th state in the $l$-th trajectory.
+Figure 2 illustrates a batch containing $n+1$ trajectories, each of length $k+1$, and $S_{l}^{t}$ is the $t$-th state in the $l$-th trajectory.
 We first conduct a search for all $S^{k}$s, followed by a search for all $S^{k-1}$s, and so on.
@@
 Since the time used to search for node $S^{t+1}_{l}$ is saved, this enhanced search process is faster than the original version.
-Notice that the analysis in Proposal 2 assumes that we know the expectation value.
-However, in our design, we use the root value as an approximate estimation.
-Still based on the stochastic bandit assumption, the following theorem analyzes the impact of the error between these two values:
-Theorem 1
-For a bandit which satisfies assumptions in Proposal 1 and a sample result for an arm, the regret upper bound can be reduced if we use the way of Equation 8 to choose arm.
-Specifically, if the sampled arm is the optimal arm with $n$ times sample mean $\hat{\mu}_{1n}$, then the regret $R_{n}$ satisfies:
-$R_{n}\leq\sum_{i:\Delta_{i}>0}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(2+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+n\exp(-\frac{n\varepsilon^{2}}{2}))$ (10)
-In other case, with the $n$ times sample mean $\hat{\mu}_{ln}$ of a sub-optimal arm $l$, the regret $R_{n}$ satisfies:
-$R_{n}$ $\leq\sum_{i:\Delta_{i}>0,i\neq l}\inf_{\varepsilon\in(0,\Delta_{i})}\Delta_{i}(3+\frac{16}{\varepsilon^{4}P_{1}^{2}}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon})$ (11)
-$+$ $\inf_{\varepsilon\in(0,\Delta_{l})}\Delta_{l}(1+\frac{16}{\varepsilon^{4}P_{1}^{2}}+n\exp(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{2}))$ (12)
-This theorem demonstrates that using the sampled mean as a substitute for the expectation value can also reduce the asymptotic upper bound of regret and maintain sub linear growth rate.
-This can be understood as follows: although there is an error between the sampled mean and the expectation value, since we use the average of $n$ samples (as the root value comes from a tree search which has $n$ simulations), this sampled mean will increasingly approximate the expectation value as $n$ grows, thereby preventing a linear increase.
-Refer to Appendix A for the proof.
+Figure 2: An illustration about the backward-view reanalyze in both node and batch view.
+We sample $n+1$ trajectories of length $k+1$ to form a batch and conduct the search in the reverse direction of trajectories.
+From the node view, we would first search $S_{0}^{1}$ and then pass root value $m_{0}^{1}$ to $S_{0}^{0}$ to evaluate the value of a child node.
+$T_{0}^{1}$ and $T_{0}^{0}$ are the corresponding search trees.
+From the batch view, we would group all $S^{1}$s into a sub batch to search together and pass the root values to the $S^{0}$s.
 Algorithm 1 Python-like code for information reuse
@@
 backpropagate()
+4.2 Theoretical Analysis
+AlphaZero selects child nodes using Equation 3 and takes the final action based on the visit counts.
+In the previous section, we replace Equation 3 with Equation 5, which undoubtedly impacts the visit distribution of child nodes.
+Additionally, since we use the root value as an approximation to true expectation, the error between the two may also affect the search results.
+To demonstrate the reliability of our algorithm, in this section, we model the root node as an non-stationary bandit and prove that, as the number of total visit increases, the visit distribution gradually concentrates on the optimal arm.
+Specifically, we have the following theorem:
+Theorem 1
+For a non-stationary bandit that conforms to the assumptions of Equation 2, denote the total number of rounds as $n$, and the times a sub-optimal arm $i$ is selected in $n$ rounds as $T_{i}(n)$, then selecting actions according to Equation 5 can ensure that $\frac{{\mathbb{E}[T_{i}(n)]}}{n}\to 0$ as $n\to\infty$.
+Specifically, if we know the $n$ times sample mean $\hat{\mu}^{*}_{n}$ of the optimal arm in advance, then $\mathbb{E}[T_{i}(n)]$ for all sub-optimal arm $i$ satisfies
+$\mathbb{E}[T_{i}(n)]\leq 2+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{C^{2}}{(\Delta_{i}-\varepsilon)^{2}}+n\exp{(-\frac{n\varepsilon^{2}}{C^{2}})}$ (6)
+Otherwise, if we have the $n$ times sample mean $\hat{\mu}_{ln}$ of a sub-optimal arm $l$, then for arm $l$,
+$\mathbb{E}[T_{l}(n)]\leq 1+\frac{4C^{4}}{\varepsilon^{4}P_{1}^{2}}+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{C^{2}})}$ (7)
+and for other sub-optimal arms,
+$\mathbb{E}[T_{i}(n)]\leq 3+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{C^{2}}{(\Delta_{i}-\varepsilon)^{2}}+\frac{4C^{4}}{\varepsilon^{4}P_{1}^{2}}$ (8)
+where $\Delta_{i}$ is the optimal gap for arm $i$, $\varepsilon$ is a constant in $(0,\Delta_{i})$, and $C$ is the constant in Equation 2.
+We provide the complete proof and draw similar conclusions for AlphaZero in Appendix A.
+Additionally, our method has a lower upper bound for $\mathbb{E}[T_{i}(n)]$.
+This implies that our algorithm may yield visit distribution more concentrated on the optimal arm.
+This is a potential worth exploring in future work, especially in offline scenarios [13] where reanalyze becomes the sole method for policy improvement.
 4.3 The ReZero Framework
@@
 Lemma A.1.
-Regret decomposition lemma:Denote $X_{t}$ as the reward at time $t$, $A_{t}$ as the arm chosen at time $t$, and $S_{n}=\sum_{t}X_{t}$, then the regret can be decomposed by regret caused by every sub-optimal arm:
-$R_{n}=\sum_{i}E[T_{i}]\Delta_{i}$ (13)
-Proof.
-Note that
-$R_{n}$ $=n\mu^{*}-E[S_{n}]$ (14)
-$=\sum_{i}\sum_{t=1}^{n}E[(\mu^{*}-X_{t})\mathbb{I}\{A_{t}=i\}]$ (15)
-and
-$E[(\mu^{*}-X_{t})\mathbb{I}\{A_{t}=i\}|A_{t}]$ (16)
-$=\mathbb{I}\{A_{t}=i\}E[\mu^{*}-X_{t}|A_{t}]$ (17)
-$=\mathbb{I}\{A_{t}=i\}(\mu^{*}-\mu_{A_{t}})$ (18)
-$=\mathbb{I}\{A_{t}=i\}(\mu^{*}-\mu_{i})$ (19)
-$=\mathbb{I}\{A_{t}=i\}\Delta_{i}$ (20)
-The proof can be completed by combining the two equations and using the definition of $T_{i}$. ∎
-This lemma demonstrates that for a certain bandit, if the regret grows sub linearly, then the expected times each sub optimal arm is chosen also grows sub linearly.
-Consequently, as $n$ increases, the proportion of times the sub optimal arm is selected converges to 00.
-In other words, the visit count distribution concentrates on the optimal arm as $n$ grows.
-Lemma A.2.
-Let $X_{1},...,X_{n}$ be a sequence of independent 1-subgaussian random variables, $\hat{\mu_{t}}=\frac{1}{t}\sum^{t}_{s=1}X_{s}$, $\varepsilon>0$, $a>0$ and
+[…]
 $\kappa=\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{t}+\sqrt{\frac{a}{t^{2}}}\geq\varepsilon\}$ (9)
-, then it holds that $E[\kappa]\leq 1+\frac{2\sqrt{a}}{\varepsilon}+\frac{2}{\varepsilon^{2}}$.
+, then it holds that $E[\kappa]\leq 1+\frac{2\sqrt{a}}{\varepsilon}+\frac{C^{2}}{\varepsilon^{2}}$.
 Proof.
 Take $u$ as $\frac{2\sqrt{a}}{\varepsilon}$, then
-$E[\kappa]$ $\leq u+\sum_{t=\lceil u\rceil}^{n}\mathbb{P}(\hat{\mu}_{t}\sqrt{\frac{a}{t^{2}}}\geq\varepsilon)$ (23)
-$\leq u+\sum_{t=\lceil u\rceil}^{n}\exp(-\frac{t(\varepsilon-\sqrt{\frac{a}{t^{2}}})^{2}}{2})$ (24)
+$E[\kappa]$ $\leq u+\sum_{t=\lceil u\rceil}^{n}\mathbb{P}(\hat{\mu}_{t}+\sqrt{\frac{a}{t^{2}}}\geq\varepsilon)$ (10)
+$\leq u+\sum_{t=\lceil u\rceil}^{n}\exp(-\frac{t(\varepsilon-\sqrt{\frac{a}{t^{2}}})^{2}}{C^{2}})$ (11)
 […]
 […]
-$=1+u+\frac{2}{\varepsilon^{2}}=1+\frac{2\sqrt{a}}{\varepsilon}+\frac{2}{\varepsilon^{2}}$ (27)
-Proof for Proposal 1:
+$=1+u+\frac{C^{2}}{\varepsilon^{2}}=1+\frac{2\sqrt{a}}{\varepsilon}+\frac{C^{2}}{\varepsilon^{2}}$ (14)
+Proof for Theorem 1:
 Proof.
-Denote $T_{i}(k)$ as the times that arm $i$ has been chosen until time $k$, $\hat{\mu}_{is}$ as the empirical mean based on the first $s$ samples from distribution $P_{i}$, and $\hat{\mu}_{i}(k)=\hat{\mu}_{iT_{i}(k)}$.
+We first present the upper bound of $\mathbb{E}[T_{i}(n)]$ for using the Equation 3 in AlphaZero.
+A slight adjustment to this proof yields the conclusion of Theorem 1.
+Denote $T_{i}(k)$ as the times that arm $i$ has been chosen until time $k$, $\hat{\mu}_{is}$ as the average of the first $s$ samples of arm $i$, which satisfies the concentration assumption in Equation 2, and $\hat{\mu}_{i}(k)=\hat{\mu}_{iT_{i}(k)}$.
 Then we have:
@@
 $+\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}$ (17)
-for 29, we have
+for 16, we have
 $E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]$ (18)
@@
 $=1+\sum_{t=2}^{n}\sum_{s=1}^{t-1}\mathbb{P}(\hat{\mu}_{1s}+P_{1}\frac{\sqrt{t-1}}{1+s}\leq\mu_{1}-\varepsilon)$ (20)
-$\leq 1+\sum_{t=2}^{n}\sum_{s=1}^{t-1}\exp(-\frac{s(\varepsilon+P_{1}\frac{\sqrt{t-1}}{1+s})^{2}}{2})$ (34)
-$\leq 1+\sum_{t=2}^{n}\exp(-\frac{1}{2}\varepsilon P_{1}\sqrt{t-1})\sum_{s=1}^{t-1}\exp(-\frac{s\varepsilon^{2}}{2})$ (35)
-$\leq 1+\sum_{t=2}^{n}\exp(-\frac{1}{2}\varepsilon P_{1}\sqrt{t-1})\frac{2}{\varepsilon^{2}}$ (36)
-$\leq 1+\frac{16}{\varepsilon^{4}P_{1}^{2}}$ (37)
-Notes: In 32, since we assume $P_{1}\geq\mu_{1}$, the probability of the term $s=0$ would be 00.
+$\leq 1+\sum_{t=2}^{n}\sum_{s=1}^{t-1}\exp(-\frac{s(\varepsilon+P_{1}\frac{\sqrt{t-1}}{1+s})^{2}}{C^{2}})$ (21)
+$\leq 1+\sum_{t=2}^{n}\exp(-\frac{1}{C^{2}}\varepsilon P_{1}\sqrt{t-1})\sum_{s=1}^{t-1}\exp(-\frac{s\varepsilon^{2}}{C^{2}})$ (22)
+$\leq 1+\sum_{t=2}^{n}\exp(-\frac{1}{C^{2}}\varepsilon P_{1}\sqrt{t-1})\frac{C^{2}}{\varepsilon^{2}}$ (23)
+$\leq 1+\frac{4C^{4}}{\varepsilon^{4}P_{1}^{2}}$ (24)
+Notes: In 19, since we assume $P_{1}\geq\mu_{1}$, the probability of the term $s=0$ would be 00.
 Thus, we can discard it.
@@
 This only changes the constant term and won’t affect the growth rate of regret.
-From 34 to 35, we just need to expand the quadratic term and do some simple inequality scaling.
-From 35 to 36, we need to notice that $\sum_{s=1}^{t-1}\exp(-\frac{s\varepsilon^{2}}{2})$ is a geometric sequence and scale it to $\frac{2}{\varepsilon^{2}}$.
-From 36 to 37, we use the inequality $\sum_{t=2}^{n}\frac{1}{e^{a\sqrt{t-1}}}\leq\int_{1}^{\infty}\frac{1}{e^{a\sqrt{t-1}}}$.
-for 30, we have
+From 21 to 22, we just need to expand the quadratic term and do some simple inequality scaling.
+From 22 to 23, we need to notice that $\sum_{s=1}^{t-1}\exp(-\frac{s\varepsilon^{2}}{C^{2}})$ is a geometric sequence and scale it to $\frac{C^{2}}{\varepsilon^{2}}$.
+From 23 to 24, we use the inequality $\sum_{t=2}^{n}\frac{1}{e^{a\sqrt{t-1}}}\leq\int_{1}^{\infty}\frac{1}{e^{a\sqrt{t-1}}}$.
+for 17, we have
 $E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (25)
@@
 with Lemma A.1, we can have
-[…]
+282+$\leq 2+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{C^{2}}{(\Delta_{i}-\varepsilon)^{2}}$ (29)
 so
-$E[T_{i}(n)]\leq 3+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+\frac{16}{\varepsilon^{4}P_{1}^{2}}$ (43)
-combined with Lemma A.1, the proof is completed. ∎
-[…]
-Proof for Proposal 2:
-Proof.
-The proof can be obtained by making slight modifications to the proof of Proposal 1. In case that $\mu_{1}$ is already known,
-$E[T_{i}(n)]=E\sum_{t=1}^{n}\mathbb{I}\{A_{t}=i\}]$ (44)
-$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\mu_{1}\leq\mu_{1}-\varepsilon\}]$ (45)
-$+E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (46)
-$=E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (47)
-$\leq 2+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}$ (48)
-In case that a sub-optimal value $\mu_{l}$ is known,
-$E[T_{l}(n)]=E[\sum_{t=1}^{n}\mathbb{I}\{A_{t}=l\}]$ (49)
-$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]$ (50)
-$+E[\sum_{t=1}^{n}\mathbb{I}\{\mu_{l}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (51)
-$=E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]$ (52)
-$\leq 1+\frac{16}{\varepsilon^{4}P_{1}^{2}}$ (53)
-and the bound of $T_{i}(n)$ for $i\neq l$ keeps unchanged. In both cases, combined with Lemma A.1, the proof is completed. ∎
-From the proof, it can be seen that if the real value $\mu_{1}$ of the optimal arm is known in advance, the regret caused by the underestimation of the optimal arm can be eliminated.
-However, the part of the regret that is reduced is $O(1)$.
-If the real value $\mu_{i}$ of a sub-optimal arm $l$ is known in advance, the regret caused by continuous exploration of this sub-optimal arm can be eliminated.
-The part of the reduced regret is $O(\sqrt{n})$.
-Intuitively, it can be understood this way: If the value of a sub-optimal arm $l$ is known, then after sufficient exploration of the optimal arm, the upper confidence bound will converge to be higher than $\mu_{l}$ with probability $1$, thus making the regret caused by arm $l$ no longer increases with $n$.
-While if the value of the optimal arm is known, then as the number of simulations $n$ increases, there will always be a moment when the upper confidence bound of the sub-optimal arm $i$ is greater than $\mu_{1}$, so arm $i$ will still be repeatedly selected, leading to the regret caused by arm $i$ increases with $n$.
-However, in this case, since the optimal arm is chosen the most time as $n$ increases, we can save more time by early stopping the search.
-Proof for Theorem 1:
-Proof.
-The proof can be obtained by making slight modifications to the proof of Proposal 2.
-In case we have $\hat{\mu}_{1n}=\frac{1}{n}\sum_{t=1}^{n}\hat{\mu}_{t}$ where $\hat{\mu}_{t}$ is sampled from the same reward distribution of arm $1$,
+$E[T_{i}(n)]\leq 3+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{C^{2}}{(\Delta_{i}-\varepsilon)^{2}}+\frac{4C^{4}}{\varepsilon^{4}P_{1}^{2}}$ (30)
+The proof of theorem 1 can be obtained by making slight modifications.
+In case we have $\hat{\mu}^{*}_{n}=\frac{1}{n}\sum_{t=1}^{n}\hat{\mu}_{t}$ where $\hat{\mu}_{t}$ is sampled from the optimal arm in advance,
 $E[T_{i}(n)]=E[\sum_{t=1}^{n}\mathbb{I}\{A_{t}=i\}]$ (31)
@@
 $+E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (33)
-$\leq n\exp{(-\frac{n\varepsilon^{2}}{2})}+E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (57)
-$\leq 2+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{2}{(\Delta_{i}-\varepsilon)^{2}}+n\exp{(-\frac{n\varepsilon^{2}}{2})}$ (58)
+$\leq n\exp{(-\frac{n\varepsilon^{2}}{C^{2}})}+E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (34)
+$\leq 2+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{C^{2}}{(\Delta_{i}-\varepsilon)^{2}}+n\exp{(-\frac{n\varepsilon^{2}}{C^{2}})}$ (35)
 In case we have $\hat{\mu}_{ln}=\frac{1}{n}\sum_{t=1}^{n}\hat{\mu}_{t}$ where $\hat{\mu}_{t}$ is sampled from the same reward distribution of arm $l$,
@@
 $+E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{ln}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (38)
-$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{2})}$ (62)
-$\leq 1+\frac{16}{\varepsilon^{4}P_{1}^{2}}+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{2})}$ (63)
-and the bound of $E[T_{i}(n)]$ for $i\neq l$ keeps unchanged.
-In both cases, combined with Lemma A.1, the proof is completed. ∎
+$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{C^{2}})}$ (39)
+$\leq 1+\frac{4C^{4}}{\varepsilon^{4}P_{1}^{2}}+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{C^{2}})}$ (40)
+and the bound of $E[T_{i}(n)]$ for $i\neq l$ keeps unchanged. ∎
 Appendix B Implementation Details
@@
 We ultimately set the batch size to 2000, which yields the fastest average search speed on our device.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: The claims made matches theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: Section 6 discusses the limitations of the paper.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes]
-  14. Justification: The proof of the theorems and relative analysis can be found in appendix.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: The experiment settings can be found in Section 5.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: Yes, We will release the experiment codes and documentations in GitHub (anonymous in the review phase).
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: Yes, the specific settings can be found in Section 5 and Appendix.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: Yes, the error bars are shown in the experiment results.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: The computer resources are provided in the Appendix.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: Yes, it does.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [N/A]
-  49. Justification: This work does not involve issues like that.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: This work does not involve this kind of issues.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: We used public benchmark and quoted it.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: We will release the experiment codes and documentations in GitHub (anonymous in the review phase).
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: This work does not involve this kind of issues.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: This work does not involve this kind of issues.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v5

@@
 To address this issue, we propose a general approach named ReZero to boost tree search operations for MCTS-based algorithms.
-Specifically, drawing inspiration from the one-armed bandit model, we reanalyze training samples through a backward-view reuse technique which obtains the value estimation of a certain child node in advance.
+Specifically, drawing inspiration from the one-armed bandit model, we reanalyze training samples through a backward-view reuse technique which uses the value estimation of a certain child node to save the corresponding sub-tree search time.
 To further adapt to this design, we periodically reanalyze the entire buffer instead of frequently reanalyzing the mini-batch.
 The synergy of these two designs can significantly reduce the search cost and meanwhile guarantee or even improve performance, simplifying both data collecting and reanalyzing.
-Experiments conducted on Atari environments and board games demonstrate that ReZero substantially improves training speed while maintaining high sample efficiency.
-The code is available as part of the LightZero benchmark at https://github.com/opendilab/LightZero.
+Experiments conducted on Atari environments, DMControl suites and board games demonstrate that ReZero substantially improves training speed while maintaining high sample efficiency.
+The code is available as part of the LightZero MCTS benchmark at https://github.com/opendilab/LightZero.
 1 Introduction
-As a pivotal subset of artificial intelligence, Reinforcement Learning (RL) [1] has acquired achievements and applications across diverse fields, including interactive gaming [2], autonomous vehicles [3], and natural language processing [4].
+As a pivotal subset of artificial intelligence, Reinforcement Learning (RL) (Sutton & Barto, 1988) has acquired achievements and applications across diverse fields, including interactive gaming (Vinyals et al., 2019), autonomous vehicles (Li et al., 2022), and natural language processing (Rafailov et al., 2023).
 Despite its successes, a fundamental challenge plaguing traditional model-free RL algorithms is their low sample efficiency.
 These algorithms typically require a large amount of data to learn effectively, making them often infeasible in the real-world scenarios.
-In response to this problem, numerous model-based RL [5, 6] methods have emerged.
+In response to this problem, numerous model-based RL methods (Janner et al., 2019; Hafner et al., 2020) have emerged.
 These methods involve learning an additional model of the environment from data and utilizing the model to assist the agent’s learning, thereby improving sample efficiency significantly.
-Within this field, Monte Carlo Tree Search [7] (MCTS) has been proven to be a efficient method for utilizing models for planning.
-It incorporates the UCB1 algorithm [8] into the tree search process and has achieved promising results in in a wide range of scenarios.
-Specifically, AlphaZero [9] plays a big role in combining deep reinforcement learning with MCTS, achieving notable accomplishments that can beat top-level human players.
-While it can only be applied to environments with perfect simulators, MuZero [10] extended the algorithm to cases without known environment models, resulting in good performances in a wider range of tasks.
-Following MuZero, many successor algorithms have emerged, enabling MuZero to be applied in continuous action spaces [11], environments with stochastic dynamics [12], offline RL training scenarios [13], and etc.
-All these MCTS-based algorithms [9, 10, 14] made valuable contributions to the universal applicability of the MCTS+RL paradigm.
+Within this field, Monte Carlo Tree Search (MCTS) (Świechowski et al., 2023) has been proven to be a efficient method for utilizing models for planning.
+It incorporates the UCB1 algorithm (Auer et al., 2002) into the tree search process and has achieved promising results in a wide range of scenarios.
+Specifically, AlphaZero (Silver et al., 2017) plays a big role in combining deep reinforcement learning with MCTS, achieving notable accomplishments that can beat top-level human players.
+While it can only be applied to environments with perfect simulators, MuZero (Schrittwieser et al., 2019) extended the algorithm to cases without known environment models, resulting in good performances in a wider range of tasks.
+Following MuZero, many successor algorithms have emerged, enabling MuZero to be applied in continuous action spaces (Hubert et al., 2021), offline RL training scenarios (Schrittwieser et al., 2021), and etc.
+All these MCTS-based algorithms made valuable contributions to the universal applicability of the MCTS+RL paradigm.
 However, the extensive tree search computations incur additional time overhead for these algorithms: During the data collection phase, the agent needs to execute MCTS to select an action every time it receives a new state.
-Furthermore, due to the characteristics of tree search, it is challenging to parallelize it using commonly used vectorized environments [15], further amplifying the speed disadvantage.
-On the other hand, during the reanalyze [13] process, in order to obtain higher-quality update targets, the latest model parameters are used to re-run MCTS on the training mini-batch.
-The time required to process samples thus increases as a trade-off for high sample efficiency.
-The excessive time cost has become a bottleneck hindering the further promotion of these algorithms.
-Recently, a segment of research endeavors are directed toward mitigating the above wall-clock time overhead about MCTS-based algorithms.
-On the one hand, SpeedyZero [16] greatly diminishes algorithms’ time overheads by deploying a parallel execution training pipeline; however, it demands additional computational resources.
+Furthermore, due to the characteristics of tree search, it is challenging to parallelize it using commonly used vectorized environments (Weng et al., 2022), further amplifying the speed disadvantage.
+On the other hand, during the reanalyze (Schrittwieser et al., 2021) process, in order to obtain higher-quality update targets, the latest models are used to re-run MCTS on the training mini-batch.
+The wall-clock time thus increases as a trade-off for high sample efficiency.
+The excessive cost has become a bottleneck hindering the further promotion of these algorithms.
+Recently, a segment of research endeavors are directed toward mitigating the above wall-clock time overhead.
+On the one hand, SpeedyZero (Mei et al., 2023) diminishes algorithms’ time overheads by deploying a parallel execution training pipeline; however, it demands additional computational resources.
 On the other hand, it remains imperative to identify methodologies that accelerate these algorithms without imposing extra demands.
-PTSAZero [17], for instance, compresses the search space via state abstraction, decreasing the time cost per search.
-Nevertheless, it is widely acknowledged that MCTS-based algorithms necessitate search times (the number of simulations) to assure their performance.
-From this perspective, our research try to discover strategies that preserve algorithmic performance while minimizing the search times as much as possible.
-In this paper, we introduce ReZero, a new approach designed to boost the MCTS-based algorithms.
-Firstly, inspired by the one-armed bandit model [18], we propose a backward-view reanalyze technique that proceeds in the reverse direction of the trajectories, utilizing previously searched root values to bypass the exploration of specific child nodes, thereby saving time.
-We present several propositions to articulate our insights and analyzed the convergence of the algorithm based on the bandit model.
+PTSAZero (Fu et al., 2023), for instance, compresses the search space via state abstraction, decreasing the time cost per search.
+In contrast, we aim to adopt a method that is orthogonal to both of the previous approaches.
+It does not require state space compression but directly reduces the search space through value estimation, and it does not introduce additional hardware overhead.
+In this paper, we introduce ReZero, a new approach/framework designed to boost the MCTS-based algorithms.
+Firstly, inspired by the one-armed bandit model (Lattimore & Szepesvári, 2020), we propose a backward-view reanalyze technique that proceeds in the reverse direction of the trajectories, utilizing previously searched root values to bypass the exploration of specific child nodes, thereby saving time.
+Additionally, we have proven the convergence of our search mechanism based on the non-stationary bandit model, i.e., the distribution of child node visits will concentrate on the optimal node.
 Secondly, to better adapt to our proposed backward-view reanalyze technique, we have devised a novel pipeline that concentrates MCTS calls within the reanalyze process and periodically reanalyzes the entire buffer after a fixed number of training iterations.
 This entire-buffer reanalyze not only reduces the number of MCTS calls but also better leverages the speed advantages of parallelization.
-Our approach can be easily applied to the MCTS-based algorithm family and synergistically implemented alongside existing acceleration methods.
+Skipping the search of specific child nodes can be seen as a pruning operation, which is common in different tree search settings.
+Reanalyze is also a common module in MCTS-based algorithms.
+Therefore, our algorithm design is universal and can be easily applied to the MCTS-based algorithm family.
 In addition, it will not bring about any overhead in computation resources.
-Empirical experiments (Section 5) show that our approach yields good results in both single-agent environment (Atari [19]) and two-player board games [14], greatly improving the training speed while maintaining or even improving sample efficiency.
+Empirical experiments (Section 5) show that our approach yields good results in both single-agent discrete-action environment (Atari (Bellemare et al., 2013)), two-player board games (Silver et al., 2016), and continuous control suites (Tunyasuvunakool et al., 2020), greatly improving the training speed while maintaining or even improving sample efficiency.
 Ablation experiments explore the impact of different reanalyze frequencies and the acceleration effect of backward-view reuse on a single search.
 The main contributions of this paper can be summarized as follows:
-  • We design a method to speed up a single search by the backward-view reanalyze technique. Theoretical support for the effectiveness of our proposed method is also provided.
-  • We propose a simplified framework with the entire-buffer reanalyze mechanims that further reduces the number of MCTS calls and enhance its parallelization, thus boosting MCTS-based algorithms.
-  • We conduct several experiments on diverse environments and investigate ReZero through ablations.
+  • We design a method to speed up a single tree search by the backward-view reanalyze technique. Theoretical support for the convergence of our proposed method is also provided.
+  • We propose an efficient framework with the entire-buffer reanalyze mechanism that further reduces the number of MCTS and enhance its parallelization, boosting MCTS-based algorithms.
+  • We conduct experiments on diverse environments and investigate ReZero through ablations.
 2 Related Work
 2.1 MCTS-based Algorithms
-AlphaGo [14] and AlphaZero [9] combined MCTS with deep RL, achieving significant results in board games, defeating world-champion players, and revealing their remarkable capabilities.
-Following these achievements, MuZero [10] combines tree search with a learned value equivalent model [20], successfully promoting the algorithm’s application to scenarios without known models, such as Atari.
+AlphaGo (Silver et al., 2016) and AlphaZero (Silver et al., 2017) combined MCTS with deep RL, achieving significant results in board games, defeating world-champion players, and revealing their remarkable capabilities.
+Following these achievements, MuZero (Schrittwieser et al., 2019) combines tree search with a learned value equivalent model (Grimm et al., 2020b), successfully promoting the algorithm’s application to scenarios without known models, such as Atari.
 Subsequent to MuZero’s development, several new works based on the MuZero framework emerged.
-EfficientZero [21] further improved MuZero’s sample efficiency, using very little data for training and still achieving outstanding performance.
-Another works [11, 12] extended the MuZero agent to environments with large action spaces and stochastic environments, further broadening the algorithm’s usage scenario.
-MuZero Unplugged [13] applied MuZero’s reanalyze operation as a policy enhancement operator and extended it to offline settings.
-Gumbel MuZero [22] changed the node selection method during tree search, so the policy obtained from the search shows a monotonic improvement, ensuring algorithm performance even with low simulation numbers.
-Despite the aforementioned excellent improvement techniques, the current time overhead of MCTS-based RL remains significant, which is the main problem this paper aims to address.
+EfficientZero (Ye et al., 2021) further improved MuZero’s sample efficiency, using very little data for training and still achieving outstanding performance.
+Another works (Hubert et al., 2021; Antonoglou et al., 2021) extended the MuZero agent to environments with large action spaces and stochastic environments, further broadening the algorithm’s usage scenario.
+MuZero Unplugged (Schrittwieser et al., 2021) applied MuZero’s reanalyze operation as a policy enhancement operator and extended it to offline settings.
+Despite the aforementioned excellent improvement techniques, the time overhead of MCTS-based RL remains significant, which is the main problem this paper aims to address.
 2.2 MCTS Acceleration
 Recent research has focused on accelerating MCTS-based algorithms.
-The work [16] reduces the algorithm’s time overhead by designing a parallel system; however, this method requires more computational resources and involves some adjustments for large batch training.
-PTSAZero [17] narrows the search space through state abstraction, which amalgamates redundant information, thereby reducing the time cost per search.
-KataGo [23] use a naive information reuse trick.
+Mei et al.
+(2023) reduces the algorithm’s time overhead by designing a parallel system; however, this method requires more computational resources and involves some adjustments for large batch training.
+Fu et al.
+(2023) narrows the search space through state abstraction, which amalgamates redundant information, thereby reducing the time cost per search.
+KataGo (Wu, 2019) use a naive information reuse trick.
 They save sub-trees of the search tree and serve as initialization for the next search.
@@
 To our knowledge, we are the first to enhance MCTS by reusing information in a backward-view.
-Our proposed approach can seamlessly integrate with various MCTS-based algorithms, many of which [16, 17, 22, 24] are orthogonal to our contributions.
+Our proposed approach can seamlessly integrate with various MCTS-based algorithms, many of which (Mei et al., 2023; Fu et al., 2023; Danihelka et al., 2022; Leurent & Maillard, 2020) are orthogonal to our contributions.
 3 Preliminaries
 3.1 MuZero
-MuZero [10] is a fundamental model-based RL algorithm that incorporates a value-equivalent model [25] and leverages MCTS for planning within a learned latent space.
+MuZero (Schrittwieser et al., 2019) is a fundamental model-based RL algorithm that incorporates a value-equivalent model (Grimm et al., 2020a) and leverages MCTS for planning within a learned latent space.
 The model consists of three core components: a representation model $h_{\theta}$, a dynamics model $g_{\theta}$, and a prediction model $f_{\theta}$:
 $\begin{array}[]{lll}\text{Representation:}&&s_{t}=h_{\theta}(o_{t-l:t})\\ \text{Dynamics:}&&s_{t+1},r_{t}=g_{\theta}(s_{t},a_{t})\\ \text{Prediction: }&&v_{t},p_{t}=f_{\theta}(s_{t})\\ \end{array}$ (1)
-The representation model transforms a last observation sequences $o_{t-l:t}$ into a corresponding latent state $s_{t}$.
-The dynamics model processes this latent state alongside an action $a_{t}$, yielding the subsequent latent state $s_{t+1}$ and an estimated reward $r_{t}$.
+The representation model transforms last observation sequences $o_{t-l:t}$ into a corresponding latent state $s_{t}$.
+The dynamics model processes this latent state alongside an action $a_{t}$, yielding the next latent state $s_{t+1}$ and an estimated reward $r_{t}$.
 Finally, the prediction model accepts a latent state and produces both the predicted policy $p_{t}$ and the state’s value $v_{t}$.
@@
 Lastly the agent samples the best action $a_{t}$ following the searched visit count distribution.
-MuZero Reanalyze is an advanced version of the original MuZero algorithm.
-This variant enhances the model’s accuracy by conducting a fresh MCTS on sampled states with the most recent version of the model, subsequently utilizing the refined policy from this search to update the policy targets.
+MuZero Reanalyze is an advanced version of the original MuZero.
+This variant enhances the model’s accuracy by conducting a fresh MCTS on sampled states with the latest model, subsequently utilizing the refined policy from this search to update the policy targets.
 Such reanalysis yields targets of superior quality compared to those obtained during the initial data collection.
-Traditional uses of the algorithm have intertwined reanalysis with training, while we suggests a novel paradigm, advocating for the decoupling of the reanalyze process from training iterations, thus providing a more flexible and efficient methodology.
+Traditional uses of the algorithm have intertwined reanalysis with training, while we suggest a novel paradigm, advocating for the decoupling of the reanalyze process from training iterations, thus providing a more flexible and efficient methodology.
+Refer to the Appendix B for more details on MuZero during the training and inference phases.
 3.2 Bandit-view Tree Search
@@
 However, as the policy is continuously improved during the search process, the reward distribution for the arms should change over time.
-Therefore, UCT [27] modeled the root node as a non-stationary stochastic bandit with a drift condition:
-$\mathbb{P}(\overline{X}_{is}-\mu_{i}\geq\varepsilon)\leq\exp(-\frac{\varepsilon^{2}s}{C^{2}})\>\text{and}\>\mathbb{P}(\overline{X}_{is}-\mu_{i}\leq-\varepsilon)\leq\exp(-\frac{\varepsilon^{2}s}{C^{2}})$ (2)
-Where $\overline{X}_{is}$ is the average reward of the first $s$ samples of arm $i$.
-$\mu_{i}$ is the limit of $\mathbb{E}[\overline{X}_{is}]$ as $s$ approaches infinity, which indicates that the expectation of the node value converges.
+Therefore, UCT (Kocsis & Szepesvári, 2006) modeled the root node as a non-stationary stochastic bandit with a drift condition:
+$\mathbb{P}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{\mu}_{is}}-\mu_{i}\geq\varepsilon)\leq\exp(-\frac{\varepsilon^{2}s}{C^{2}})\>\text{and}\>\mathbb{P}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{\mu}_{is}}-\mu_{i}\leq-\varepsilon)\leq\exp(-\frac{\varepsilon^{2}s}{C^{2}})$ (2)
+Where ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{\mu}_{is}}$ is the average reward of the first $s$ samples of arm $i$.
+$\mu_{i}$ is the limit of $\mathbb{E}[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{\mu}_{is}}]$ as $s$ approaches infinity, which indicates that the expectation of the node value converges.
 $C$ is an appropriate constant characterizes the rate of concentration.
-Based on this modeling, UCT uses the bandit algorithm UCB1 [8] to select child nodes. AlphaZero [9] inherits this concept and employs a variant formula:
+Based on this modeling, UCT uses the bandit algorithm UCB1 (Auer et al., 2002) to select child nodes. AlphaZero inherits this concept and employs a variant formula:
 $UCB_{score}(s,a)=Q(s,a)+cP(s,a)\frac{\sqrt{{\sum}_{b}N(s,b)}}{1+N(s,a)}$ (3)
@@
 $Q(s,a)$ is the mean return of choosing action $a$, $P(s,a)$ is the prior score of action $a$, $N(s,a)$ is the total time that $a$ has been chosen, ${\sum}_{b}N(s,b)$ is the total time that $s$ has been visited.
+Viewing tree search from the bandit-view inspired us to use techniques from the field of bandits to improve the tree search.
+In next section, We use the idea of the one-armed bandit model to design our algorithm and prove the convergence of our algorithm based on the non-stationary bandit model.
 4 Method
+In this section, we introduce the specific design of ReZero.
+Section 4.1 describes the inspiration and specific operations of backward-view reanalyze.
+It performs a reverse search on the trajectory and uses the value estimation of a certain child node to save the corresponding sub-tree search time.
+Section 4.2 analyzes the node selection method introduced in Section 4.1 and ultimately proves the convergence of the method.
+Section 4.3 introduced the overall framework of ReZero, which adopts the buffer reanalyze setting to better integrate with the method described in Section 4.1.
 4.1 Backward-view Reanalyze
-Our algorithm design stems from a simple inspiration: if we could know the true state-value of a child node in advance, we could save the search for it, thus conserving search time.
-As shown on the right side of Figure 1, we directly use the true expected value $\mu_{A}$ to evaluate the quality of Arm A, thereby eliminating the need for the back-propagated value to calculate the score in 3.
-This results in the process occurring within the gray box being omitted.
-Indeed, this situation can be well modeled as an one-armed bandit[18], which also facilitates our subsequent theoretical analysis.
+Our algorithm design stems from a simple inspiration: if we could know the true state-value (e.g., expected long-term return) of a child node in advance, we could save the search for it, thus conserving search time.
+As shown on the right side of Figure 1, we directly use the true expectation $\mu_{A}$ to evaluate the quality of Arm A, thereby eliminating the need for the back-propagated value to calculate the score in Eq.
+3.
+This results in the process occurring within the gray box in Figure 1 being omitted.
+Indeed, this situation can be well modeled as an one-armed bandit (Lattimore & Szepesvári, 2020), which also facilitates our subsequent theoretical analysis.
 Figure 1: The connection between MCTS and bandits.
-Left shows the tree search in a bandit-view.
-When action $A$ is selected, a reward $r_{A}$ will be returned, where $r_{A}=\sum_{t=1}^{3}\gamma^{t-1}r_{t}^{a}+\gamma^{3}v_{a}$.
+Left shows tree search in a bandit-view.
+When the action $A$ is selected, a return $r_{A}$ will be returned, where $r_{A}=\sum_{t=1}^{3}\gamma^{t-1}r_{t}^{a}+\gamma^{3}v_{a}$.
 For the root node, the traversal, evaluation and back-propagation occurring in the sub-tree can be approximated as sampling from an non-stationary distribution.
-Thus the root node can be seen as a non-stationary bandit.
+Thus it can be seen as a non-stationary bandit.
 Right shows the one-armed bandit case.
-Once the reward $\mu_{A}$ is known, we can evaluate arm A using $\mu_{A}$, thereby eliminating the need to rely on subsequent tree search processes.
-Driven by the aforementioned motivation, we aspire to obtain the reward expectation of a child node in advance.
-However, the true expectation is always unknown.
+Once the true value $\mu_{A}$ is known, we can evaluate arm A using $\mu_{A}$, thereby eliminating the need to rely on subsequent tree search processes.
+Driven by the aforementioned motivation, we aspire to obtain the expected return of a child node in advance.
+However, the true value is always unknown.
 Consequently, we resort to using the root value obtained from MCTS as an approximate substitute.
@@
 Fortunately, this is possible during the reanalyze process.
-Algorithm 1 shows the specific design with Python-like code.
-During reanalyze, since the trajectories was already collected in the replay buffer, we can perform tree search in a backward-view, which searches the trajectory in a reverse order.
+During reanalyze, since the trajectories were already collected in the replay buffer, we can perform tree search in a backward-view, which searches the trajectory in a reverse order.
 Figure 2 illustrates a batch containing $n+1$ trajectories, each of length $k+1$, and $S_{l}^{t}$ is the $t$-th state in the $l$-th trajectory.
@@
 $a_{root}=\arg\max_{a}I_{l}^{t}(a)$ (4)
-$I_{t}(a)=\begin{cases}UCB_{score}(S^{t}_{l},a),&a\neq a_{l}^{t}\\ r_{l}^{t}+\gamma m_{l}^{t+1},&a=a_{l}^{t}\end{cases}$ (5)
+${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}I_{l}^{t}(a)}=\begin{cases}UCB_{score}(S^{t}_{l},a),&a\neq a_{l}^{t}\\ r_{l}^{t}+\gamma m_{l}^{t+1},&a=a_{l}^{t}\end{cases}$ (5)
 where $a$ refers to the action associated with a child node, $a_{l}^{t}$ is the action corresponding to $S_{l}^{t+1}$, and $r_{l}^{t}$ signifies the reward predicted by the dynamic model.
-If an action distinct from $a_{l}^{t}$ is selected, the simulation continues its traversal in accordance with the original setting [10].
+If an action distinct from $a_{l}^{t}$ is selected, the simulation continues its traversal with the original setting as in MuZero.
 If action $a_{l}^{t}$ is selected, this simulation is terminated immediately.
 Since the time used to search for node $S^{t+1}_{l}$ is saved, this enhanced search process is faster than the original version.
-Figure 2: An illustration about the backward-view reanalyze in both node and batch view.
+Algorithm 1 shows the specific design with Python-like code.
+[^3] Our enhanced search process is implemented through $reuse\_MCTS()$, where $select\_root\_child()$ performs the action selection method of Equation 5.
+The $traverse()$ and $backpropagate()$ represent the forward search and backward propagation processes in standard MCTS.
+Figure 2: An illustration about the backward-view reanalyze in node and batch view.
 We sample $n+1$ trajectories of length $k+1$ to form a batch and conduct the search in the reverse direction of trajectories.
@@
 Theorem 1
-For a non-stationary bandit that conforms to the assumptions of Equation 2, denote the total number of rounds as $n$, and the times a sub-optimal arm $i$ is selected in $n$ rounds as $T_{i}(n)$, then selecting actions according to Equation 5 can ensure that $\frac{{\mathbb{E}[T_{i}(n)]}}{n}\to 0$ as $n\to\infty$.
-Specifically, if we know the $n$ times sample mean $\hat{\mu}^{*}_{n}$ of the optimal arm in advance, then $\mathbb{E}[T_{i}(n)]$ for all sub-optimal arm $i$ satisfies
+For a non-stationary bandit that conforms to the assumptions of Equation 2, denote the total number of rounds as $n$, the prior score for arm $i$ as $P_{i}$, and the number of times a sub-optimal arm $i$ is selected in $n$ rounds as $T_{i}(n)$, then use a sampled estimation instead of UCB value to evaluate a specific arm(like we do in Equation 5) can ensure that $\frac{{\mathbb{E}[T_{i}(n)]}}{n}\to 0$ as $n\to\infty$.
+Specifically, if we know the $n$ times sample mean $\hat{\mu}^{*}$ of the optimal arm in advance, then $\mathbb{E}[T_{i}(n)]$ for all sub-optimal arm $i$ satisfies
 $\mathbb{E}[T_{i}(n)]\leq 2+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{C^{2}}{(\Delta_{i}-\varepsilon)^{2}}+n\exp{(-\frac{n\varepsilon^{2}}{C^{2}})}$ (6)
-Otherwise, if we have the $n$ times sample mean $\hat{\mu}_{ln}$ of a sub-optimal arm $l$, then for arm $l$,
-$\mathbb{E}[T_{l}(n)]\leq 1+\frac{4C^{4}}{\varepsilon^{4}P_{1}^{2}}+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{C^{2}})}$ (7)
+Otherwise, if we have the $n$ times sample mean $\hat{\mu}_{l}$ of a sub-optimal arm $l$, then for arm $l$,
+$\mathbb{E}[T_{l}(n)]\leq 1+\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2C^{6}}}{\varepsilon^{4}P_{1}^{2}}+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{C^{2}})}$ (7)
 and for other sub-optimal arms,
-$\mathbb{E}[T_{i}(n)]\leq 3+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{C^{2}}{(\Delta_{i}-\varepsilon)^{2}}+\frac{4C^{4}}{\varepsilon^{4}P_{1}^{2}}$ (8)
-where $\Delta_{i}$ is the optimal gap for arm $i$, $\varepsilon$ is a constant in $(0,\Delta_{i})$, and $C$ is the constant in Equation 2.
+$\mathbb{E}[T_{i}(n)]\leq 3+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{C^{2}}{(\Delta_{i}-\varepsilon)^{2}}+\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2C^{6}}}{\varepsilon^{4}P_{1}^{2}}$ (8)
+where $\Delta_{i}$ is the optimal gap for arm $i$, $\varepsilon$ is a constant in $(0,\Delta_{i})$, and $C$ is the constant in Eq. 2.
 We provide the complete proof and draw similar conclusions for AlphaZero in Appendix A.
@@
 This implies that our algorithm may yield visit distribution more concentrated on the optimal arm.
-This is a potential worth exploring in future work, especially in offline scenarios [13] where reanalyze becomes the sole method for policy improvement.
+This is a potential worth exploring in future work, especially in offline scenarios (Schrittwieser et al., 2021) where reanalyze becomes the sole method for policy improvement.
 4.3 The ReZero Framework
 The technique introduced in Section 4.1 will bring a new problem in practice.
-The experiments in the Appendix C.3 demonstrate that batching MCTS allows for parallel model inference and data processing, thereby accelerating the average search speed.
+The experiments in the Appendix D.3 demonstrate that batching MCTS allows for parallel model inference and data processing, thereby accelerating the average search speed.
 However, as shown in the Figure 2, to conduct the backward-view reanalyze, we need to divide the batch into $\frac{1}{k+1}$ of its original size.
@@
 This shift can be interpreted as an alternative method to augment exploration in the collect phase, akin to the noise injection at the root node in MCTS-based algorithms.
-Figure 10 in Appendix C.3 indicates that this modification does not significantly compromise performance during evaluation.
-For the reanalyze process, we introduce the periodical entire-buffer reanalyze mechanism.
+Figure 11 in Appendix D.3 indicates that this modification does not significantly compromise performance during evaluation.
+For the reanalyze process, we introduce the periodical entire-buffer reanalyze.
 As shown in Figure 3, we reanalyze the whole buffer after a fixed number of training iterations.
-For each training iteration, we don’t need to run MCTS to reanalyze the mini-batch, only need to sample the mini-batch and execute the gradient descent.
-Overall, this design offers two significant advantages: 1 The entire buffer reanalyze is akin to the fixed target net mechanism in DQN [28], maintaining a constant policy target for a certain number of training iterations.
+For each iteration, we do not need to run MCTS to reanalyze the mini-batch, only need to sample the mini-batch and execute the gradient descent.
+Overall, this design offers two significant advantages: 1 The entire buffer reanalyze is akin to the fixed target net mechanism in DQN (Mnih et al., 2013), maintaining a constant policy target for a certain number of training iterations.
 Reducing the frequency of policy target updates correspondingly decreases the number of MCTS calls.
@@
 Figure 12 shows both excessively large and excessively small batch sizes can lead to a decrease in search speed.
-We choose the batch size of $2000$ according to the experiment in Appendix C.3.
+We choose the batch size of $2000$ according to the experiment in Appendix D.3.
 Experiments show that our algorithm maintains high sample efficiency and greatly save the running time of the algorithm. Our pipeline also has the following potential improvement directions:
   • When directly using policy for data collection, action selection is no longer bound by tree search.
-Thus, previous vectorized environments like [15] can be seamlessly integrated.
-Besides, this design makes MCTS-based algorithms compatible to existing RL exploration methods like [29].
+Thus, previous vectorized environments like Weng et al.
+(2022) can be seamlessly integrated.
+Besides, this design makes MCTS-based algorithms compatible with existing RL exploration methods like Badia et al.
+(2020).
   • Our method no longer needs to reanalyze the mini-batch for each iteration, thus decoupling the process of reanalyze and training.
@@
 5 Experiment
-Figure 4: Time-efficiency of ReZero-M vs.
-MuZero on six representative Atari games and two board games (Connect4, Gomoku).
-The horizontal axis represents Wall-clock Time (hours), while the vertical axis indicates the Episode Return over 5 episodes.
-ReZero-M achieves higher time-efficiency than the baseline on most Atari games and Gomoku, and achieve similar time-efficiency on some simpler game like Connect4.
-Mean of 5 runs, shaded areas are 95% confidence intervals.
 Efficiency in RL usually refers to two aspects: sample efficiency, the agent’s ability to learn effectively from a limited number of environmental interactions, gauged by the samples needed to reach a successful policy using equal compute resources; time efficiency, which is how swiftly an algorithm learns to make optimal decisions, indicated by the wall-clock time taken to achieve a successful policy.
 Our primary goal is to improve time efficiency without compromising sample efficiency.
-To empirically validate the efficiency of our introduced two boosting techniques for MCTS-based algorithms, we have incorporated the ReZero with two prominent algorithms detailed in [30]: MuZero with a Self-Supervised Learning loss (SSL) and EfficientZero, yielding the enhanced variants ReZero-M and ReZero-E respectively.
-For brevity, when the context is unambiguous, we simply refer to them as ReZero, omitting the symbol representing the baseline algorithm.
-In order to validate the applicability of our method across multiple types of decision-making environments, we opted for six representative Atari environments characterized by classic image-input observation and discrete action spaces, in addition to the strategic board games Connect4 and Gomoku with special state spaces.
-For a baseline comparison, we employed the original implementations of MuZero and EfficientZero as delineated in the LightZero benchmark [30].
-The full implementation details available in Appendix B.
-Besides, we emphasize that to ensure an fair comparison of wall-clock time, all experimental trials were executed on a fixed single worker hardware settings.
-In the next subsections, we sequentially explore and try to answer the following three questions:
+Here we first give a toy example case in Section 5.1 to intuitively demonstrate the acceleration effect of our design, and the corresponding code example is placed in the Appendix D.4 for readers to quickly understand the algorithm design.
+And then, to validate the efficiency of our introduced two boosting techniques for MCTS-based algorithms, we have incorporated the ReZero with two prominent algorithms detailed in Niu et al.
+(2023): MuZero with a Self-Supervised Learning loss (SSL) and EfficientZero, yielding the enhanced variants ReZero-M and ReZero-E respectively.
+For brevity, when the context is unambiguous, we simply refer to them as ReZero, omitting the symbol representing the baseline.
+It is important to highlight that ReZero can be seamlessly integrated with various algorithm variants of MuZero.
+In this context, we have chosen to exemplify this integration using the above two instances.
+In order to validate the applicability of our method across various decision-making environments, we opt for 26 representative Atari environments characterized by classic image-input observation and discrete action spaces, in addition to the strategic board games Connect4 and Gomoku with special state spaces, and two continuous control tasks in DMControl (Tunyasuvunakool et al., 2020).
+For a baseline comparison, we employed the original implementations of MuZero and EfficientZero as delineated in the LightZero benchmark (Niu et al., 2023).
+The full implementation details available in Appendix C.
+Besides, we emphasize that to ensure a fair comparison of wall-clock time, all experimental trials were executed on a fixed single worker hardware settings.
+In Section 5.2, Section 5.3 and Section 5.4, we sequentially explore and try to answer the following three questions:
   • How much can ReZero-M/ReZero-E improve time efficiency compared to MuZero/EfficientZero, while maintaining equivalent levels of high sample efficiency?
@@
   • How much search budgets can be saved in practice by the Backward-view Reanalyze technique?
-Atari Board Games
-avg. wall time (h) to 100k env. steps $\downarrow$ Pong Breakout MsPacman Qbert Seaquest UpNDown Connet4 Gomoku
-ReZero-M (ours) $\mathbf{1}\scriptstyle{\pm 0.6}$ $\mathbf{7.7}\scriptstyle{\pm 0.3}$ $\mathbf{3.8}\scriptstyle{\pm 0.2}$ $\mathbf{5.1\scriptstyle{\pm 0.3}}$ $\mathbf{6.1}\scriptstyle{\pm 0.5}$ $\mathbf{6.9}\scriptstyle{\pm 0.5}$ $\mathbf{10}\scriptstyle{\pm 0.8}$ $\mathbf{9\scriptstyle{\pm 0.8}}$
-MuZero [10] $4.1\scriptstyle{\pm 0.2}$ $12.4\scriptstyle{\pm 0.8}$ $12.0\scriptstyle{\pm 0.9}$ $10.2\scriptstyle{\pm 0.5}$ $11.75\scriptstyle{\pm 0.9}$ $12.4\scriptstyle{\pm 0.9}$ $11\scriptstyle{\pm 0.9}$ $23\scriptstyle{\pm 2}$
+5.1 Toy Case
+Figure 4: Acceleration effect on the toy case.
+Left is a simple maze environment where the agent starts at point A and receives a reward of size $1$ upon reaching the end point G.
+Middle shows the search time corresponding to each position when set as the root node.
+Meanwhile, the root node values obtained during the search are preserved.
+Right shows the corresponding search time when these root node values are used to assist the search.
+The comparison shows that the search duration is generally reduced.
+For specific experimental settings and code, please refer to the Appendix D.4.
+Intuitively, it is evident that our proposed method can achieve a speed gain because we eliminate the search of a certain subtree, especially when this subtree corresponds to the optimal action (which often implies that the subtree has a larger number of nodes).
+We conduct an experiment on a toy example case and include the experimental code in the Appendix D.4.
+This helps to visually illustrate the speed gain achieved by skipping subtree search and allows readers to quickly understand the algorithm design through simple code.
+As shown in Figure 4 (Left), we implement a simple $7\times 7$ maze environment where the agent starts at point A and receives a reward of $1$ upon reaching point G.
+We perform an MCTS with each position in the maze as the root node and recorded the search time in Figure 4 (Middle).
+It can be seen that regions farther from the end point require more time to search (this is related to our simulation settings, see the appendix for details).
+For comparision, we also performed searches with each position as the root node, but during the search, we used the root node values obtained from the search in Figure 4 (Middle) to evaluate specific actions.
+The experimental time was recorded in Figure 4 (Right).
+It can be seen that after eliminating the search of specific subtrees, the search time was generally reduced.
+This simple result validates the rationality of our algorithm design.
+In the next section, we will specifically validate the time efficiency of the ReZero framework in diverse decision-making tasks in Section 5.2.
+Figure 5: Time-efficiency of ReZero-M vs.
+MuZero on four representative Atari games, two continuous control tasks of DMControl (ball_in_cup-catch, walker-stand), and two board games (Connect4, Gomoku).
+The horizontal axis represents Wall-clock Time (hours), while the vertical axis indicates the Episode Return over 5 evaluation episodes.
+ReZero-M demonstrates superior time-efficiency compared to the baseline across a diverse set of games, encompassing both image and state observations, discrete and continuous actions, and scenarios involving sparse rewards.
+These figures compute mean of 5 runs, and shaded areas are 95% confidence intervals.
+5.2 Time Efficiency
+Setup: We aim to evaluate the performance of ReZero against classical MCTS-based algorithms MuZero and EfficientZero, focusing on the wall-clock time reduction required to achieve the comparable performance level.
+In order to facilitate a fair comparison in terms of wall-time, we not only utilize identical computational resources but also maintain consistent hyper-parameter settings across the algorithms (unless specified cases).
+Key parameters are aligned with those from original papers.
+After each data collection phase, the model is trained for multiple iterations according to the replay ratio, which is denoted as one training epoch.
+Specifically, we set the replay ratio (the ratio between environment steps and training steps) (D’Oro et al., 2022) to 0.25, and the reanalyze ratio (the ratio between targets computed from the environment and by reanalysing existing data) (Schrittwieser et al., 2021) is set to 1.
+For detailed hyper-paramater configurations, please refer to the Appendix C.
+Results: Our experiments shown in Figure 5 illustrates the training curves and performance comparisons in terms of wall-clock time between ReZero-M and the MuZero across various decision-making environments.
+Note that on continuous control tasks we use the sampled version of ReZero-M and MuZero Hubert et al.
+(2021).
+The data clearly indicates that ReZero-M achieves a significant improvement in time efficiency, attaining a near-optimal policy in significantly less time for these eight diverse decision-making tasks.
+To provide an alternative perspective on the time cost of training on the same number of environment steps, we also provide an in-depth comparison of the wall-clock training time up to 100k environment steps for ReZero-M against MuZero in Table 1.
+For a more comprehensive understanding of our findings, we offer the complete table of the 26 Atari games usually used for sample-efficient RL in Appendix D.1.
+It reveals that, on most games, ReZero require between 2-4 times less wall-clock time per 100k steps compared to the baselines, while maintaining comparable or even superior performance in terms of episode return.
+Additional results about another algorithm comparison instance ReZero-E and EfficientZero are detailed in Appendix D.2, specifically in Figure 8 and 9.
+This setting also supports the time efficiency of our proposed ReZero framework.
+Atari DMControl Board Games
+avg. wall time (h) to 100k env. steps $\downarrow$ Pong Breakout MsPacman Seaquest ball_in_cup-catch walker-stand Connet4 Gomoku
+ReZero-M (ours) $\mathbf{1.0}\scriptstyle{\pm 0.1}$ $\mathbf{3.0}\scriptstyle{\pm 0.8}$ $\mathbf{1.4}\scriptstyle{\pm 0.2}$ $\mathbf{1.9\scriptstyle{\pm 0.4}}$ $\mathbf{2.1}\scriptstyle{\pm 0.2}$ $\mathbf{4.3}\scriptstyle{\pm 0.3}$ $\mathbf{5.5}\scriptstyle{\pm 0.6}$ $\mathbf{4.5\scriptstyle{\pm 0.5}}$
+MuZero (Schrittwieser et al., 2019) $4.0\scriptstyle{\pm 0.5}$ $4.9\scriptstyle{\pm 1.8}$ $6.9\scriptstyle{\pm 0.3}$ $10.1\scriptstyle{\pm 0.5}$ $5.6\scriptstyle{\pm 0.4}$ $9.5\scriptstyle{\pm 0.6}$ $9.1\scriptstyle{\pm 0.8}$ $15.3\scriptstyle{\pm 1.5}$
 Table 1: Average wall-time of ReZero-M vs.
 MuZero on various tasks.
-(left) Six Atari games, (right) two board games.
+(left) Four Atari games, (middle) two control tasks, (right) two board games.
 The time represents the average total wall-time to 100k environment steps for each algorithm.
 Mean and standard deviation over 5 runs.
-5.1 Time Efficiency
-Setup In this section, we aim to evaluate the performance of ReZero against classical MCTS-based algorithms MuZero and EfficientZero, focusing on the wall-clock time reduction required to achieve the comparable performance level.
-In order to facilitate a fair comparison in terms of wall-time, for each run, we not only utilize identical computational resources but also maintain consistent hyper-parameter settings across the algorithms (unless otherwise specified cases).
-Key parameters are aligned with those from original papers.
-Specifically, we set the replay ratio (the ratio between environment steps and training steps) [31, 32] to 0.25, and the reanalyze ratio (the ratio between targets computed from interactions with the environment and by reanalysing existing data) [13] is set to 1.
-For detailed hyper-paramater configurations, please refer to the Appendix B.
-Results Figure 4 illustrates the performance comparison in terms of wall-clock time between ReZero-M and the original MuZero across a variety of RL environments.
-It is apparent that ReZero-M achieves a significant improvement in time efficiency, reaching a near-optimal policy in substantially less time for the games Pong, Breakout and Seaquest, and obtained a better policy than MuZero on the remaining tasks.
-Additional results about ReZero-E and EfficientZero are detailed in Appendix C.2, specifically in Figure 7 and 8.
-For a more comprehensive understanding of our findings, we provide an in-depth comparison of the wall-clock training time up to 100k environment steps for ReZero-M against MuZero in Table 1.
-The data reveals that, on most games, ReZero require between two to four times less wall-clock time per 100k steps compared to the baseline methods.
-Figure 5: The effect of Reanalyze Frequency in ReZero-M on the MsPacman game.
+Figure 6: The ablation experiment of Reanalyze Frequency in ReZero-M on the Atari MsPacman game.
 The proper reanalyze frequency can improve time and sample efficiency while obtaining the comparable return with MuZero (reanalyze ratio=1).
@@
 Firstly, we will analyze how the entire-buffer reanalyze technique manages to simplify the original iterative mini-batch reanalyze scheme to the periodical updated version while maintaining high sample efficiency.
-Here we adjust the periodical reanalyze frequency(reanalyze the buffer 1 time every $n$ training iterations) in ReZero to a lower or higher value on the MsPacman environment, specifically reanalyze frequency=$\{2,1,1/3,0\}$, where $1$ denotes the setting $n=1000$.
+Here, we adjust the periodic reanalyze frequency—which determines how often the buffer is reanalyzed during a training epoch—in ReZero for the MsPacman environment.
+Specifically, we set reanalyze frequency to $\{0,\frac{1}{3},1,2\}$.
 The original MuZero variants with the reanalyze ratio of 1 is also included in this ablation experiment as a baseline.
@@
 MuZero $1.08\pm 0.09$ $13284$ $256$ $455$
-Table 2: Comparisons about detailed time indicators between MuZero and ReZero-M inside the tree search.
+Table 2: Comparisons about the detailed time cost indicators between MuZero and ReZero-M inside the tree search.
 To further validate and understand the advantages and significance of the backward-view reanalyze technique proposed in ReZero, we meticulously document a suite of statistical indicators of the tree search process in Table 2.
@@
 Consequently, this advanced technique in leveraging subsequent time step data contributes to save the tree search time in various MCTS-based algorithms, further leading to overall wall-clock time gains.
+The complexity and implementation of the tree search process directly influences the efficiency gains achieved through the backward-view reanalyze technique.
+As the tree search becomes more intricate and sophisticated, e.g.
+Sampled MuZero (Hubert et al., 2021), the time savings realized through this method are correspondingly amplified.
 Additionally, Theorem 1 demonstrates that the backward-view reanalyze can reduce the regret upper bound, which indicates a better search result.
-Figure 9 in Appendix C.3 shows a comparison of sample efficiency between using backward-view reanalyze and origin reanalyze process in MsPacman.
+Besides, Figure 10 in Appendix D.3 shows a comparison of sample efficiency between using backward-view reanalyze and origin reanalyze process in MsPacman.
 The experimental results reveal that our method not only enhances the speed of individual searches but also improves sample efficiency.
@@
 In this paper, we have delved into the efficiency and scalability of MCTS-based algorithms.
-Unlike most existing works, we incorporates information reuse and periodic reanalyze techniques to reduces wall-clock time costs while preserving sample efficiency.
+Unlike most existing works, we incorporate information reuse and periodic reanalyze techniques to reduces wall-clock time costs while preserving sample efficiency.
 Our theoretical analysis and experimental results confirm that ReZero efficiently reduces the time cost and maintains or even improves performance across different decision-making domains.
 However, our current experiments are mainly conducted on the single worker setting, there exists considerable optimization scope to apply our approach into distributed RL training, and our design harbors the potential of better parallel acceleration and more stable convergence in large-scale training tasks.
-Also, the combination between ReZero and frameworks akin to AlphaZero, or its integration with some recent offline datasets such as RT-X [33], constitutes a fertile avenue for future research.
-These explorations could broadening the application horizons of MCTS-based algorithms.
+Also, the combination between ReZero and frameworks akin to AlphaZero, or its integration with some recent offline datasets such as RT-X (Padalkar et al., 2023), constitutes a fertile avenue for future research.
+These explorations could broaden the application horizons of MCTS-based algorithms.
+Additionally, since in offline training scenarios, reanalyze becomes the only means of policy improvement, this makes the acceleration of the reanalyze phase in ReZero even more critical.
+Moreover, the potential improvement in search results by ReZero may further improve the training result, rather than merely accelerating the training.
+Therefore, combining ReZero with MuZero Unplugged (Schrittwieser et al., 2021) is a direction worth exploring for building foundation models for decision-making.
 Appendix A Proof Materials
 In this section, we will present the complete supplementary proofs.
-Lemma A.1.
-[…]
-$\kappa=\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{t}+\sqrt{\frac{a}{t^{2}}}\geq\varepsilon\}$ (9)
-, then it holds that $E[\kappa]\leq 1+\frac{2\sqrt{a}}{\varepsilon}+\frac{C^{2}}{\varepsilon^{2}}$.
+Lemma A: Let $w_{t}$ be a random variable that satisfies the concentration condition in Equation 2 with zero expectation, $\varepsilon>0$, $a>0$ and
+$\kappa=\sum_{t=1}^{n}\mathbb{I}\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}w_{t}}+\sqrt{\frac{a}{t^{2}}}\geq\varepsilon\}$ (9)
+then it holds that $E[\kappa]\leq 1+\frac{2\sqrt{a}}{\varepsilon}+\frac{C^{2}}{\varepsilon^{2}}$.
 Proof.
 Take $u$ as $\frac{2\sqrt{a}}{\varepsilon}$, then
-$E[\kappa]$ $\leq u+\sum_{t=\lceil u\rceil}^{n}\mathbb{P}(\hat{\mu}_{t}+\sqrt{\frac{a}{t^{2}}}\geq\varepsilon)$ (10)
+$E[\kappa]$ $\leq u+\sum_{t=\lceil u\rceil}^{n}\mathbb{P}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}w_{t}}+\sqrt{\frac{a}{t^{2}}}\geq\varepsilon)$ (10)
 $\leq u+\sum_{t=\lceil u\rceil}^{n}\exp(-\frac{t(\varepsilon-\sqrt{\frac{a}{t^{2}}})^{2}}{C^{2}})$ (11)
@@
 A slight adjustment to this proof yields the conclusion of Theorem 1.
-Denote $T_{i}(k)$ as the times that arm $i$ has been chosen until time $k$, $\hat{\mu}_{is}$ as the average of the first $s$ samples of arm $i$, which satisfies the concentration assumption in Equation 2, and $\hat{\mu}_{i}(k)=\hat{\mu}_{iT_{i}(k)}$.
+Denote $T_{i}(k)$ as the number of times that arm $i$ has been chosen until time $k$, $A_{t}$ as the arm selected at time $t$, $\hat{\mu}_{is}$ as the average of the first $s$ samples of arm $i$, $\mu_{i}$ as the limit of $\mathbb{E}[\hat{\mu}_{is}]$, which satisfies the concentration assumption in Equation 2, and $\hat{\mu}_{i}(k)=\hat{\mu}_{iT_{i}(k)}$.
+Without loss of generality, we assume that arm $1$ is the optimal arm.
 Then we have:
@@
 $+\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}$ (17)
-for 16, we have
-$E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]$ (18)
+for Equation 16, we have
+$\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]$ (18)
 $\leq 1+\sum_{t=2}^{n}\sum_{s=0}^{t-1}\mathbb{P}(\hat{\mu}_{1s}+P_{1}\frac{\sqrt{t-1}}{1+s}\leq\mu_{1}-\varepsilon)$ (19)
@@
 $\leq 1+\sum_{t=2}^{n}\exp(-\frac{1}{C^{2}}\varepsilon P_{1}\sqrt{t-1})\frac{C^{2}}{\varepsilon^{2}}$ (23)
-$\leq 1+\frac{4C^{4}}{\varepsilon^{4}P_{1}^{2}}$ (24)
-Notes: In 19, since we assume $P_{1}\geq\mu_{1}$, the probability of the term $s=0$ would be 00.
+$\leq 1+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{2C^{6}}{\varepsilon^{4}P_{1}^{2}}}$ (24)
+Notes: In Equation 19, since we assume $P_{1}\geq\mu_{1}$, the probability of the term $s=0$ would be 00.
 Thus, we can discard it.
@@
 This only changes the constant term and won’t affect the growth rate of regret.
-From 21 to 22, we just need to expand the quadratic term and do some simple inequality scaling.
-From 22 to 23, we need to notice that $\sum_{s=1}^{t-1}\exp(-\frac{s\varepsilon^{2}}{C^{2}})$ is a geometric sequence and scale it to $\frac{C^{2}}{\varepsilon^{2}}$.
-From 23 to 24, we use the inequality $\sum_{t=2}^{n}\frac{1}{e^{a\sqrt{t-1}}}\leq\int_{1}^{\infty}\frac{1}{e^{a\sqrt{t-1}}}$.
-for 17, we have
-$E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (25)
-$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\sqrt{\frac{n-1}{(1+T_{i}(t-1))^{2}}}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (26)
-$\leq 1+E[\sum_{s=1}^{n-1}\mathbb{I}\{\hat{\mu}_{is}+P_{i}\sqrt{\frac{n-1}{(1+s)^{2}}}\geq\mu_{1}-\varepsilon\}]$ (27)
-$\leq 1+E[\sum_{s=1}^{n}\mathbb{I}\{\hat{\mu}_{is}-\mu_{i}+P_{i}\sqrt{\frac{n-1}{s^{2}}}\geq\Delta_{i}-\varepsilon\}]$ (28)
-with Lemma A.1, we can have
+From Equation 21 to Equation 22, we just need to expand the quadratic term and do some simple inequality scaling.
+From Equation 22 to Equation 23, we need to notice that $\sum_{s=1}^{t-1}\exp(-\frac{s\varepsilon^{2}}{C^{2}})$ is a geometric sequence and scale it to $\frac{C^{2}}{\varepsilon^{2}}$.
+From Equation 23 to Equation 24, we use the inequality $\sum_{t=2}^{n}\frac{1}{e^{a\sqrt{t-1}}}\leq\int_{1}^{\infty}\frac{1}{e^{a\sqrt{t-1}}}$.
+And for Equation 17, we have
+$\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (25)
+$\leq\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\sqrt{\frac{n-1}{(1+T_{i}(t-1))^{2}}}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (26)
+$\leq 1+\mathbb{E}[\sum_{s=1}^{n-1}\mathbb{I}\{\hat{\mu}_{is}+P_{i}\sqrt{\frac{n-1}{(1+s)^{2}}}\geq\mu_{1}-\varepsilon\}]$ (27)
+$\leq 1+\mathbb{E}[\sum_{s=1}^{n}\mathbb{I}\{\hat{\mu}_{is}-\mu_{i}+P_{i}\sqrt{\frac{n-1}{s^{2}}}\geq\Delta_{i}-\varepsilon\}]$ (28)
+with Lemma A, we can have
 282+$\leq 2+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{C^{2}}{(\Delta_{i}-\varepsilon)^{2}}$ (29)
-so
-$E[T_{i}(n)]\leq 3+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{C^{2}}{(\Delta_{i}-\varepsilon)^{2}}+\frac{4C^{4}}{\varepsilon^{4}P_{1}^{2}}$ (30)
+so we have
+$\mathbb{E}[T_{i}(n)]\leq 3+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{C^{2}}{(\Delta_{i}-\varepsilon)^{2}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{2C^{6}}{\varepsilon^{4}P_{1}^{2}}}$ (30)
 The proof of theorem 1 can be obtained by making slight modifications.
-In case we have $\hat{\mu}^{*}_{n}=\frac{1}{n}\sum_{t=1}^{n}\hat{\mu}_{t}$ where $\hat{\mu}_{t}$ is sampled from the optimal arm in advance,
-$E[T_{i}(n)]=E[\sum_{t=1}^{n}\mathbb{I}\{A_{t}=i\}]$ (31)
-$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1n}\leq\mu_{1}-\varepsilon\}]$ (32)
-$+E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (33)
-$\leq n\exp{(-\frac{n\varepsilon^{2}}{C^{2}})}+E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (34)
+In case we have drawn $n$ samples from the same non-stationary distribution as arm $1$, and the average of these first $n$ samples is $\hat{\mu}_{1}$,
+$\mathbb{E}[T_{i}(n)]=\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}\{A_{t}=i\}]$ (31)
+$\leq\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{\mu}_{1}}\leq\mu_{1}-\varepsilon\}]$ (32)
+$+\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (33)
+$\leq n\exp{(-\frac{n\varepsilon^{2}}{C^{2}})}+\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{i}(t-1)+P_{i}\frac{\sqrt{t-1}}{1+T_{i}(t-1)}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (34)
 $\leq 2+\frac{2P_{i}\sqrt{n-1}}{\Delta_{i}-\varepsilon}+\frac{C^{2}}{(\Delta_{i}-\varepsilon)^{2}}+n\exp{(-\frac{n\varepsilon^{2}}{C^{2}})}$ (35)
-In case we have $\hat{\mu}_{ln}=\frac{1}{n}\sum_{t=1}^{n}\hat{\mu}_{t}$ where $\hat{\mu}_{t}$ is sampled from the same reward distribution of arm $l$,
-$E[T_{l}(n)]=E[\sum_{t=1}^{n}\mathbb{I}\{A_{t}=l\}]$ (36)
-$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]$ (37)
-$+E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{ln}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}=i\}]$ (38)
-$\leq E[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{C^{2}})}$ (39)
-$\leq 1+\frac{4C^{4}}{\varepsilon^{4}P_{1}^{2}}+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{C^{2}})}$ (40)
+In case we have drawn $n$ samples from the same non-stationary distribution as arm $l$, and the average of these first $n$ samples is $\hat{\mu}_{l}$,
+$\mathbb{E}[T_{l}(n)]=\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}\{A_{t}=l\}]$ (36)
+$\leq\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]$ (37)
+$+\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\{\hat{\mu}_{l}}\geq\mu_{1}-\varepsilon\ \text{and}\ A_{t}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}l}\}]$ (38)
+$\leq\mathbb{E}[\sum_{t=1}^{n}\mathbb{I}\{\hat{\mu}_{1}(t-1)+P_{1}\frac{\sqrt{t-1}}{1+T_{1}(t-1)}\leq\mu_{1}-\varepsilon\}]+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{C^{2}})}$ (39)
+$\leq 1+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{2C^{6}}{\varepsilon^{4}P_{1}^{2}}}+n\exp{(-\frac{n(\Delta_{l}-\varepsilon)^{2}}{C^{2}})}$ (40)
 and the bound of $E[T_{i}(n)]$ for $i\neq l$ keeps unchanged. ∎
-Appendix B Implementation Details
-B.1 Environments
+Appendix B MuZero
+During the inference phase, the representation model transforms a sequence of the last $l$ observations $o_{t-l:t}$ into a corresponding latent state representation $s_{t}$.
+The dynamics model processes this latent state alongside an action $a_{t}$, yielding the subsequent latent state $s_{t+1}$ and an estimated reward $r_{t}$.
+Finally, the prediction model accepts a latent state and produces both the predicted policy $p_{t}$ and the state’s value estimate $v_{t}$.
+These outputs are instrumental in guiding the agent’s action selection process throughout its MCTS.
+Lastly the agent selects or samples the best action $a_{t}$ following the searched visit count distribution.
+During the training phase, given a training sequence $\{o_{t-l:t+K},a_{t+1:t+K},u_{t+1:t+K},\pi_{t+1:t+K},z_{t+1:t+K}\}$ at time $t$ sampled from the replay buffer, where $u_{t+k}$ denotes the actual reward obtained from the environment, $\pi_{t+k}$ represents the target policy obtained through MCTS during the agent-environment interaction, and $z_{t+k}$ is the value target computed using n-step bootstrapping Hessel et al.
+(2018).
+The representation model initially converts the sequence of observations $o_{t-l:t}$ into the latent state $s_{t}^{0}$.
+Subsequently, the dynamic model executes $K$ latent space rollouts based on the sequence of actions $a_{t+1:t+K}$.
+The latent state derived after the $k$-th rollout is denoted as $s_{t}^{k}$, with the corresponding predicted reward indicated as $r_{t}^{k}$.
+Upon receiving $s_{t}^{k}$, the prediction model generates a predicted policy $p_{t}^{k}$ and a estimated value $v_{t}^{k}$.
+The final training loss encompasses three components: the policy loss ($l_{p}$), the value loss ($l_{v}$), and the reward loss ($l_{r}$):
+$L_{\text{MuZero}}=$ $\sum_{k=0}^{K}l_{p}(\pi_{t+k},p_{t}^{k})+\sum_{k=0}^{K}l_{v}(z_{t+k},v_{t}^{k})+\sum_{k=1}^{K}l_{r}(u_{t+k},r_{t}^{k})$ (41)
+MuZero Reanalyze, as introduced in Schrittwieser et al.
+(2021), is an advanced iteration of the original MuZero algorithm.
+This variant enhances the model’s accuracy by conducting a fresh Monte Carlo Tree Search on sampled states with the most recent version of the model, subsequently utilizing the refined policy from this search to update the policy targets.
+Such reanalysis yields targets of superior quality compared to those obtained during the initial data collection phase.
+The Schrittwieser et al.
+(2021) expands upon this approach, formalizing it as a standalone method for policy refinement.
+This innovation opens avenues for its application in offline settings, where interactions with the environment are not possible.
+Appendix C Implementation Details
+C.1 Environments
 In this section, we first introduce various types of reinforcement learning environments evaluated in the main paper and their respective characteristics, including different observation/action/reward space and transition functions.
@@
 With their high-dimensional visual input and discrete action space features, Atari environments are widely used to evaluate the capability of reinforcement learning algorithms in handling visual inputs and discrete control problems.
+DMControl: This continuous control suite comprises 39 continuous control tasks.
+Our focus here is to validate the effectiveness of ReZero in the continuous action space.
+Consequently, we have utilized two representative tasks (ball_in_cup-catch and walker-stand) for illustrative purposes.
+A comprehensive benchmark for this domain will be included in future versions.
 Board Games: This types of environment includes Connect4, Gomoku, where uniquely marked boards and explicitly defined rules for placement, movement, positioning, and attacking are employed to achieve the game’s ultimate objective.
@@
 In practice, algorithms utilize action mask to indicate reasonable actions.
-B.2 Algorithm Implementation Details
-Our algorithm’s implementation is based on the open-source code of LightZero [30].
+C.2 Algorithm Implementation Details
+Our algorithm’s implementation is based on the open-source code of LightZero (Niu et al., 2023).
 Given that our proposed theoretical improvements are applicable to any MCTS-based RL method, we have chosen MuZero and EfficientZero as case studies to investigate the practical improvements in time efficiency achieved by integrating the ReZero boosting techniques: just-in-time reanalyze and speedy reanalyze (temporal information reuse).
 To ensure an equitable comparison of wall-clock time, all experimental trials were executed on a fixed single worker hardware setting consisting of a single NVIDIA A100 GPU with 30 CPU cores and 120 GiB memory.
-Besides, we emphasize that to ensure a fair comparison of time efficiency and sample efficiency, the model architecture and hyper-parameters used in the experiments of Section 5 are essentially consistent with the settings in LightZero [30].
+Besides, we emphasize that to ensure a fair comparison of time efficiency and sample efficiency, the model architecture and hyper-parameters used in the experiments of Section 5 are essentially consistent with the settings in LightZero.
 For specific hyper-parameters of ReZero-M and MuZero on Atari, please refer to the Table 3.
+The main different hyper-parameters in the DMControl task are set out in Tables 4.
 The main different hyper-parameters for the ReZero-M algorithm in the Connect4 and Gomoku environment are set out in Tables 5.
-In addition to employing an LSTM network with a hidden state dimension of 512 to predict the value prefix [21], all hyperparameters of ReZero-E are essentially identical to those of ReZero-M in Table 3.
+In addition to employing an LSTM network with a hidden state dimension of 512 to predict the value prefix (Ye et al., 2021), all hyperparameters of ReZero-E are essentially identical to those of ReZero-M in Table 3.
 Wall-time statistics Note that all our current tests are conducted in the single-worker case. Therefore, the wall-time reported in Table 1 and Table 7 for reaching 100k env steps includes:
-  • collect time: The total time spent by an agent interacting with the environment to gather experience data.[15] can be integrated to speed up.
-Besides, this design also makes MCTS-based algorithms compatible to existing RL exploration methods like [34].
-  • reanalyze time: The time used to reanalyze collected data with the current policy or value function for more accurate learning targets [13].
+  • collect time: The total time spent by an agent interacting with the environment to gather experience data.Weng et al.
+(2022) can be integrated to speed up.
+Besides, this design also makes MCTS-based algorithms compatible to existing RL exploration methods like Burda et al.
+(2018).
+  • reanalyze time: The time used to reanalyze collected data with the current policy or value function for more accurate learning targets (Schrittwieser et al., 2021).
   • train time: The duration for performing updates to the agent’s policy, value functions and model based on collected data.
@@
 This approach is versatile, enabling seamless integration with single-worker serial execution environments as well as multi-worker asynchronous frameworks.
-The exploration of ReZero’s extensions and its evaluations in a multi-worker [16] paradigm are earmarked for future investigation.
+The exploration of ReZero’s extensions and its evaluations in a multi-worker (Mei et al., 2023) paradigm are earmarked for future investigation.
 Board games settings Given that our primary objective is to test the proposed techniques for improvements in time efficiency, we consider a simplified version of single-player mode in all the board games.
@@
 Hyperparameter Value
-Replay Ratio [32] 0.25
-Reanalyze ratio [13] 1
+Replay Ratio (Schwarzer et al., 2023) 0.25
+Reanalyze frequency 1
 Num of frames stacked 4
 Num of frames skip 4
-Reward clipping [28] True
-Optimizer type Adam [35]
+Reward clipping (Mnih et al., 2013) True
+Optimizer type Adam
 Learning rate $3\times 10^{-3}$
@@
 Weight of policy entropy loss 0
-Weight of SSL (self-supervised learning) loss [21] 2
+Weight of SSL (self-supervised learning) loss (Ye et al., 2021) 2
 Batch size 256
@@
 Discrete action encoding type One Hot
-Normalization type Batch Normalization [36]
-Priority exponent coefficient [37] 0.6
+Normalization type Layer Normalization
+Priority exponent coefficient (Schaul et al., 2015) 0.6
 Priority correction coefficient 0.4
@@
 Categorical distribution in value and reward modeling True
-The scale of supports used in categorical distribution [38] 300
+The scale of supports used in categorical distribution (Pohlen et al., 2018) 300
 Table 3: Key hyperparameters of ReZero-M on Atari environments.
 Hyperparameter Value
-Replay Ratio [32] 0.25
-Reanalyze ratio [13] 1
+Replay ratio (Schwarzer et al., 2023) 0.25
+Reanalyze frequency 1
+Batch size 64
+Num of frames stacked 1
+Num of frames skip 2
+Discount factor 0.997
+Length of game segment 8
+Use augmentation False
+Number of simulations in MCTS (sim) 50
+Number of sampled actions (Hubert et al., 2021) 20
+Table 4: Key hyperparameters of ReZero-M on two DMControl tasks (ball_in_cup-catch and walker-stand).
+More experiments about this hyper-parameter will be explored in the future version.
+Other unmentioned parameters are the same as that in Atari settings.
+Hyperparameter Value
+Replay ratio (Schwarzer et al., 2023) 0.25
+Reanalyze frequency 1
 Board size 6x7; 6x6
@@
 Table 5: Key hyperparameters of ReZero-M on Connect4 and Gomoku environments. If the parameter settings of these two environments are different, they are separated by a semicolon.
-Appendix C Additional Experiments
-C.1 Sample-efficiency of ReZero-M
-Except for the results reported in main paper for time efficiency, we also shows the detailed sample efficiency results in this section.
+Appendix D Additional Experiments
+D.1 ReZero-M
+In this section, we provide additional experimental results for ReZero-M. As a supplement to Table 1, Table 6 presents the complete experimental results on the 26 Atari environments.
+Env. Name ReZero-M MuZero
+Alien $\mathbf{1.6}\scriptstyle{\pm 0.2}$ $8.6\scriptstyle{\pm 0.4}$
+Amidar $\mathbf{1.5}\scriptstyle{\pm 0.2}$ $8.1\scriptstyle{\pm 0.3}$
+Assault $\mathbf{1.5}\scriptstyle{\pm 0.1}$ $7.5\scriptstyle{\pm 0.1}$
+Asterix $\mathbf{1.3}\scriptstyle{\pm 0.1}$ $7.2\scriptstyle{\pm 0.2}$
+BankHeist $\mathbf{2.9}\scriptstyle{\pm 0.3}$ $8.9\scriptstyle{\pm 0.6}$
+BattleZone $\mathbf{2.2}\scriptstyle{\pm 0.3}$ $9.6\scriptstyle{\pm 0.6}$
+ChopperCommand $\mathbf{3.4}\scriptstyle{\pm 0.4}$ $9.0\scriptstyle{\pm 0.7}$
+CrazyClimber $\mathbf{2.7}\scriptstyle{\pm 0.1}$ $9.1\scriptstyle{\pm 0.4}$
+DemonAttack $\mathbf{1.1}\scriptstyle{\pm 0.1}$ $6.8\scriptstyle{\pm 0.8}$
+Freeway $\mathbf{1.0}\scriptstyle{\pm 0.0}$ $6.1\scriptstyle{\pm 0.2}$
+Frostbite $\mathbf{2.2}\scriptstyle{\pm 0.4}$ $10.9\scriptstyle{\pm 0.8}$
+Gopher $\mathbf{3.2}\scriptstyle{\pm 0.6}$ $8.1\scriptstyle{\pm 0.8}$
+Hero $\mathbf{2.5}\scriptstyle{\pm 0.4}$ $9.9\scriptstyle{\pm 0.6}$
+Jamesbond $\mathbf{2.2}\scriptstyle{\pm 0.3}$ $9.1\scriptstyle{\pm 0.5}$
+Kangaroo $\mathbf{2.0}\scriptstyle{\pm 0.2}$ $8.6\scriptstyle{\pm 0.8}$
+Krull $\mathbf{1.8}\scriptstyle{\pm 0.1}$ $7.7\scriptstyle{\pm 0.3}$
+KungFuMaster $\mathbf{1.3}\scriptstyle{\pm 0.1}$ $7.6\scriptstyle{\pm 0.7}$
+PrivateEye $\mathbf{1.0}\scriptstyle{\pm 0.1}$ $5.8\scriptstyle{\pm 0.5}$
+RoadRunner $\mathbf{1.5}\scriptstyle{\pm 0.2}$ $9.0\scriptstyle{\pm 0.3}$
+UpNDown $\mathbf{1.4}\scriptstyle{\pm 0.1}$ $7.2\scriptstyle{\pm 0.4}$
+Pong $\mathbf{1.0}\scriptstyle{\pm 0.1}$ $4.0\scriptstyle{\pm 0.5}$
+MsPacman $\mathbf{1.4}\scriptstyle{\pm 0.2}$ $6.9\scriptstyle{\pm 0.3}$
+Qbert $\mathbf{1.3}\scriptstyle{\pm 0.1}$ $7.0\scriptstyle{\pm 0.3}$
+Seaquest $\mathbf{1.9}\scriptstyle{\pm 0.4}$ $10.1\scriptstyle{\pm 0.5}$
+Boxing $\mathbf{1.1}\scriptstyle{\pm 0.0}$ $6.6\scriptstyle{\pm 0.1}$
+Breakout $\mathbf{3.0}\scriptstyle{\pm 0.8}$ $4.9\scriptstyle{\pm 1.8}$
+Table 6: Average wall-time(hours) of ReZero-M vs.
+MuZero on 26 Atari game environments.
+The time represents the average total wall-clock time to 100k environment steps.
+Mean and standard deviation over 5 runs.
 Figure 7 displays the performance over environment interaction steps of the ReZero-M algorithm compared with the original MuZero algorithm across six representative Atari environments and two board games.
@@
 Mean of 5 runs; shaded areas are 95% confidence intervals.
-C.2 ReZero-E
+D.2 ReZero-E
 The enhancements of ReZero we have proposed are universally applicable to any MCTS-based reinforcement learning approach theoretically.
@@
 ReZero-E achieves higher time-efficiency than the baseline method.
-Mean of 3 runs; shaded areas are 95% confidence intervals.
+Mean of 5 runs; shaded areas are 95% confidence intervals.
 Figure 9: Sample-efficiency of ReZero-E vs.
@@
 ReZero-E achieves similar sample-efficiency than the baseline method.
-Mean of 3 runs; shaded areas are 95% confidence intervals.
+Mean of 5 runs; shaded areas are 95% confidence intervals.
 avg. wall time (h) to 100k env. steps $\downarrow$ Pong MsPacman Seaquest UpNDown
 ReZero-E (ours) $\mathbf{2.3}\scriptstyle{\pm 1.4}$ $\mathbf{3}\scriptstyle{\pm 0.3}$ $\mathbf{3.1}\scriptstyle{\pm 0.1}$ $\mathbf{3.6}\scriptstyle{\pm 0.2}$
-EfficientZero [21] $10\scriptstyle{\pm 0.2}$ $12\scriptstyle{\pm 1.3}$ $15\scriptstyle{\pm 2.3}$ $15\scriptstyle{\pm 0.7}$
+EfficientZero (Ye et al., 2021) $10\scriptstyle{\pm 0.2}$ $12\scriptstyle{\pm 1.3}$ $15\scriptstyle{\pm 2.3}$ $15\scriptstyle{\pm 0.7}$
 Table 7: Average wall-time of ReZero-E vs.
@@
 Mean and standard deviation over 5 runs.
-C.3 Ablations
+D.3 More Ablations
 Figure 10: The results of the ablation study comparing the use of backward-view reanalyze versus its absence. Mean of 3 runs; shaded areas are 95% confidence intervals.
@@
 We ultimately set the batch size to 2000, which yields the fastest average search speed on our device.
+D.4 Toy Case
+We offer the complete code for the toy case.
+Readers can compare the core functions $search()$ and $reuse\_search()$ of the MCTS class to understand how root node values are utilized and compare $select()$ and $reuse\_select()$ to understand how the search process is prematurely halted.
+import random
+import math
+import time
+import numpy as np
+import matplotlib.pyplot as plt
+class Node:
+def __init__(self, state, parent=None):
+self.state = state
+self.parent = parent
+self.children = []
+self.visits = 0
+self.value = 0
+def is_fully_expanded(self):
+return len(self.children) == len(self.state.get_possible_actions())
+def add_child(self, child_state):
+child = Node(child_state, self)
+self.children.append(child)
+return child
+class MCTS:
+def __init__(self, exploration_weight=1.0):
+self.exploration_weight = exploration_weight
+self.gamma = 0.9
+# the search process in origin MCTS
+def search(self, initial_state, max_iter=100):
+root = Node(initial_state)
+for _ in range(max_iter):
+node = self.select(root)
+reward = self.simulate(node.state)
+self.backpropagate(node, reward)
+best_child = self.get_best_child(root, 0)
+return best_child.state.current_pos, root.value
+# the search process in our accelerated MCTS by reuse the root value
+def reuse_search(self, initial_state, value, action, max_iter=100):
+root = Node(initial_state)
+for _ in range(max_iter):
+node = self.reuse_select(root,action)
+if node.state.current_pos == action:
+reward = value
+else: reward = self.simulate(node.state)
+self.backpropagate(node, reward)
+best_child = self.get_best_child(root, 0)
+return best_child.state.current_pos, root.value
+# select nodes in origin MCTS
+def select(self, node):
+while not node.state.is_terminal():
+if not node.is_fully_expanded():
+return self.expand(node)
+else:
+node = self.get_best_child(node, self.exploration_weight)
+return node
+# select nodes in our accelerated MCTS by reuse the root value
+def reuse_select(self, node,actionpos):
+while not (node.state.is_terminal() or node.state.current_pos == actionpos):
+if not node.is_fully_expanded():
+return self.expand(node)
+else:
+node = self.get_best_child(node, self.exploration_weight)
+return node
+def expand(self, node):
+actions = node.state.get_possible_actions()
+for action in actions:
+if not any(child.state.current_pos == node.state.copy().step(action)[0] for child in node.children):
+new_state = node.state.copy()
+new_state.step(action)
+return node.add_child(new_state)
+return None
+def simulate(self, state):
+sim_state = state.copy()
+count = 1
+while not sim_state.is_terminal():
+action = random.choice(sim_state.get_possible_actions())
+sim_state.step(action)
+count += 1
+return sim_state.get_reward()/count
+def backpropagate(self, node, reward):
+gamma = self.gamma
+while node is not None:
+node.visits += 1
+node.value += reward * gamma
+node = node.parent
+gamma *= self.gamma
+def get_best_child(self, node, exploration_weight):
+best_value = -float(’inf’)
+best_children = []
+for child in node.children:
+exploit = child.value / child.visits
+explore = math.sqrt(2.0 * math.log(node.visits) / child.visits)
+value = exploit + exploration_weight * explore
+if value > best_value:
+best_value = value
+best_children = [child]
+elif value == best_value:
+best_children.append(child)
+return random.choice(best_children)
+class GridWorld:
+def __init__(self):
+self.grid = [[0 for _ in range(4)] for _ in range(4)]
+self.start_pos = (0, 0)
+self.goal_pos = (0, 3)
+self.current_pos = self.start_pos
+self.actions = [’down’, ’up’, ’left’, ’right’]
+def reset(self):
+self.current_pos = self.start_pos
+return self.current_pos
+def step(self, action):
+if action not in self.actions:
+raise ValueError("Invalidaction")
+x, y = self.current_pos
+if action == ’up’:
+x = max(0, x - 1)
+elif action == ’down’:
+x = min(3, x + 1)
+elif action == ’left’:
+y = max(0, y - 1)
+elif action == ’right’:
+y = min(3, y + 1)
+self.current_pos = (x, y)
+reward = 1 if self.current_pos == self.goal_pos else 0
+done = self.current_pos == self.goal_pos
+return self.current_pos, reward, done
+def is_terminal(self):
+return self.current_pos == self.goal_pos
+def get_reward(self):
+return 1 if self.current_pos == self.goal_pos else 0
+def get_possible_actions(self):
+x, y = self.current_pos
+possible_actions = []
+if x > 0:
+possible_actions.append(’up’)
+if x < 3:
+possible_actions.append(’down’)
+if y > 0:
+possible_actions.append(’left’)
+if y < 3:
+possible_actions.append(’right’)
+return possible_actions
+def copy(self):
+new_grid = GridWorld()
+new_grid.current_pos = self.current_pos
+return new_grid
+def render(self):
+for i in range(4):
+for j in range(4):
+if (i, j) == self.current_pos:
+print("A", end="")
+elif (i, j) == self.goal_pos:
+print("G", end="")
+else:
+print(".", end="")
+print()
+print()
+class GridWorldWithWalls:
+def __init__(self):
+self.grid = [[0 for _ in range(7)] for _ in range(7)]
+self.start_pos = (0, 0)
+self.goal_pos = (6, 6)
+self.current_pos = self.start_pos
+self.actions = [’down’, ’up’, ’left’, ’right’]
+self.walls = [(2, 2), (2, 3), (1,1), (3, 2), (3, 4), (3,2), (0,5), (4, 4), (6,3), (5,3)]
+for wall in self.walls:
+self.grid[wall[0]][wall[1]] = 1
+def reset(self):
+self.current_pos = self.start_pos
+return self.current_pos
+def step(self, action):
+if action not in self.actions:
+raise ValueError("Invalidaction")
+x, y = self.current_pos
+if action == ’up’:
+x = max(0, x - 1)
+elif action == ’down’:
+x = min(6, x + 1)
+elif action == ’left’:
+y = max(0, y - 1)
+elif action == ’right’:
+y = min(6, y + 1)
+if (x, y) not in self.walls:
+self.current_pos = (x, y)
+reward = 1 if self.current_pos == self.goal_pos else 0
+done = self.current_pos == self.goal_pos
+return self.current_pos, reward, done
+def is_terminal(self):
+return self.current_pos == self.goal_pos
+def get_reward(self):
+return 1 if self.current_pos == self.goal_pos else 0
+def get_possible_actions(self):
+x, y = self.current_pos
+possible_actions = []
+if x > 0 and (x - 1, y) not in self.walls:
+possible_actions.append(’up’)
+if x < 6 and (x + 1, y) not in self.walls:
+possible_actions.append(’down’)
+if y > 0 and (x, y - 1) not in self.walls:
+possible_actions.append(’left’)
+if y < 6 and (x, y + 1) not in self.walls:
+possible_actions.append(’right’)
+return possible_actions
+def copy(self):
+new_grid = GridWorldWithWalls()
+new_grid.current_pos = self.current_pos
+return new_grid
+def render(self):
+for i in range(7):
+for j in range(7):
+if (i, j) == self.current_pos:
+print("A", end="")
+elif (i, j) == self.goal_pos:
+print("G", end="")
+elif (i, j) in self.walls:
+print("#", end="")
+else:
+print(".", end="")
+print()
+print()
+env = GridWorldWithWalls()
+mcts = MCTS()
+# plot the grid
+plt.figure(figsize=(6, 6))
+plt.matshow(env.grid, cmap=’binary’, fignum=0, vmin=0, vmax=1)
+for i in range(7):
+for j in range(7):
+if (i, j) == env.start_pos:
+plt.text(j, i, ’A’, ha=’center’, va=’center’, color=’black’, fontsize=12)
+elif (i, j) == env.goal_pos:
+plt.text(j, i, ’G’, ha=’center’, va=’center’, color=’black’, fontsize=12)
+elif (i, j) in env.walls:
+plt.text(j, i, ’’, ha=’center’, va=’center’, color=’black’, fontsize=12)
+else:
+plt.text(j, i, ’’, ha=’center’, va=’center’, color=’black’, fontsize=12)
+for wall in env.walls:
+plt.gca().add_patch(plt.Rectangle((wall[1] - 0.5, wall[0] - 0.5), 1, 1, color=’black’))
+for i in range(7):
+for j in range(7):
+plt.gca().add_patch(plt.Rectangle((j - 0.5, i - 0.5), 1, 1, fill=False, edgecolor=’black’, linewidth=2))
+plt.title(’GridWorldEnvironment’, fontsize=24)
+plt.xticks(np.arange(7), [’0’, ’1’, ’2’, ’3’, ’4’, ’5’, ’6’])
+plt.yticks(np.arange(7), [’0’, ’1’, ’2’, ’3’, ’4’, ’5’, ’6’])
+plt.xlabel(’Column’, fontsize=24)
+plt.ylabel(’Row’, fontsize=24)
+plt.show()
+# record the search time
+search_times = np.zeros((7, 7))
+reuse_times = np.zeros((7, 7))
+for i in range(7):
+for j in range(7):
+if (i, j) == env.goal_pos or (i,j) in env.walls:
+continue
+env.current_pos = (i, j)
+print(f"StartingMCTSfromposition:{env.current_pos}")
+start_time = time.time()
+reuse_action, root_value = mcts.search(env)
+end_time = time.time()
+search_time = end_time - start_time
+search_times[i, j] = search_time
+env.current_pos = reuse_action
+if reuse_action == env.goal_pos:
+reuse_value = 1
+else: _, reuse_value = mcts.search(env)
+env.current_pos = (i, j)
+start_time = time.time()
+best_action, root_value = mcts.reuse_search(env, reuse_value, reuse_action)
+end_time = time.time()
+reuse_time = end_time - start_time
+reuse_times[i, j] = reuse_time
+print(f"Bestactionleadstoposition:{reuse_action}")
+print(f"Reusesearchbestactionleadstoposition:{best_action}")
+print(f"Searchtime:{search_time:.4f}seconds")
+[…]
+# plot the time heatmap
+plt.figure(figsize=(6, 6))
+plt.imshow(search_times, cmap=’viridis’, vmin=0, vmax=0.2, interpolation=’nearest’)
+plt.colorbar(label=’SearchTime(seconds)’)
+plt.title(’OriginSearchTime’, fontsize=20)
+plt.xticks(np.arange(7), [’0’, ’1’, ’2’, ’3’, ’4’, ’5’, ’6’])
+plt.yticks(np.arange(7), [’0’, ’1’, ’2’, ’3’, ’4’, ’5’, ’6’])
+plt.xlabel(’Column’, fontsize=20)
+plt.ylabel(’Row’, fontsize=20)
+plt.figure(figsize=(6, 6))
+plt.imshow(reuse_times, cmap=’viridis’, vmin=0, vmax=0.2, interpolation=’nearest’)
+plt.colorbar(label=’SearchTime(seconds)’)
+plt.title(’AcceleratedSearchTime’, fontsize=20)
+plt.xticks(np.arange(7), [’0’, ’1’, ’2’, ’3’, ’4’, ’5’, ’6’])
+plt.yticks(np.arange(7), [’0’, ’1’, ’2’, ’3’, ’4’, ’5’, ’6’])
+plt.xlabel(’Column’, fontsize=20)
+plt.ylabel(’Row’, fontsize=20)
+plt.show()